### v7

Abstract. We prove a central limit theorem for network formation models with strategic interactions and homophilous agents. Since data often consists of observations on a single large network, we consider an asymptotic framework in which the network size diverges. We argue that a modification of “stabilization” conditions from the literature on geometric graphs provides a useful high-level formulation of weak dependence which we utilize to establish an abstract central limit theorem. Using results in branching process theory, we derive interpretable primitive conditions for stabilization. The main conditions restrict the strength of strategic interactions and equilibrium selection mechanism. We discuss practical inference procedures justified by our results.
JEL Codes: C31, C57, D85
Keywords: social networks, strategic interactions, weak dependence, network formation
1 Introduction
Network models have attracted considerable attention as tractable representations of non-market interactions, such as peer effects and social learning, and of economic relationships, such as financial and trade networks. The economic perspective on networks emphasizes the importance of strategic interactions or externalities (Jackson et al., 2021). One strand of the literature studies social interactions, how an individual’s behavior interacts with those of her social contacts. A second strand studies network formation, why an individual chooses particular social contacts, and how their choices affect those of others. In both cases, externalities generate a wedge between choices that are optimal from the individual’s perspective and those that are efficient for society. This has important consequences for policymaking, for instance motivating policies of “associational redistribution” that intervene on the network structure (Durlauf, 1996). However, when preferences are misaligned with the policy objective, the endogenous response to a policy intervention may diverge from the intended outcome (Carrell et al., 2013). It is therefore of interest to develop econometric methodologies for measuring network externalities.
The focus of our paper is models of network formation, which have diverse applications ranging from risk sharing in the rural Philippines (Fafchamps and Gubert, 2007) to research partnerships in the biotechnology industry (Powell et al., 2005). They can be used to distinguish between different theoretical mechanisms for link formation, including preferential attachment (Barabási and Albert, 1999), strategic transitivity (Mele, 2022; Ridder and Sheng, 2022), and homophily (McPherson et al., 2001). They enable policymakers to forecast the effects of counterfactual interventions on a social network (Mele, 2020) and to account for dependence between the network and unobservables of a social interactions model (Badev, 2021).
We augment a class of latent space (Hoff et al., 2002) and geometric graph (Penrose and Yukich, 2003) models with strategic interactions. In such models, nodes are positioned in a possibly latent social or economic space such that the rate of link formation decays with distance in space. However, due to strategic interactions, link formation depends not only on the attributes of the node pair involved in the link but also on links between other pairs. This induces a non-standard form of cross-sectional dependence between network units. The challenge for large-sample theory lies in establishing conditions under which the “amount of independent information” increases with the network size.
We establish a CLT for a large class of network moments computed from a single large network. The moments are averages of node-level statistics $n^{-1}\sum_{i=1}^{n}\psi_{i}$, where $\psi_{i}$ is some function of the network and node types. Examples include the degree of node $i$, which is the number of links involving $i$; the average clustering coefficient; and subnetwork counts which have been utilized for inference in econometric models of network formation (Sheng, 2020). We discuss practical inference procedures justified by the result.
We emphasize two technical contributions. The first is an abstract CLT that holds under high-level conditions, which is an extension of limit theorems from the literature on geometric graphs (e.g. Penrose, 2007). The key condition is a modification of “stabilization” assumptions from that literature that provide a useful formulation of weak dependence for our purposes. Stabilization essentially requires $\psi_{i}$ to only be a function of a random subset of nodes whose size has a distribution with exponential tails. This implies that the size is asymptotically bounded, so node $i$’s statistic only depends on a small fraction of alters. In this sense, $\{\psi_{i}\}_{i=1}^{n}$ is weakly dependent.
In our applications, this random subset involves a union of network components, which are disconnected subnetworks that are challenging combinatorial objects to analyze (a formal definition is provided at the end of this section). Our primary contribution is to develop a methodology for deriving the tail bounds required for stabilization. We employ techniques from branching process theory commonly used to bound component sizes in the literature on random graphs. Using the methodology, we derive interpretable primitive sufficient conditions for stabilization in models of strategic network formation.
One of the main conditions restricts the strength of strategic interactions. These induce cross-sectional dependence since the realization of a link can depend on the existence of “neighboring” links, which in turn can depend on other links, and so on. The longer the lengths of these chains of dependent links, the stronger the degree of cross-sectional dependence. We adopt a well-known technique in random graph theory to bound the lengths of these chains by branching processes whose sizes can more tractably be proven to have exponential tails, provided the processes are “subcritical” or non-explosive in growth. We argue that conditions for subcriticality in our context are analogous to weak dependence conditions imposed on spatial and temporal autoregressive models that bound the magnitude of the autoregressive parameter below one.
The other main condition is a restriction on the equilibrium selection mechanism. Even if strategic interactions are sufficiently weak, strong cross-sectional dependence may exist if nodes can globally “coordinate” on the equilibrium network through a common signal such as the type of a particular node. We instead require equilibrium selection to be sufficiently “decentralized,” which holds for instance under myopic best-response dynamics. This is the single-network analog of the usual requirement under many-network asymptotics that equilibrium selection is independent across networks.
A growing literature studies frequentist inference in network formation models when the econometrician observes a single network. Leung (2019b) and Menzel (2024) develop laws of large numbers for models of strategic network formation. The former paper modifies a weaker stabilization condition due to Penrose and Yukich (2003) and uses branching processes to derive primitive sufficient conditions. Our paper tackles the more difficult problem of obtaining a normal approximation, which naturally requires a stronger stabilization condition, and discusses practical inference procedures newly justified by the result. We prove the CLT by adapting results from the literature on geometric graphs (Penrose and Yukich, 2008; Penrose, 2007); a more detailed discussion of our contributions relative to this literature can be found in § 3.3. Kuersteiner (2019) takes a different approach, using a novel conditional mixingale type assumption defined in terms of a random metric of distance.
Leung (2015) and Ridder and Sheng (2022) study strategic network formation under incomplete information. In this setting, links are independent conditional on observables, whereas the models we study can be microfounded as games of complete information. These allow for unobserved heterogeneity, which generates dependence between potential links even conditional on observables.
Charbonneau (2017), Dzemski (2019), Graham (2017), and Jochmans (2018) consider dyadic link formation models without strategic interactions but allow for node-level fixed effects. A large literature in statistics studies models without strategic interactions, for example stochastic block models (Bickel et al., 2011) and latent-space models (Hoff et al., 2002). These are useful for their parsimony and tasks such as community detection. Chandrasekhar and Jackson (2021) and Boucher and Mourifié (2017) study statistical models allowing for interdependence between links.
Kojevnikov et al. (2021) prove a CLT for node-level data conditional on the network. This does not apply to network formation models since the network is the outcome. We prove an unconditional CLT that may be applied to network formation as well as network processes (Leung, 2019a).
The next section presents a model of strategic network formation and defines network moments. In § 3, we state high-level conditions for a CLT and outline its proof. Readers interested in the low-level CLT for strategic network formation may skip § 3 for § 4, which presents primitive sufficient conditions. We also outline a general methodology for verifying the high-level condition that can be applied to other network models. In § 5, we discuss practical inference procedures. We present results from a simulation study in § 6, and § 7 concludes. All proofs are stated in the supplemental appendix.
We introduce standard notation and terminology for networks. We represent a network on a set of $n$ nodes by an $n\times n$ adjacency matrix, where the $ij$th entry $A_{ij}$, termed the potential link, is an indicator for whether nodes $i,j$ are connected. We assume $A_{ii}=0$ for all nodes $i$, meaning that there are no self-links, and we focus on undirected networks, so $A_{ij}=A_{ji}$. For two networks $\bm{A},\bm{A}^{\prime}$, we say that $\bm{A}$ is a subnetwork of $\bm{A}^{\prime}$ if every link in $\bm{A}$ is a link in $\bm{A}^{\prime}$. A path in a network from node $i$ to $j$ is a sequence of distinct nodes starting with $i$ and ending with $j$ such that for each consecutive node pair $k,k^{\prime}$ in this sequence, $A_{kk^{\prime}}=1$. Its length is the number of links it involves. The path distance between two nodes $i\neq j$ in $\bm{A}$ is the length of the shortest path that connects them if a path exists and $\infty$ otherwise. The path distance between a node and itself is defined as zero. The $K$-neighborhood of a node $i$ in $\bm{A}$, denoted by $\mathcal{N}_{\bm{A}}(i,K)$, is the set of nodes of path distance at most $K$ from $i$. Finally the component of a node $i$ in a network $\bm{A}$ is the set of nodes at finite path distance from $i$.
2 Model
Let $\mathcal{N}_{n}=\{1,\dots,n\}$ be a set of nodes, and endow each $i\in\mathcal{N}_{n}$ with an i.i.d. vector-valued type $(X_{i},Z_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}$. We distinguish $X_{i}$ as the position of node $i$, a continuously distributed vector of homophilous (defined below) attributes with density $f$ that has bounded support. Endow each node pair $(i,j)$ with an i.i.d. $\mathbb{R}$-valued random utility shock $\zeta_{ij}=\zeta_{ji}$ independent of types.
The network $\bm{A}$ satisfies, for all $i\neq j$,
$A_{ij}=\mathbf{1}\left\{V_{ij}>0\right\}\quad\text{for}\quad V_{ij}\equiv V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,S_{ij},Z_{i},Z_{j},\zeta_{ij}),$ (1)
where the joint-surplus function $V(\cdot)$ is an $\mathbb{R}$-valued function, $\lVert\cdot\rVert$ a norm on $\mathbb{R}^{d}$, and $S_{ij}$ a vector of statistics that captures strategic interactions through its dependence on $\bm{A}$. As discussed below, the sparsity of $\bm{A}$ will in part be determined by the positive constant $r_{n}$. Model (1) corresponds to the well-known pairwise-stability solution concept under transferable utility (Jackson, 2010).
Example 1.
Consider the linear joint surplus function
$V_{ij}=\theta_{1}+\theta_{2}S_{ij}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{ij}\quad\text{with}\quad S_{ij}=\max_{k}A_{ik}A_{jk}.$ (2)
Our theory will require $\theta_{3}>0$ which captures homophily in position since it disincentivizes link formation between positionally dissimilar nodes. The term $S_{ij}$ is an indicator for whether $i$ and $j$ share a common neighbor (used for example by Menzel, 2024). If $\theta_{2}>0$, then node pairs $(i,j)$ sharing a common neighbor are more likely to form a link. Both rationalize the well-known stylized fact that networks are commonly clustered in that nodes with common neighbors are themselves typically neighbors (Jackson, 2010).
Example 2.
Sheng (2020) studies a specification similar to
$V_{ij}=\beta_{0}+(Z_{i}+Z_{j})^{\prime}\beta_{1}+\beta_{2}r_{n}^{-1}\lVert X_{i}-X_{j}\rVert\\ +\gamma_{1}\sum_{k=1}^{n}(A_{ik}+A_{jk})+\gamma_{2}\sum_{k=1}^{n}A_{ik}A_{jk}+\zeta_{ij},$ (3)
which corresponds to linear joint surplus function with
$S_{ij}=\bigg{(}\sum_{k=1}^{n}A_{ik}A_{jk},\sum_{k=1}^{n}A_{ik},\sum_{k=1}^{n}A_{jk}\bigg{)}.$ (4)
The first component of (4) plays a role analogous to $S_{ij}$ in (2), being a count of the number of common neighbors. The second and third components are respectively the degrees (number of neighbors) of $i$ and $j$. If $\gamma_{1}>0$, this captures a popularity effect.
As explained in the next subsection, the assumptions we impose for a CLT in many cases require $S_{ij}$ to be uniformly bounded over $i,j\in\mathcal{N}_{n}$ and $n\in\mathbb{N}$. Existing work on large-network asymptotics for network formation shares this limitation (Boucher and Mourifié, 2017; Menzel, 2024; Ridder and Sheng, 2022). Statistics such as (4) may be modified to satisfy uniform boundedness by truncation, for example $\min\{\sum_{k=1}^{n}A_{ik},\Delta\}$ for some user-specified constant $\Delta$.
The next subsection states a restriction on $V(\cdot)$ that generates homophily and sparsity. In § 2.2, we formalize how $S_{ij}$ may depend on $\bm{A}$ and introduce the equilibrium selection mechanism. Finally, in § 2.3, we define the class of network moments for which we provide a CLT.
2.1 Homophily and Sparsity
A common feature of social networks is homophily, the tendency for those with similar characteristics to associate. We require homophily in position, specifically that the joint-surplus function $V(\cdot)$ is decreasing in the first component, so nodes dissimilar in position are less likely to form links. Positions may represent latent node locations in an abstract “social space,” as in latent-space models (Hoff et al., 2002), or attributes such as income and geographic location.
Another common feature is sparsity, meaning that the number of connections formed by the typical node is of significantly smaller order than $n$ (Barabási, 2015). This is often accomplished by scaling the sequence of models such that the expected degree $n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}{\bf E}[A_{ij}]$ is asymptotically bounded.
Our first assumption restricts the joint-surplus function $V(\cdot)$ and the distribution of $\zeta_{ij}$, so that the network generated by model (1) exhibits homophily in positions and sparsity. Let $\bar{\Phi}_{\zeta}(z)={\bf P}(\zeta_{ij}\geq z)$. Define $\bar{V}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,\zeta_{ij})=\sup_{s,z,z^{\prime}}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,z,z^{\prime},\zeta_{ij})$, where the supremum is taken over $(s,z,z^{\prime})$ in the support of $(S_{ij},Z_{i},Z_{j})$ (which will be identical across $i,j$ under assumptions below). Finally, recall that $d$ is the dimension of $X_{1}$.
Assumption 1 (Homophily and Sparsity).
For any $\delta\in\mathbb{R}_{+}$, $\bar{V}(\delta,\cdot)$ is invertible in its second argument, and its inverse $\bar{V}^{-1}(\delta,\cdot)$ satisfies $\limsup_{\delta\rightarrow\infty}\delta^{-1}\log\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))<0$. Furthermore, there exists $\kappa>0$ such that, for any $n\in\mathbb{N}$,
$r_{n}\equiv(\kappa/n)^{1/d}.$ (5)
Homophily is a consequence of the first sentence because ${\bf P}(A_{ij}=1\mid r_{n}^{-1}\lVert X_{i}-X_{j}\rVert=\delta)\leq\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))$, which is required to decrease exponentially with $\delta$. If $V(\cdot)$ is not too nonlinear in its first component and the distribution of $\zeta_{ij}$ has exponential tails, then this assumption is satisfied, as shown in the next example.
Example 3.
Consider Example 1, and suppose $S_{ij}$ has uniformly bounded support. Then for some universal constant $M$,
${\bf P}(V_{ij}>0\mid r_{n}^{-1}\lVert X_{i}-X_{j}\rVert=\delta)\leq{\bf P}(\underbrace{M-\theta_{3}\delta^{2}+\zeta_{ij}}_{\bar{V}(\delta,\zeta_{ij})}>0)=\bar{\Phi}_{\zeta}(\underbrace{\theta_{3}\delta^{2}-M}_{\bar{V}^{-1}(\delta,0)}),$
which decays to zero exponentially with $\delta$ if $\theta_{3}>0$ and the distribution of $\zeta_{ij}$ has exponential tails.
Example 4.
In the previous example, Assumption 1 implicitly imposes restrictions on the support of $(S_{ij},Z_{i},Z_{j},\zeta_{ij})$. Consider a variant of (2) in which we replace $\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}$ with the “random geometric graph” penalty $\infty\cdot\bm{1}\{r_{n}^{-1}\lVert X_{i}-X_{j}\rVert>c\}$ for some $c>0$, with the convention that $\infty\cdot 0=0$. That is, nodes do not link with alters whose scaled positions are sufficiently far from the ego’s. Then ${\bf P}(V_{ij}>0\mid r_{n}^{-1}\lVert X_{i}-X_{j}\rVert=\delta)\leq\bm{1}\{\delta\leq c\}$ which satisfies Assumption 1 without support restrictions.
Lastly we discuss the role of (5) for sparsity. As $n$ increases, there are more opportunities to form links, which increases expected degree, corresponding to a denser network. By sending $r_{n}$ to zero with $n$, we increase the “cost” of link formation due to homophily, thereby decreasing the expected degree. Our choice of $r_{n}$ balances these two forces to achieve a sparse network in which the expected degree is asymptotically bounded. To see this, notice that expected degree is
$\frac{1}{n}\sum_{i=1}^{n}$ $\sum_{j=1}^{n}{\bf E}[A_{ij}]=(n-1){\bf P}\big{(}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,S_{ij},Z_{i},Z_{j},\zeta_{ij})>0\big{)}$
$\leq(n-1){\bf P}\big{(}\zeta_{ij}>\bar{V}^{-1}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,0)\big{)}$
$=(n-1)r_{n}^{d}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}{\bf P}\big{(}\zeta_{ij}>\bar{V}^{-1}(\lVert x-x^{\prime}\rVert,0)\big{)}f(x)f(x+r_{n}(x^{\prime}-x))\,\text{d}x\,\text{d}x^{\prime}$ (6)
by a change of variables $x^{\prime}\mapsto x+r_{n}(x^{\prime}-x)$. By (5), if $f$ is continuous, this converges to
$\kappa\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\bar{\Phi}_{\zeta}\big{(}\bar{V}^{-1}(\lVert x-x^{\prime}\rVert,0)\big{)}f(x)^{2}\,\text{d}x^{\prime}\,\text{d}x,$
which is finite because the integrand decays exponentially with $\lVert x-x^{\prime}\rVert$ by Assumption 1.
2.2 Strategic Interactions and Equilibrium Selection
We next define the statistics $S_{ij}$ that capture strategic interactions. For any $H\subseteq\mathcal{N}_{n}$, let $r_{n}^{-1}\bm{T}_{H}=((r_{n}^{-1}X_{i},Z_{i}))_{i\in H}$ be the array of (scaled) types for nodes in $H$ and $\bm{\zeta}_{H}=(\zeta_{ij})_{i,j\in H}$ (with $\zeta_{ii}\equiv 0$ for all $i$) be the corresponding array of random-utility shocks. In the case where $H=\mathcal{N}_{n}$, we abbreviate $r_{n}^{-1}\bm{T}_{n}\equiv r_{n}^{-1}\bm{T}_{\mathcal{N}_{n}}$ and $\bm{\zeta}_{n}\equiv\bm{\zeta}_{\mathcal{N}_{n}}$. Let $\bm{A}_{H}=(A_{ij})_{i,j\in H}$ denote the subnetwork of $\bm{A}$ on $H$. We define
$S_{ij}=S_{n}(i,j,r_{n}^{-1}\bm{T}_{n},\bm{A})$ (7)
where $\{S_{n}\}_{n\in\mathbb{N}}$ is a sequence of permutation-invariant functions. [^1] For any injection $\pi\colon\mathcal{N}_{n}\rightarrow\mathcal{N}_{n}$, we abuse notation by defining $\pi(r_{n}^{-1}\bm{T}_{n})=((r_{n}^{-1}X_{\pi(i)},Z_{\pi(i)}))_{i\in\mathcal{N}_{n}}$ and $\pi(\bm{A})=(A_{\pi(i)\pi(j)})_{i,j}$. We say $S_{n}(\cdot)$ is permutation-invariant if $S_{n}(i,j,r_{n}^{-1}\bm{T}_{n},\bm{A})=S_{n}(\pi(i),\pi(j),\pi(r_{n}^{-1}\bm{T}_{n}),\pi(\bm{A}))$. This means that the function does not directly depend on node labels which have no economic content in our model.
We next impose the common restriction that $S_{ij}$ only depends on its arguments through the 1-neighborhoods of $i$ and $j$, recalling the definition from § 1.
Assumption 2 (Local Externalities).
For any $r>0$, $n\in\mathbb{N}$, $i,j\in\mathcal{N}_{n}$, and $H=\mathcal{N}_{\bm{A}}(i,1)\cup\mathcal{N}_{\bm{A}}(j,1)$, $S_{n}(i,j,r^{-1}\bm{T}_{n},\bm{A})=S_{\lvert H\rvert}(i,j,r^{-1}\bm{T}_{H},\bm{A}_{H})$.
That is, $S_{ij}$ is only a function of nodes linked to either $i$ or $j$. Most statistics used in the literature satisfy this assumption, including those in Examples 1 and 2, which appear to be the most common choices (Christakis et al., 2020; Leung, 2019b; Mele, 2017; Sheng, 2020). Dependence of $S_{n}(\cdot)$ on types also allows for covariate-weighted versions of these examples, such as the weighted degree $\sum_{k=1}^{n}A_{ik}Z_{k}$.
The model thus far is incomplete since multiple networks $\bm{A}$ may satisfy (1) due to strategic interactions. Let $\mathcal{E}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ denote the set of such networks, which corresponds to the set of pairwise stable equilibria. To complete the model, we introduce a selection mechanism, which is the reduced-form mapping from primitives to the observed network. It is a representation of the latent social process by which nodes coordinate on an equilibrium.
Assumption 3 (Equilibrium Selection).
For any $r>0$ and $n\in\mathbb{N}$, an equilibrium exists in that $\mathcal{E}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ is non-empty, and there exists a permutation-equivariant function $\lambda_{n}(\cdot)$ (the equilibrium selection mechanism) such that $\bm{A}=\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})\in\mathcal{E}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$. [^2] Following the notation in footnote 1, we say $\lambda_{n}(\cdot)$ is permutation-equivariant if $\pi(\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n}))=\lambda_{n}(\pi(r^{-1}\bm{T}_{n}),\pi(\bm{\zeta}_{n}))$, where $\pi(\bm{\zeta}_{n})$ is defined similarly to $\pi(\bm{A})$.
Remark 1.
The empirical games literature typically represents the equilibrium selection mechanism as a conditional distribution $\sigma(\cdot\mid r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ over $\mathcal{E}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$. We represent it as a deterministic function, which is more convenient for our purposes, especially for formulating Assumption 8 below. We next show that for any conditional distribution, we can construct a deterministic function that induces the same distribution over equilibria. The main idea is that types can always include payoff-irrelevant components that can be used to randomize over equilibria. Specifically, let $\{\nu_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{U}([0,1])$ be independent of all structural primitives. These will serve to generate a distribution over equilibria conditional on structural primitives. Redefine $r^{-1}\bm{T}_{n}=((r^{-1}X_{i},Z_{i},\nu_{i}))_{i\in\mathcal{N}_{n}}$ (scaled types with payoff-irrelevant components), and let $r^{-1}\tilde{\bm{T}}_{n}=((r^{-1}X_{i},Z_{i}))_{i\in\mathcal{N}_{n}}$ (scaled types as originally defined).
The set of equilibria is given by $\mathcal{E}(r^{-1}\tilde{\bm{T}}_{n},\bm{\zeta}_{n})=\{\bm{A}_{k}\}_{k=1}^{m}$. Consider any conditional distribution $\sigma$ over this set, where we abbreviate $\sigma_{k}\equiv\sigma(\bm{A}_{k}\mid r^{-1}\tilde{\bm{T}}_{n},\bm{\zeta}_{n})$. Define $\nu^{*}=F(1-\max\{\nu_{i}\}_{i=1}^{n})$, where $F$ is the CDF of $1-\max\{\nu_{i}\}_{i=1}^{n}$. Partition $[0,1]$ into $m$ intervals of lengths $\sigma_{1},\ldots,\sigma_{m}$, and let $\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ be the function that selects equilibrium $\bm{A}_{k}$ if $\nu^{*}$ falls within the interval associated with $\sigma_{k}$. Since $\nu^{*}\sim\mathcal{U}([0,1])$ for any $n$, we have ${\bf P}(\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})=\bm{A}_{k}\mid r^{-1}\tilde{\bm{T}}_{n},\bm{\zeta}_{n})=\sigma_{k}$. In other words, our deterministic construction of $\lambda_{n}(\cdot)$ generates the desired conditional distribution over equilibria.
2.3 Network Moments
Our objective is to prove a CLT for network moments that are averages of node statistics
$\frac{1}{n}\sum_{i=1}^{n}\psi_{n}(i,r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n},\bm{A}),$
where $\{\psi_{n}\}_{n\in\mathbb{N}}$ is a sequence of $\mathbb{R}^{d_{\psi}}$-valued, permutation-invariant functions (see footnote 1). We often abbreviate $\psi_{i}(\mathcal{N}_{n})\equiv\psi_{n}(i,r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n},\bm{A})$.
Example 5 (Subnetwork Counts).
A simple network moment is the average degree for which the node statistic is the degree $\psi_{i}(\mathcal{N}_{n})=\sum_{j=1}^{n}A_{ij}$. This is a permutation-invariant function of the network. Average degree is proportional to the dyad count (number of links). More generally, we can count any other connected subnetwork, such as the number of triangles, $k$-stars, or complete networks on $k$-tuples. For instance, the triangle count is proportional to $\sum_{i,j,k}A_{ij}A_{jk}A_{ik}$, with corresponding node statistic $\psi_{i}(\mathcal{N}_{n})=\sum_{j,k}A_{ij}A_{jk}A_{ik}$. See § A for a formal definition of subnetwork counts.
More generally, we consider the following class of node statistics that includes the previous examples. Recall that $\mathcal{N}_{\bm{A}}(i,K)$ is $i$’s $K$-neighborhood, defined in § 1.
Assumption 4 (Node Statistics).
For some $K\in\mathbb{Z}_{+}$ and any $r>0$, $n\in\mathbb{N}$, $i\in\mathcal{N}_{n}$, and $H=\mathcal{N}_{\bm{A}}(i,K)$, $\psi_{n}(i,r^{-1}\bm{T}_{n},\bm{\zeta}_{n},\bm{A})=\psi_{\lvert H\rvert}(i,r^{-1}\bm{T}_{H},\bm{\zeta}_{H},\bm{A}_{H})$.
This states that the node statistic only depends on the types, random-utility shocks, and subnetwork of nodes on $i$’s $K$-neighborhood. The average degree and triangle count satisfy this for $K=1$. A more complex example is the number of nodes at most path distance $D$ from node $i$, which satisfies this assumption for $K=D$.
Example 6.
Subnetwork counts are useful for structural inference. Suppose $V(\cdot)$ is known up to some vector of parameters $\theta_{0}$. Sheng (2020) defines an identified set for $\theta_{0}$ in terms of moment inequalities of the form
$\frac{1}{n}\sum_{i_{1}=1}^{n}\dots\sum_{i_{m}=1}^{n}{\bf E}[(G_{i_{1},\ldots,i_{m}}-H_{i_{1},\ldots,i_{m}}(\theta_{0}))q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})]\leq 0.$ (8)
Here $\sum_{i_{1}=1}^{n}\ldots\sum_{i_{m}=1}^{n}G_{i_{1},\ldots,i_{m}}$ is proportional to a count of a particular connected subnetwork of size $m$. For example, for counting triangles, which are fully connected subnetworks with $m=3$, $G_{i_{1},\ldots,i_{m}}=A_{i_{1}i_{2}}A_{i_{2}i_{3}}A_{i_{3}i_{1}}$. The upper bound $H_{i_{1},\ldots,i_{m}}(\theta_{0})$ and instrument function $q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})$ are known, deterministic functions of the observed component of $\bm{T}_{n}$.
Furthermore, $n^{-1}\sum_{i_{1}=1}^{n}\ldots\sum_{i_{m}=1}^{n}H_{i_{1},\ldots,i_{m}}(\theta_{0})q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})$ is proportional to a U-statistic of order $m$ with a kernel that is a deterministic function of $\{(X_{i_{k}},Z_{i_{k}})\}_{k=1}^{m}$. By the Hoeffding decomposition, it equals $n^{-1}\sum_{i=1}^{n}J_{i}(\theta_{0})+o_{p}(n^{-1/2})$ for some $J_{i}(\theta_{0})$ with the same mean that is a deterministic function of $i$’s type $(X_{i},Z_{i})$. Hence, $J_{i}(\theta_{0})$ satisfies Assumption 4 for $K=0$. Additionally, $\tilde{\psi}_{i_{1}}(\mathcal{N}_{n})\equiv\sum_{i_{2}=1}^{n}\dots\sum_{i_{m}=1}^{n}G_{i_{1},\ldots,i_{m}}q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})$ is a node statistic satisfying Assumption 4 for some $K\leq m-1$ since the subnetwork is connected. We have therefore shown that
$\frac{1}{\sqrt{n}}\sum_{i_{1}=1}^{n}\dots\sum_{i_{m}=1}^{n}\big{[}(G_{i_{1},\ldots,i_{m}}-H_{i_{1},\ldots,i_{m}}(\theta_{0}))q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})\\ -{\bf E}[(G_{i_{1},\ldots,i_{m}}-H_{i_{1},\ldots,i_{m}}(\theta_{0}))q_{i_{1},\ldots,i_{m}}(\bm{T}_{n})]\big{]}\\ =\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\big{(}\underbrace{\tilde{\psi}_{i}(\mathcal{N}_{n})-J_{i}(\theta_{0})}_{\psi_{i}(\mathcal{N}_{n})}-{\bf E}[\tilde{\psi}_{i}(\mathcal{N}_{n})-J_{i}(\theta_{0})]\big{)}+o_{p}(1).$
Our CLT can be applied to the first term on the last line.
3 Stabilization
This section provides high-level conditions for a CLT. Readers interested in the application to strategic network formation may wish to skip to § 4 where we state primitive sufficient conditions.
Recall from Assumption 3 that for any $H\subseteq\mathcal{N}_{n}$, $\lambda_{\lvert H\rvert}(r_{n}^{-1}\bm{T}_{H},\bm{\zeta}_{H})$ is the equilibrium network under the counterfactual model in which the set of nodes is $H$ rather than $\mathcal{N}_{n}$. Abusing notation, for $i\in H$, we define $i$’s counterfactual node statistic
$\psi_{i}(H)\equiv\psi_{\lvert H\rvert}\big{(}i,r_{n}^{-1}\bm{T}_{H},\bm{\zeta}_{H},\lambda_{\lvert H\rvert}(r_{n}^{-1}\bm{T}_{H},\bm{\zeta}_{H})\big{)}.$ (9)
For $H=\mathcal{N}_{n}$, this coincides with the original definition. For $H\subset\mathcal{N}_{n}$, the counterfactual interpretation is due entirely to the last argument. Contrast (9) with
$\psi_{\lvert H\rvert}(i,r_{n}^{-1}\bm{T}_{H},\bm{\zeta}_{H},\bm{A}_{H}).$ (10)
For example, if $\psi_{i}(\mathcal{N}_{n})$ is the degree of node $i$ on the full network, then (10) is $i$’s degree in the observed subnetwork restricted to $H$, which can be directly computed from observation of $\bm{A}$. Generally, this does not equal (9) because the networks $\bm{A}_{H}$ and $\lambda_{\lvert H\rvert}(r_{n}^{-1}\bm{T}_{H},\bm{\zeta}_{H})$ do not coincide due to strategic interactions, which induce nodes in $H$ to form different links when nodes $\H\mathcal{N}_{n}\backslash$ H are absent from the model.
Recall that $d$ is the dimension of $X_{1}$. For any $B\subseteq\mathbb{R}^{d}$, let $\mathcal{N}_{n}(B)=\{i\in\mathcal{N}_{n}\colon r_{n}^{-1}X_{i}\in B\}$, the set of nodes with scaled positions lying in the set $B$. For $x\in\mathbb{R}^{d}$ and $R\geq 0$, let $B(x,R)=\{y\in\mathbb{R}^{d}\colon\lVert x-y\rVert\leq R\}$.
Definition 1.
For any $r>0$, $n\in\mathbb{N}$, and $i\in\mathcal{N}_{n}$, the radius of stabilization
$\bm{R}_{i}(n,r)\equiv\bm{R}(i,r^{-1}\bm{T}_{n},\bm{\zeta}_{n},\lambda_{n})$
is the smallest integer-valued $R\geq 0$ such that $\psi_{i}(\mathcal{N}_{n})=\psi_{i}(\mathcal{N}_{n}(\mathcal{B}_{i}))$ for all sets $\mathcal{B}_{i}\subseteq\mathbb{R}^{d}$ containing $B(r^{-1}X_{i},R)$.
This is the smallest radius $R$ such that $i$’s node statistic has the same value under all counterfactual models that drop nodes positioned outside of $i$’s neighborhood $B(r^{-1}X_{i},R)$.
The main idea is as follows. If this radius were “small,” then $\psi_{i}(\mathcal{N}_{n})$ would primarily depend on a small fraction of nodes, specifically those positioned near $i$, in which case node statistics should be weakly dependent. To establish a law of large numbers (LLN), Leung (2019b) requires the radius to be $O_{p}(1)$, analogous to a condition due to Penrose and Yukich (2003). However, a CLT seems to require stronger conditions on the distribution of the radius, in particular the thickness of its tails (Penrose, 2007).
3.1 Abstract CLT
We first define some notation and the sequence along which we take limits. Let $\{(X_{i},Z_{i})\}_{i\in\mathbb{N}}$ and $\{\zeta_{ij}\colon i\in\mathcal{N}_{n},j>i\}$ both be i.i.d. and independent, and assume the density $f$ of $X_{1}$ is bounded. We consider a sequence indexed by $n\in\mathbb{N}$ such that the $n$th element of the sequence is given by the tuple
$(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n},\lambda_{n},\psi_{n}).$
The first three components are structural primitives that determine $\bm{A}$, while the last is the node statistic function.
For technical reasons, we need to consider sequences of models indexed by $n$ in which the network size is not $n$ but rather of the same asymptotic order as $n$. Let $\{N_{n}\}_{n\in\mathbb{N}}$ be a sequence of random variables independent of all model primitives such that
$N_{n}\sim\text{Poisson}(n).$ (11)
A de-Poissonization argument discussed below requires us to consider two different network sizes: a random size $N_{n}+k$, where $k$ is a constant, and a non-random size $m+k$, where $m=m_{n}$ with $m_{n}/n\rightarrow c\in(0,\infty)$. Our original setup with $n$ nodes corresponds to $m=n$ and $k=0$, and no intuition is lost to the reader who only considers this case. The need to consider models with different network sizes is purely for technical reasons clarified in § 3.2.
Define $\mathcal{T}=\text{supp}(X_{1},Z_{1})$, $\mathcal{T}^{k}=\bigtimes_{i=1}^{k}\mathcal{T}$, and $\bm{T}_{k}=((X_{i},Z_{i}))_{i=1}^{k}$. Our first assumption is the key high-level weak dependence condition that controls the tail behavior of the distribution of the radius of stabilization, or more formally the conditional probability ${\bf P}(\bm{R}_{1}(m+k,r_{n})>w\mid\bm{T}_{k}=\bm{t}_{k})$. Since probabilities conditional on continuous random variables are not uniquely defined, we require some additional notation to phrase the assumption in terms of a version of the conditional probability. Let $m,k\in\mathbb{N}$ and $\bm{t}_{k}\in\mathcal{T}^{k}$. Construct $r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k})$ from $r_{n}^{-1}\bm{T}_{m+k}$ by replacing the types of the first $k$ nodes $\bm{T}_{k}$ with fixed values $\bm{t}_{k}$. Define $\bm{R}_{1}^{\star}(m+k,r_{n};\bm{t}_{k})\equiv\bm{R}(1,r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k}),\bm{\zeta}_{m+k},\lambda_{m+k})$. Then ${\bf P}(\bm{R}_{1}^{\star}(m+k,r_{n};\bm{t}_{k})>w)$ is a version of the conditional probability ${\bf P}(\bm{R}_{1}(m+k,r_{n})>w\mid\bm{T}_{k}=\bm{t}_{k})$.
Assumption 5 (Exponential Stabilization).
For any $k\in\{1,\dots,4\}$ and $k^{\prime}\in\{1,2\}$, there exist $n_{0},\epsilon>0$ and $\eta\in(0,1]$ such that
$\limsup_{w\rightarrow\infty}w^{-\eta}\max\{\log\tau_{b,\epsilon}(w),\log\tau_{p}(w)\}<0,\quad\text{where}$
$\tau_{b,\epsilon}(w)=\sup_{n>n_{0}}\sup_{m\in((1-\epsilon)n,(1+\epsilon)n)}\sup_{\bm{t}_{k}\in\mathcal{T}^{k}}{\bf P}\big{(}\bm{R}_{1}^{\star}(m+k,r_{n};\bm{t}_{k})>w\big{)},$
$\tau_{p}(w)=\sup_{n>n_{0}}\sup_{\bm{t}_{k^{\prime}}\in\mathcal{T}^{k^{\prime}}}{\bf P}\big{(}\bm{R}_{1}^{\star}(N_{n}+k^{\prime},r_{n};\bm{t}_{k^{\prime}})>w\big{)}.$
For $\eta=1$, this says that the distribution of the radius of stabilization has an exponential tail under models with $N_{n}+k$ and $m+k$ nodes, uniformly over the types of the first $k$ nodes. [^4] The first $k$ is an arbitrary choice of $k$ nodes. Note that nodes are exchangeable in our model since types and random-utility shocks are i.i.d. and the equilibrium selection mechanism is permutation-equivariant. Technically the assumption allows for slower than exponential decay with $\eta<1$, but it is otherwise analogous to conditions used by Penrose and Yukich (2008) and Penrose (2007). Our method of proof for the result below is based on theirs but with important differences in setup discussed in § 3.3.
Contrast Assumption 5 with the more familiar concept of $m$-dependence, which states that an observation is only correlated with nodes in some non-random neighborhood of known radius $m$. Stabilization generalizes this to allow neighborhoods to be node-specific, random, and complex functions of the primitives. In general, bounding the size of this set is far from trivial, and one of our main contributions is to demonstrate that branching processes can be used for this purpose to derive primitive sufficient conditions, as discussed in § 4.5.
[…]
Assumption 6.
(a) There exist $p>2,\epsilon>0$, $M<\infty$, and $n_{0}\in\mathbb{N}$ such that for all $n>n_{0}$, $k\in\{1,2,3\}$, $m\in((1-\epsilon)n,(1+\epsilon)n)$, and $\bm{t}_{k}\in\mathcal{T}^{k}$,
$\max\big{\{}{\bf E}[\lVert\psi_{1}^{\star}(\mathcal{N}_{m+k};\bm{t}_{k})\rVert_{\infty}^{p}],{\bf E}[\lVert\psi_{1}^{\star}(\mathcal{N}_{N_{n}+k}\cap(H_{n}\cup\{1\});\bm{t}_{k})\rVert_{\infty}^{p}]\big{\}}<M$
for any sequence of sets $H_{n}\subseteq\mathbb{N}$. (b) For any $r>0$, $n\in\mathbb{N}$, $i\in\mathcal{N}_{n}$, and $x\in\mathbb{R}^{d}$,
$\psi_{n}(i,r^{-1}\bm{T}_{n},\bm{\zeta}_{n},\bm{A})=\psi_{n}\big{(}i,((r^{-1}X_{j}+x,Z_{j}))_{j=1}^{n},\bm{\zeta}_{n},\bm{A}\big{)}\quad\text{and}$
$\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})=\lambda_{n}\big{(}(r^{-1}X_{j}+x,Z_{j}))_{j=1}^{n},\bm{\zeta}_{n}\big{)}.$
Part (a) requires node statistics to have bounded $p>2$ moments, uniformly over the types of $k$ nodes. We provide primitive conditions in § A for the case of subnetwork counts (Example 5).
Part (b) says that node statistics and the selection mechanism are invariant to additive shifts in scaled positions $\{r^{-1}X_{i}\}_{i=1}^{n}$, which holds if these only enter through scaled distances $r^{-1}\lVert X_{i}-X_{j}\rVert$. Because $X_{i}$ and $Z_{i}$ may be arbitrarily dependent, this still allows $X_{i}$ (unscaled) to enter the model as a subvector of $Z_{i}$. In Examples 1 and 2, the latent index only depends on types through scaled distances or $Z_{i}$, so requiring the selection mechanism and node statistics to satisfy the same property does not apparently rule out any economically interesting applications. [^5] For example, Assumption 6(b) disallows network moments such as $n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}r^{-1}X_{j}A_{ij}$, but since $X_{i}$ and $Z_{i}$ may be arbitrarily correlated, it allows for $n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}X_{j}A_{ij}$.
Let $\bm{\Sigma}_{n}=n^{-1}\text{Var}(\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n}))$, $\lambda_{\text{min}}(\bm{\Sigma}_{n})$ be its smallest eigenvalue, and $\bm{I}$ denote the $d_{\psi}$-dimensional identity matrix.
Theorem 1.
Under Assumptions 5 and 6, $\sup_{n}\lVert\bm{\Sigma}_{n}\rVert_{\infty}<\infty$. Further suppose that $\liminf_{n\rightarrow\infty}\lambda_{\text{min}}(\bm{\Sigma}_{n})>0$. Then
$\bm{\Sigma}_{n}^{-1/2}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\big{(}\psi_{i}(\mathcal{N}_{n})-{\bf E}[\psi_{i}(\mathcal{N}_{n})]\big{)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(\bm{0},\bm{I}).$ (12)
Proof. See § SA.1.
The proof does not rely on the setup or assumptions in § 2, other than the distributional assumptions at the top of § 2 and the requirement that $\lambda_{n}(\cdot)$ is a deterministic, permutation-equivariant function of the structural primitives. The other assumptions will be used in the next section to verify Assumption 5 for strategic network formation. [^6] Theorem 1 can be applied to other network models. For instance, in Leung (2019a), $\lambda_{n}(\cdot)$ is the reduced-form mapping that takes as input the structural primitives and outputs both the network and the outcome of a second-stage social interactions model given the network.
3.2 Outline of Proof
Step 1. We first establish a CLT for the “Poissonized” model in which the number of nodes is $N_{n}$ defined in (11), so-called because $\{X_{i}\}_{i=1}^{N_{n}}$ has the same distribution as a Poisson point process with intensity function $nf(\cdot)$ (Penrose, 2003, Proposition 1.5). Specifically, we show
$\tilde{\bm{\Sigma}}_{n}^{-1/2}\frac{1}{\sqrt{n}}\left(\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})-{\bf E}\left[\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})\right]\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(\bm{0},\bm{I}),$ (13)
where $\tilde{\bm{\Sigma}}_{n}=n^{-1}\text{Var}(\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}}))$. This is simpler to prove directly because, unlike $\{X_{i}\}_{i=1}^{n}$, the Poisson process $\{X_{i}\}_{i=1}^{N_{n}}$ possesses a well-known spatial independence property: for any disjoint subsets $S_{1},S_{2}\subseteq\mathbb{R}^{d}$, $\lvert\{X_{i}\}_{i=1}^{N_{n}}\cap S_{1}\rvert\perp\!\!\!\perp\lvert\{X_{i}\}_{i=1}^{N_{n}}\cap S_{2}\rvert$.
We prove (13) by adapting a theorem due to Penrose and Yukich (2008) (see our Theorem SA.1.1). We proceed by first partitioning the support of $X_{1}$ into cubes $Q_{1},\dots,Q_{V_{n}}$ of slowly growing volume and representing the moment $n^{-1}\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{N_{n}})$ as a double sum over cubes and nodes within cubes $n^{-1}\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})$, where $N_{i}$ is the number of nodes positioned in $Q_{i}$ and $ij$ indexes the $j$th node in $Q_{i}$. Spatial independence of the Poisson process implies independence of node positions across cubes. However, node statistics are complex functionals of the process and hence are not generally independent across cubes.
Since stabilization implies that node statistics $\psi_{ij}(\mathcal{N}_{n})$ are primarily determined by nodes relatively proximate to $ij$, and hence cubes relatively proximate $Q_{i}$, it can be shown that the dependence structure of $\{\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})\}_{i=1}^{V_{n}}$ is “approximately” characterized by a certain “dependency graph.” This is a network in which, roughly speaking, two observations are linked if and only if they are dependent. In our case, observations are cubes, which we connect if and only if they are relatively proximate. Careful construction of the cubes ensures a small approximation error for the dependency graph characterization, and applying a CLT for dependency graphs (Chen and Shao, 2004) delivers the result. For precise details, see § SA.1.1.
Step 2. Since $N_{n}/n\stackrel{{\scriptstyle p}}{{\longrightarrow}}1$, $\{X_{i}\}_{i=1}^{n}$ and $\{X_{i}\}_{i=1}^{N_{n}}$ should be “similar,” so given (13), we expect a similar result for the original model. The second “de-Poissonization” step of the proof, which follows Penrose (2007), shows that this intuition is correct, provided we properly adjust $\tilde{\bm{\Sigma}}_{n}$ downward to obtain the correct variance $\bm{\Sigma}_{n}$ (see our Theorem SA.1.2). This is needed because $N_{n}$ contributes additional randomness to the asymptotic distribution.
More specifically, define the add-one cost
$\Xi_{n}=\psi_{n+1}(\mathcal{N}_{n+1})+\sum_{i=1}^{n}\big{(}\psi_{i}(\mathcal{N}_{n+1})-\psi_{i}(\mathcal{N}_{n})\big{)}.$ (14)
This is the aggregate counterfactual impact on the total $\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$ from adding a new node labeled $n+1$ to the model. The first term is the direct effect of adding $n+1$, which is its own node statistic. The second term is the indirect effect, which is the new node’s impact on the statistics of all other nodes. A key step of the proof establishes that
$n^{-1/2}\left(\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})-{\bf E}\left[\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})\right]\right)\\ =n^{-1/2}\left(\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})-{\bf E}\left[\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})\right]\right)\\ +n^{-1/2}(N_{n}-n){\bf E}[\Xi_{N_{n}}]+o_{p}(1).$
This may be viewed as a first-order expansion in the number of nodes, comparing $N_{n}$ to $n$. The “derivative” is ${\bf E}[\Xi_{N_{n}}]$ since it captures the change in moments as a result of a unit increment in the number of nodes. By (13), the left-hand side is asymptotically normal, and by the Poisson CLT, so is $n^{-1/2}(N_{n}-n){\bf E}[\Xi_{N_{n}}]$. Because $N_{n}$ is independent of all other primitives, we can then establish that the first term on the right-hand side is asymptotically normal by an argument using characteristic functions; for details see the end of § SA.1.2.
3.3 Related Literature
The proof is closely based on arguments in Penrose and Yukich (2008) and Penrose (2007), whose results pertain to geometric graphs without strategic interactions. The innovation in Theorem 1 is primarily conceptual, namely, the recognition that an appropriate modification of stabilization allows us to adapt their results to econometric models. Our main technical innovation will be discussed in § 4, namely the use of branching processes to derive primitive conditions for stabilization in models with strategic interactions. Leung (2019b) uses branching processes to establish an LLN, while we tackle the more difficult task of proving a CLT, which requires us to establish new tail bounds for the radius of stabilization (Lemma SA.2.8).
The setup and assumptions used in Penrose’s work are not directly applicable to our setting, so we cannot simply verify their conditions. However, we show that their proofs can be translated to our setting, which differs in three main aspects. The first is the definition of the radius of stabilization. We reformulate the definition in terms of counterfactual models (9) and require invariance of $i$’s node statistic to the removal of nodes outside of $\mathcal{N}_{n}(\mathcal{B}_{i})$. Existing definitions demand invariance to the removal and addition of new nodes, but invariance to addition is typically violated in the models we study due to strategic interactions. Second, $X_{i}$ may be correlated with $Z_{i}$ in our setup, whereas the literature requires independence, but this turns out to have little effect on the proofs. Third, our model includes pair-specific shocks $\zeta_{ij}$, which pose little problem due to their high degree of independence. These are independent across pairs, in contrast to type pairs $((X_{i},Z_{i}),(X_{j},Z_{j}))$ which are correlated across pairs sharing a common node, for example $(i,j)$ and $(i,k)$.
4 CLT for Network Formation
Exponential stabilization (Assumption 5) provides a high-level formulation of weak dependence. This section derives primitive conditions for the network formation model in § 2, so throughout this section we work under its setup. We begin in § 4.1 by introducing key definitions used in § 4.2 to explain two sources of cross-sectional dependence induced by the model. These motivate the weak dependence conditions stated in § 4.3 and § 4.4. In § 4.5 we present the main result, that these conditions imply exponential stabilization. We outline the method of proof in § 4.6.
4.1 Strategic Neighborhood
Recall the definition of the joint surplus from (1), and let
$D_{ij}=\bm{1}\big{\{}\sup_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0\big{\}}\\ \times\bm{1}\big{\{}\inf_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})\leq 0\big{\}}.$ (15)
This is an indicator for whether the potential link $A_{ij}$ is non-robust. If $\inf_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0$, then $A_{ij}=1$, and the link is robust in that the joint surplus is positive regardless of what other links are formed. This is because $\bm{A}$ enters $V(\cdot)$ only through $S_{ij}$. Likewise, if $\sup_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})\leq 0$, then $A_{ij}=0$, and the link is robustly absent in that the joint surplus is negative regardless of what other links are formed. In either case, $D_{ij}=0$. If instead $D_{ij}=1$, then $A_{ij}$ may be 0 or 1, and the potential link is non-robust in that the sign of the joint surplus is responsive to links formed by others.
Let $\bm{D}$ be the network of non-robustness indicators with $ij$th entry $D_{ij}$. Let $C_{i}$ denote $i$’s component in $\bm{D}$, recalling from § 1 that a component is a connected subnetwork that is disconnected from the rest of the network. Let $\bm{\Pi}$ be the network of robust link indicators with $ij$th entry
$\Pi_{ij}=\bm{1}\big{\{}\inf_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0\big{\}}.$
Recall that $\mathcal{N}_{\bm{\Pi}}(i,1)$ denotes $i$’s 1-neighborhood in $\bm{\Pi}$, which includes $i$ itself. A crucial concept for what follows is a node’s strategic neighborhood, given by
$C_{i}^{+}=\bigcup\left\{\mathcal{N}_{\bm{\Pi}}(j,1)\colon j\in C_{i}\right\}.$ (16)
This adds to $C_{i}$ the set of all nodes that possess a robust link to some member of $C_{i}$.
Example 7.
Consider Example 1, and suppose $\theta_{2}>0$. Then $A_{ij}$ is robust if $\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{ij}>0$ and robustly absent if $\theta_{1}+\theta_{2}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{ij}\leq 0$. The non-robust indicator is
$D_{ij}=\bm{1}\big{\{}-\theta_{2}<\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{12}\leq 0\big{\}}.$
As the strength of strategic interactions $\theta_{2}$ increases, so does the right-hand side, and hence, the likelihood of non-robustness. In the case of no strategic interactions ($\theta_{2}=0$), there are no non-robust potential links.
We can compute $C_{i}$ as follows. Initialize $C_{i}=\{i\}$, add $i$’s neighbors in the network $\bm{D}$ to the set, and then iteratively add neighbors of neighbors in the manner of a breadth-first search until there are no new nodes to add. To compute $C_{i}^{+}$, we set $\Pi_{ij}=\bm{1}\{\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert)^{2}+\zeta_{ij}>0\}$ for all $i\neq j$ and add to $C_{i}$ all nodes that are neighbors under $\bm{\Pi}$ of some member of $C_{i}$.
The previous example illustrates how $D_{ij}$ is increasing in the strength of strategic interactions. Stronger interactions then imply that $C_{i}^{+}$ is a larger set for any given realization of the primitives. This suggests that, when strategic neighborhoods are likely small in size, we expect weaker strategic interactions and hence weaker cross-sectional dependence. We elaborate on this point in the next subsection.
Figure 1: A network with three strategic neighborhoods.
Example 8.
Let the primitives $(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ be given, which fully determine the realizations of non-robustness indicators $\bm{D}$ and robust link indicators $\bm{\Pi}$ defined in Example 7. Suppose their realizations are depicted in Figure 1, where thin lines represent non-robust potential links (those of $\bm{D}$), solid lines represent robust links (those of $\bm{\Pi}$), and the absence of a line between two nodes represents a robustly absent link. The observed network $\bm{A}$ is a subnetwork of the depicted network because if $D_{ij}=1$, then $A_{ij}$ is either 1 or 0, while if $D_{ij}=0$, then the link between $i,j$ is either robustly absent, in which case $A_{ij}=0$, or robust ($\Pi_{ij}=1$), in which case $A_{ij}=1$. The components of $\bm{D}$ are the three subnetworks obtained by removing the solid lines: $\{0,1,2\}$, $\{3,4,5,6\}$, and $\{7,8,9\}$. The strategic neighborhoods are the three circled subnetworks $\{0,1,2,3\}$, $\{2,\ldots,7\}$, and $\{6,7,8,9\}$, obtained by adding to each component the set of nodes with solid lines to any member of the component. Notice that strategic neighborhoods may have nodes in common, whereas components necessarily partition $\mathcal{N}_{n}$.
4.2 Dependence Structure
Models of strategic network formation induce two forms of cross-sectional dependence, one due to link interdependencies directly induced by strategic interactions, and the other due to equilibrium selection. To illustrate, consider a model with $n=9$ nodes and realized structural primitives such that Figure 1 depicts the resulting robust and non-robust potential links. Consider the impact on the realized network $\bm{A}$ of a hypothetical intervention that perturbs the value of $\zeta_{56}$.
4.2.1 Best-Response Chains
The perturbation changes the joint surplus $V_{56}$ to some new value $V_{56}^{(1)}$, both evaluated under $\bm{A}$. If the signs of these values differ, then $A_{56}$ is no longer pairwise stable, so suppose in response the pair updates their potential link $A_{56}$ to $A_{56}^{(1)}=\bm{1}\{V_{56}^{(1)}>0\}$, resulting in the network $\bm{A}_{1}$. This update may affect the joint surplus of other potential links, for instance $V_{67}$ via $S_{67}$. However, because $A_{67}$ is a robust link, $V_{67}$ is always positive, whether evaluated under $\bm{A}$ or $\bm{A}_{1}$, so there is no subsequent update to $A_{67}$. On the other hand, $A_{45}$ is non-robust ($D_{45}=1$), so the update to $A_{56}$ may change the value of $V_{45}$ evaluated under $\bm{A}$ to a new value $V_{45}^{(2)}$ evaluated under $\bm{A}_{1}$ that has a different sign. Suppose in response the pair updates $A_{45}$ to $A_{45}^{(2)}=\bm{1}\{V_{45}^{(2)}>0\}$, resulting in a new network $\bm{A}_{2}$. This in turn affects the joint surplus of any other node pair containing nodes 4 and 5. The only such pairs with non-robust potential links are $(3,4)$ and $(5,6)$ since $D_{34}=D_{56}=1$, so only those potential links may update in response to the update to $A_{45}$.
Suppose we iterate these best-response dynamics indefinitely. At no point do potential links other than those of pairs $(3,4)$, $(4,5)$, and $(5,6)$ update during the process. This is because they are the only pairs connected to nodes 5 or 6, the target of the initial perturbation, through a chain of non-robust potential links, or more formally a path in $\bm{D}$. These paths constitute the furthest extent that best-response chains extending from nodes 5 and 6 can travel. By definition, the component of $\bm{D}$ containing nodes 5 and 6 contains all such paths.
The previous example pertains to interventions that perturb non-robust potential links, specifically $A_{56}$. Interventions that perturb robust links, such as $A_{67}$, trigger best-response dynamics in all strategic neighborhoods containing the link, which would be the two right-most components in Figure 1.
These examples demonstrate that the sizes of components $C_{i}$ and strategic neighborhoods $C_{i}^{+}$ (since they contain the components) are indicative of the degree of cross-sectional dependence. The key idea is that if strategic neighborhoods are relatively small, then best-response chains cannot extend too far from the initial perturbation, corresponding to weaker dependence. To control the length of best-response chains, in § 4.3, we state a “subcriticality” condition that ensures that component sizes are asymptotically bounded. From the discussion following Example 7, this should hold if strategic interactions are sufficiently weak.
To derive the condition, we employ a well-known technique used in random graph theory for bounding the size of a component, which is to construct a branching process (see § SA.2.1 for a formal definition) whose size stochastically dominates that of the component (e.g. Bollobás and Riordan, 2012). The basic idea is to explore each $C_{i}$ via breadth-first search by starting at $i$, branching to its neighbors, neighbors of its neighbors, and so on. This is akin to growing a branching process, a model of population growth in which individuals in a given generation independently produce a random number of offspring, which corresponds to a node’s neighborhood size. Subcriticality ensures that the average number of offspring is less than one, in which case the size of the process does not diverge and $\lvert C_{i}\rvert$ is asymptotically bounded.
While this is enough to establish an LLN (Leung, 2019b), a CLT additionally requires the distribution of $\lvert C_{i}\rvert$ to have exponential tails. We utilize a tail bound for subcritical branching processes to obtain the desired result; see Lemma SA.2.3 which is proven by Leung (2020) using an argument due to Turova (2012). This is a key ingredient for verifying Assumption 5.
4.2.2 Coordination
The second source of dependence is due to equilibrium selection. Let us shut down the first source of dependence by supposing the perturbation to $\zeta_{56}$ is small enough not to change the sign of $V_{56}$, so there is no change to the network at any point in the best-response dynamics outlined above. However, if equilibrium selection is governed by a mechanism different from best-response dynamics, the perturbation may still substantially affect the network structure.
Suppose $\lvert\mathcal{E}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})\rvert=2$. Since the perturbation does not change $V_{56}$, the set of equilibria is identical before and after the perturbation. However, one can construct a selection mechanism $\lambda_{n}(\cdot)$ that outputs one equilibrium under the structural primitives before the perturbation but outputs the other after the perturbation. Hence, the perturbation can alter potential links involving nodes external to $C_{5}$ even though any best-response chain must be limited to this component. This illustrates how, under unrestricted equilibrium selection, all nodes may coordinate on the same “signal” ($\zeta_{56}$), resulting in strongly dependent potential links whose realizations all depend on this random variable.
In § 4.4, we state a condition that ensures coordination is “decentralized” in that strategic neighborhoods “separately select” their pairwise stable subnetworks. A consequence is that any perturbation to a node pair’s types or random-utility shocks only affects the selection of the equilibrium subnetwork on strategic neighborhoods involving the pair. For example, in Figure 1, it will be the case that a perturbation to $\zeta_{56}$ only changes the equilibrium subnetwork on $\{2,\ldots,7\}$ since no other neighborhood contains $\{5,6\}$. Combined with the subcriticality condition that ensures strategic neighborhood sizes are asymptotically bounded, this implies that the perturbation can only shift equilibrium selection within a bounded subset of nodes, ensuring weak cross-sectional dependence. This will allow us to construct a radius of stabilization that is asymptotically bounded.
4.3 Strength of Interactions
We next state a condition that controls the sizes of strategic neighborhoods, and hence length of best-response chains described in § 4.2.1, by restricting the magnitude of strategic interactions. We measure strategic interaction strength by
$p_{r_{n}}(X_{i},Z_{i},X_{j},Z_{j})=\,\,$ ${\bf P}\big{(}\sup_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0\mid X_{i},Z_{i},X_{j},Z_{j}\big{)}$
-\,\, ${\bf P}\big{(}\inf_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0\mid X_{i},Z_{i},X_{j},Z_{j}\big{)}.$ (17)
This is the effect on link formation of changing $S_{ij}$ from its “lowest” to its “highest” possible value, conditional on types. In other words, it is the maximal change in linking probability induced by the strategic component of $V(\cdot)$.
Let $\Phi_{z}(\cdot\,|\,x)$ be the conditional distribution of $Z_{i}$ given $X_{i}=x$. Recall that $f$ is the density of $X_{1}$, $\mathcal{T}=\text{supp}(X_{1},Z_{1})$, and $d_{z}$ is the dimension of $Z_{i}$. For any $h\colon\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}$, define the mixed norm
$\lVert h\rVert_{\mathbf{m}}=\sup_{x\in\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d_{z}}}h(x,z)^{2}\,\text{d}\Phi^{*}(z)\right)^{1/2},$
where $\Phi^{*}$ is a measure defined in the next assumption. Finally, for $\bar{f}=\sup_{x\in\mathbb{R}^{d}}f(x)$ and $\kappa$ in (5), let
$h_{r_{n}}(x,z)=n\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d_{z}}}p_{r_{n}}(x,z,x^{\prime},z^{\prime})^{2}\,\text{d}\Phi_{z}(z^{\prime}\mid x^{\prime})\right)^{1/2}f(x^{\prime})\,\text{d}x^{\prime}\quad\text{and}\\ h^{*}(x,z)=\kappa\bar{f}\,\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d_{z}}}p_{1}(x,z;x^{\prime},z^{\prime})^{2}\,\text{d}\Phi^{*}(z^{\prime})\right)^{1/2}\,\text{d}x^{\prime}.$
Assumption 7 (Subcriticality).
(a) There exists a measure $\Phi^{*}$ on $\mathbb{R}^{d_{z}}$ such that for all $(x,z)\in\mathcal{T}$ and $n\in\mathbb{N}$,
$n\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}p_{r_{n}}(x,z;x^{\prime},z^{\prime})\,\text{d}\Phi_{z}(z^{\prime}\mid x^{\prime})f(x^{\prime})\,\text{d}x^{\prime}\leq\kappa\bar{f}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}p_{1}(x,z;x^{\prime},z^{\prime})\,\text{d}\Phi^{*}(z^{\prime})\,\text{d}x^{\prime},$
and $h_{r_{n}}(x,z)\leq h^{*}(x,z)$. (b) $\lVert h^{*}\rVert_{\mathbf{m}}<1$.
Part (a) is a regularity condition corresponding to Assumption 2 of Leung (2020) that replaces $\Phi_{z}(\cdot\mid x^{\prime})$ with a “dominating” measure $\Phi^{*}$. The substantive requirement is (b), which is a slightly stronger version of Assumption 6 of Leung (2019b). Its formulation is as primitive as possible at this level of generality, but we may obtain further insight by specializing to a particular joint surplus function. The next example demonstrates how the condition restricts the magnitude of strategic interactions.
Example 9.
Consider Example 1, and suppose $X_{i}\sim\mathcal{U}([0,1]^{2})$ and $\zeta_{ij}\sim\mathcal{N}(0,1)$. By a change of variables $x^{\prime}\mapsto x+r_{n}(x^{\prime}-x)$ as in (6),
$n\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}p_{r_{n}}(x,z;x^{\prime},z^{\prime})\,\text{d}\Phi_{z}(z^{\prime}\mid x^{\prime})f(x^{\prime})\,\text{d}x^{\prime}\\ =n\int_{\mathbb{R}^{d}}{\bf P}\big{(}-\theta_{2}<\theta_{1}-\theta_{3}(r_{n}^{-1}\lVert x-x^{\prime}\rVert)^{2}+\zeta_{12}\leq 0\big{)}\bm{1}\{x^{\prime}\in[0,1]^{2}\}\,\text{d}x^{\prime}\\ \leq\kappa\int_{\mathbb{R}^{d}}{\bf P}\big{(}-\theta_{2}<\theta_{1}-\theta_{3}\lVert x-x^{\prime}\rVert^{2}+\zeta_{12}\leq 0\big{)}\,\text{d}x^{\prime}.$
Notice the second line equals $h_{r_{n}}(x,z)$ and the third line $h^{*}(x,z)$. Letting $\Phi(\cdot)$ denote the normal CDF,
$h^{*}(x,z)$ $=\kappa\int_{\mathbb{R}^{2}}\left[\Phi(\theta_{1}+\theta_{2}-\theta_{3}\lVert x\rVert^{2})-\Phi(\theta_{1}-\theta_{3}\lVert x\rVert^{2})\right]\,\text{d}x$
$=2\kappa\pi\int_{0}^{\infty}w\left[\Phi(\theta_{1}+\theta_{2}-\theta_{3}w^{2})-\Phi(\theta_{1}-\theta_{3}w^{2})\right]\,\text{d}w$
by a change of variables to polar coordinates, where $\pi$ is the transcendental constant. Let $q(x)=x\Phi(x)+\phi(x)$, where $\phi(\cdot)$ is the normal PDF. The integral has a closed-form expression
$h^{*}(x^{\prime},z^{\prime})=\frac{\kappa\pi}{\theta_{3}}\big{(}q(\theta_{1}+\theta_{2})-q(\theta_{1})\big{)}.$
Assumption 7 requires this to be less than one, which holds if $\theta_{2}$ is sufficiently small in magnitude. This is transparently a restriction on the strength of strategic interactions.
The example illustrates how $\lVert h^{*}\rVert_{\mathbf{m}}<1$ constitutes the network formation analog of well-known weak dependence conditions for linear spatial or temporal autoregressive models, which require the magnitude of the autoregressive coefficient to be bounded below one. For instance, in the linear-in-means model of peer effects, it is assumed that the endogenous peer effect satisfies $\lvert\beta\rvert<1$ (Bramoullé et al., 2009). For nonlinear models such as ours, the analogous condition is necessarily more complicated to state. Equation (9) of de Jong and Woutersen (2011) states the condition for dynamic binary choice time series models, which shares some visual similarities with our assumption.
[…]
4.4 Coordination
The second weak dependence condition restricts the selection mechanism, ruling out coordination of the type described in § 4.2.2. It requires selection to be “decentralized” in that each strategic neighborhood $C^{+}$ selects its pairwise stable subnetwork based on the types and random-utility shocks of nodes in $C^{+}$ alone. In this sense, there is no coordination across disjoint strategic neighborhoods.
Before stating the assumption, we need to clarify why such a restriction is possible. Initially, it may appear incoherent because the pairwise stability of a subnetwork on $H\subseteq\mathcal{N}_{n}$ can depend on the state of the network outside of $H$. If nodes $\H\mathcal{N}_{n}\backslash$ H are removed from the model, we would expect nodes in $H$ to adjust their links in best response. This suggests that coordination across subsets of nodes is generally unavoidable.
However, if $H$ is specifically a strategic neighborhood, then under Assumption 2, its subnetwork is in fact pairwise stable regardless of the state of the ambient network. Consider Figure 1 and a counterfactual intervention that removes all nodes from the network outside of $C_{5}^{+}=\{2,\ldots,7\}$. The intervention has no impact on the pairwise stability of $A_{23}$ and $A_{67}$ since these are robust links. It also has no impact on the pairwise stability of non-robust potential links between nodes in $C_{5}^{+}$, for example $A_{45}$ since neither nodes 4 or 5 are linked to $\mathcal{N}_{n}\backslash C_{5}^{+}$ in $\bm{A}$ and strategic interactions are local by Assumption 2. That is, the non-robust potential links in $C_{5}^{+}$ are separated from $\mathcal{N}_{n}\backslash C_{5}^{+}$ by a “buffer” of robust links, so the intervention does not trigger a chain of best-responses that affects $\bm{A}_{C_{5}^{+}}$. This illustrates the following result.
Proposition 1 (Leung (2019b), Proposition 1).
Under Assumption 2, for any $r>0$, $n\in\mathbb{N}$, and strategic neighborhood $C^{+}$,
$\mathcal{E}(r^{-1}\bm{T}_{C^{+}},\bm{\zeta}_{C^{+}})=\{\bm{A}_{C^{+}}:\bm{A}\in\mathcal{E}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})\}.$
The left-hand side is the set of pairwise stable networks under the counterfactual model in which the set of nodes is $C^{+}$, rather than $\mathcal{N}_{n}$. The right-hand side takes the set of pairwise stable networks on $\mathcal{N}_{n}$ and restricts them to $C^{+}$. The proposition asserts the two sets are equivalent, which says that the pairwise stability of subnetworks on $C^{+}$ only depends on the structural primitives of nodes in $C^{+}$. We emphasize that this property is unique to strategic neighborhoods and is not true for arbitrary subsets of nodes. Leung (2020) exploits Proposition 1 to devise an algorithm that computes $\mathcal{E}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})$ in polynomial time under Assumption 7. We utilize it to establish a CLT.
Recall that our objective is to impose the assumption that strategic neighborhoods “separately select” their own pairwise stable subnetworks. Proposition 1 ensures that it is coherent to refer to the pairwise stability of a subnetwork on a strategic neighborhood in isolation from the rest of the network. However, a second concern is that strategic neighborhoods do not necessarily partition $\mathcal{N}_{n}$, so such an assumption may still appear incoherent. For instance, in Figure 1, the left and middle strategic neighborhoods share nodes 2 and 3 in common, so it is not clear how the two neighborhoods can separately select equilibria.
Recall that components of $\bm{D}$ do partition $\mathcal{N}_{n}$, while strategic neighborhoods are obtained by adding nodes that are robustly linked to components. Then the link between any node pair that lies in multiple strategic neighborhoods must necessarily be robust and therefore have the same realization under any pairwise stable equilibrium. That is, any pairwise stable subnetwork on the left strategic neighborhood in Figure 1 sets $A_{23}=1$, as does any pairwise stable subnetwork on the middle strategic neighborhood. Consequently, it is a logically coherent operation to select an equilibrium subnetwork for each strategic neighborhood and then take a “Cartesian product” to obtain the overall network $\bm{A}$ (see §SA.4.2 of Leung, 2020, for a detailed elaboration of this idea).
We are now prepared to state the assumption. Let $\lambda_{n}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})\big{|}_{H}$ be the restriction of the range of $\lambda_{n}(\cdot)$ to subnetworks on $H$. For example, under Assumption 3, $\lambda_{n}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})=\bm{A}$, so $\lambda_{n}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})\big{|}_{H}=\bm{A}_{H}$.
Assumption 8 (Decentralized Selection).
For any $r>0$, $n\in\mathbb{N}$, and strategic neighborhood $C^{+}$ constructed under the structural primitives $(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})$, we have $\lambda_{n}(r^{-1}\bm{T}_{n},\bm{\zeta}_{n})\big{|}_{C^{+}}=\lambda_{\lvert C^{+}\rvert}(r^{-1}\bm{T}_{C^{+}},\bm{\zeta}_{C^{+}})$.
[…]
Sheng (2020) does not require this assumption because she considers a setting with many small independent networks, which means equilibrium selection is necessarily independent across network observations. Assumption 8 is the single-network analog of this requirement, requiring selection to operate separately across latent strategic neighborhoods.
The assumption rules out selection mechanisms in which all nodes coordinate through a common signal, such as a single node’s type. Coordination is only allowed to occur within strategic neighborhoods. In the special case where there exists a unique equilibrium on $\mathcal{N}_{n}$, for instance if there are no strategic interactions, the assumption holds trivially. More generally, the condition is satisfied by variants of myopic best-response dynamics, which are widely used in the theoretical and econometric literature on dynamic network formation (e.g. Jackson, 2010; Mele, 2017).
Example 10 (Best-Response Dynamics).
An example of myopic best-response dynamics is the following. Arbitrarily order all node pairs and begin at an arbitrary network $\bm{A}_{0}$. At step $t$, update the previous network $\bm{A}_{t-1}$ by setting the $ij$th component to $\bm{1}\{V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,S_{ij}^{t-1},Z_{i},Z_{j},\zeta_{ij})>0\}$, where $S_{ij}^{t-1}=S_{n}(i,j,r_{n}^{-1}\bm{T}_{n},\bm{A}_{t-1})$, the network statistics evaluated at the prior network. Repeat for all pairs of nodes to obtain $\bm{A}_{t}$. Repeat this process until convergence to a network $\bm{A}\in\mathcal{E}(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})$. [^7] The usual method for proving the existence of a pairwise stable network is to establish non-existence of “closed cycles,” which implies that myopic best-response dynamics always converge to an equilibrium. See for example Proposition 2.1 of Sheng (2020). This constitutes a selection mechanism $\lambda_{n}(\cdot)$ since it maps structural primitives to an equilibrium network. To ensure that $\lambda_{n}(\cdot)$ is permutation-equivariant (Assumption 3), we may suppose that potential links in $\bm{A}_{0}$ and the ordering of node pairs are only functions of the types of the nodes in the pair. [^8] More generally, the initial network and ordering of node pairs may be determined by any permutation-equivariant functions of $(r_{n}^{-1}\bm{T}_{n},\bm{\zeta}_{n})$.
To understand how this satisfies Assumption 8, first consider the ideal scenario in which the data consists of two independent network observations formed by these dynamics. The equilibrium on each network may be generated in two equivalent ways. First, the dynamics may be run separately on each network. Second, the two networks may be concatenated into one, defining the joint surplus between pairs of nodes in different networks as -$\infty$, and the dynamics may be run on the entire entity. These produce the same output given the initial network and node pair ordering.
Now suppose the data consists of a single network comprised of two strategic neighborhoods. If the neighborhoods do not share a pair of nodes in common, then this is the same situation as the two-network case. If they do, then both ways of running the dynamics still produce the same result because any pair of nodes shared by both neighborhoods must form a robust link. This reasoning immediately extends to an arbitrary number of strategic neighborhoods, so Assumption 8 holds.
4.5 Main Results
The last assumption we require is a regularity condition.
Assumption 9 (Regularity).
Either $p_{r}(X_{1},Z_{1};X_{2},Z_{2})=0$ a.s. for any $r>0$ in a neighborhood of zero, or $\inf\{\liminf_{n\rightarrow\infty}n\,{\bf E}[p_{r_{n}}(x,z;X_{2},Z_{2})]\colon(x,z)\in\mathcal{T}\}>0$.
The case $p_{r}(X_{1},Z_{1};X_{2},Z_{2})=0$ corresponds to a model without strategic interactions, which is only mentioned for completeness. In the more interesting case, the assumption essentially requires that strategic interactions are sufficiently nontrivial for all nodes in that $p_{r_{n}}(X_{1},Z_{1};X_{2},Z_{2})$ is at least order $n^{-1}$. This is a mild requirement that is typically satisfied because the two probabilities in (17) are upper and lower bounds on the probability of link formation, and the upper bound is order $n^{-1}$ under sparsity by (6).
Example 11.
Consider Example 9. Following the derivation there,
$n\,{\bf E}[p_{r}(x,z;X_{2},Z_{2})]=2nr^{2}\pi\int_{0}^{\infty}w\left[\Phi(\theta_{1}+\theta_{2}-\theta_{3}w^{2})-\Phi(\theta_{1}-\theta_{3}w^{2})\right]\,\text{d}w.$
If $\theta_{2}=0$, this corresponds to the case $p_{r}(X_{1},Z_{1};X_{2},Z_{2})=0$ a.s. Otherwise, Assumption 9 holds if $\theta_{2}\neq 0$ since $nr_{n}^{2}=\kappa>0$ by (5).
Theorem 2.
Assumptions 14 and 79 imply Assumption 5.
Proof. See § SA.2. We sketch the proof below.
From Theorems 1 and 2 we immediately obtain the following CLT for strategic network formation.
Corollary 1.
Under Assumptions 14 and 69, $\sup_{n}\lVert\bm{\Sigma}_{n}\rVert_{\infty}<\infty$, and if additionally
$\liminf_{n\rightarrow\infty}\lambda_{\text{min}}(\bm{\Sigma}_{n})>0$, then (12) holds.
At this level of generality, these conditions are close to as primitive as possible, but with additional structure we can derive lower-level conditions.
Corollary 2.
Consider the network formation model in Example 9, and assume Assumption 6(b) and the following weak dependence conditions hold.
  1. (a)
    (Strength of interactions) For $q(x)$ defined in the example,
    $\frac{\kappa\pi}{\theta_{3}}\big{(}q(\theta_{1}+\theta_{2})-q(\theta_{1})\big{)}<1.$
  2. (b)
    (Decentralized selection) The equilibrium is selected via myopic best-response dynamics as in Example 10.
If $\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$ is a vector of connected subnetwork counts (Example 5), then (12) holds.
Proof. We verify the conditions of Corollary 1. Assumption 1 holds because
$\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\delta,0))=\bar{\Phi}_{\zeta}(\theta_{3}\delta^{2}-\theta_{1}-\theta_{2}),$
which decays to zero exponentially with $\delta$ since $\bar{\Phi}_{\zeta}$ is the complementary CDF of the standard normal distribution. Assumption 2 holds by choice of $S_{ij}$. Assumptions 3 and 8 hold by (b). Assumption 4 holds because we consider connected subnetwork counts (see Example 6). Assumption 6(a) follows from Proposition A.1. Assumption 7 follows from (a) (see Example 9). Finally, Assumption 9 holds, as shown in Example 11.
4.6 Method of Proof
We present the method of proof for Theorem 2 for the case of the average degree, whose node statistic is $\psi_{i}(\mathcal{N}_{n})=\sum_{j=1}^{n}A_{ij}$. The approach can be applied to other network models with strategic interactions. [^9] A previous version of the paper presents applications to dynamic network formation (Leung and Moon, 2019), and Leung (2019a) applies the methodology to games on networks. We construct an upper bound on the radius of stabilization that has a distribution with sufficiently thin tails.
Step 1. Recalling (9), we construct a set of nodes $J_{i}\subseteq\mathcal{N}_{n}$ positioned near $i$ such that
$\psi_{i}(\mathcal{N}_{n})=\psi_{i}(J_{i}).$ (18)
That is, $i$’s node statistic is invariant to the counterfactual removal of nodes outside of $J_{i}$. The challenge is to find a set that is relatively small so that the distribution of the set’s size has thin tails. Supposing such a set could be found, the radius of stabilization $\bm{R}_{i}(n,r_{n})$ (Definition 1) would be upper bounded by the radius of the smallest ball centered at $r_{n}^{-1}X_{i}$ containing the positions of $J_{i}$ (plus one):
$\bm{R}_{i}(n,r_{n})\leq\tilde{\bm{R}}_{i}(n,r_{n})\equiv\max_{j\in J_{i}}r_{n}^{-1}\lVert X_{i}-X_{j}\rVert+1.$ (19)
Since $\psi_{i}(\mathcal{N}_{n})=\sum_{j=1}^{n}A_{ij}$ is simply the 1-neighborhood size, an initial guess for $J_{i}$ might be $i$’s 1-neighborhood. However, this does not generally satisfy (18). To see why, consider node $i=6$ in Figure 1 whose 1-neighborhood is contained in $\{4,5,7\}$. Since $D_{56}=1$, the potential link $A_{56}$ is non-robust and therefore may differ under the counterfactual that removes node 3 from the model, even though 3 is not $i$’s neighbor.
[…]
Step 2. Next we show that $\lvert J_{i}\rvert$ has a distribution with exponential tails. First, § SA.2.1 establishes that $\lvert J_{i}\rvert$ is stochastically dominated by the size of a certain branching process when $J_{i}$ is constructed from components and $K$-neighborhoods of dyadic networks. Lemma SA.2.6 shows that the size of the branching process has a distribution with exponential tails using Assumptions 7 and 9.
In the case of average degree, our choice of $J_{i}=C_{i}^{+}$ is the union of a component of $\bm{D}$ and 1-neighborhoods in $\bm{\Pi}$, as defined in § 4, which are dyadic networks. For more general $K$-neighborhood node statistics satisfying Assumption 4, the construction of $J_{i}$ is more complicated but still involves similar objects. Lemma SA.2.1 in § SA.2.2 shows how to construct $J_{i}$ for a general node statistic satisfying Assumption 4. The lemmas in § SA.2.2 can then be applied to obtain the desired tail bounds.
Step 3. We translate the tail bound for $\lvert J_{i}\rvert$ into one for $\tilde{\bm{R}}_{i}(n,r_{n})$. Intuitively, if $\lvert J_{i}\rvert$ is small, then so will be $\tilde{\bm{R}}_{i}(n,r_{n})$ since nodes are homophilous in positions (Assumption 1), so each $j\in J_{i}$ will typically be close to $i$ in terms of distance $r_{n}^{-1}\lVert X_{i}-X_{j}\rVert$. Lemma SA.2.8 provides the formal argument.
5 Applications to Inference
We discuss two inference procedures applicable to network data generated by the model in § 2. Our results provide the first formal justification for their use in the subsequent applications.
Since our objective is to establish a CLT, thus far we have made no distinction between what aspects of the model are known or observed by the econometrician. In what follows, the only requirement is that the econometrician must be able to compute the relevant network moments, but it is application-specific which structural primitives need to be observed for this to be possible.
In the network statistics applications, only $\bm{A}$ needs to be observed. In structural applications, $V(\cdot)$ is typically known up to a vector of parameters, $\bm{A}$ is observed, and $\zeta_{ij}$ is unobserved, but any subvector of $(X_{i},Z_{i})$ could potentially be unobserved. In particular, the procedures can be implemented when positions are unobserved, as in the literature on latent-space models. Typically the distribution of the unobserved component of types, conditional on observables, is assumed known up to a vector of parameters.
5.1 Network Statistics
Define $\mu_{0}={\bf E}[\psi_{1}(\mathcal{N}_{n})]$ and $\hat{\mu}=n^{-1}\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$. Consider testing the null hypothesis
$H_{0}\colon\mu_{0}=\mu.$
This is relevant for the reporting of stylized facts in the networks literature. Such facts are obtained by computing various statistics from $\bm{A}$ but are seldom accompanied by formal uncertainty quantification due to a lack of available methods. Leung (2022) discusses two particular examples: testing for nontrivial clustering (his §3.3) and testing for a power law degree distribution (his §3.4). We next discuss two generic tests justified by our CLT. Confidence regions for $\mu_{0}$ can be obtained by test inversion.
[…]
$T_{U}(\mu;\pi)=\frac{1}{\sqrt{d_{\psi}R_{n}}}\sum_{r=1}^{R_{n}}(\psi_{\pi_{r}(1)}(\mathcal{N}_{n})-\mu)^{\prime}\hat{\bm{V}}^{-1}(\psi_{\pi_{r}(2)}(\mathcal{N}_{n})-\mu),$
recalling that $d_{\psi}$ is the dimension of the range of $\psi_{n}(\cdot)$. Let $z_{1-\alpha}$ be the $1-\alpha$ quantile of the standard normal distribution. The test rejects if and only if
$T_{U}(\mu;\pi)>z_{1-\alpha}.$ (20)
Validity of this test hinges on the high-level weak dependence condition that $\hat{\mu}$ is $\sqrt{n}$-consistent. Leung (2019b) provides an LLN for $\hat{\mu}$ but not a rate of convergence. Our paper is the first to provide primitive conditions for $\sqrt{n}$-consistency in the context of strategic network formation.
For intuition on the importance of $\sqrt{n}$-consistency for the test’s validity, consider a simpler test statistic studied by Leung (2022):
$\tilde{T}_{M}(\mu;\pi)=\frac{1}{\sqrt{R_{n}^{M}}}\sum_{r=1}^{R_{n}^{M}}\hat{\bm{V}}^{-1/2}\big{(}\psi_{\pi_{r}(1)}(\mathcal{N}_{n})-\mu\big{)}$
for $R_{n}^{M}=\sqrt{n}$ rounded to the nearest integer. To understand its asymptotic behavior, we add and subtract its mean conditional on the data:
$\frac{1}{\sqrt{R_{n}^{M}}}\sum_{r=1}^{R_{n}^{M}}\hat{\bm{V}}^{-1/2}\big{(}\psi_{\pi_{r}(1)}(\mathcal{N}_{n})-{\bf E}[\psi_{\pi_{r}(1)}(\mathcal{N}_{n})\mid\{\psi_{i}(\mathcal{N}_{n})\}_{i=1}^{n}]\big{)}\\ +\frac{1}{\sqrt{R_{n}^{M}}}\sum_{r=1}^{R_{n}^{M}}\hat{\bm{V}}^{-1/2}\big{(}{\bf E}[\psi_{\pi_{r}(1)}(\mathcal{N}_{n})\mid\{\psi_{i}(\mathcal{N}_{n})\}_{i=1}^{n}]-\mu\big{)}.$
Conditional on the data $\{\psi_{i}(\mathcal{N}_{n})\}_{i=1}^{n}$, the permutations $\pi_{r}$ are independent, so the first term is an average of $R_{n}^{M}$ conditionally independent observations and can be shown to be asymptotically $\mathcal{N}(\bm{0},\bm{I})$ under $H_{0}$. The second is a bias term that can be shown to equal $(R_{n}^{M}/n)^{1/2}\hat{\bm{V}}^{-1/2}\sqrt{n}(\hat{\mu}-\mu)$. Hence, under $H_{0}$ and $\sqrt{n}$-consistency, the bias is order $(R_{n}^{M}/n)^{1/2}=o(1)$.
Multiple large networks. A drawback of the previous procedure is that it is inefficient, having a rate of convergence slower than $\sqrt{n}$ since $R_{n}^{M}=o(n)$. If the sample consists of sufficiently many independent large networks, then more powerful methods are available from the cluster-robust inference literature. Consider a sequence of $L$ independent networks indexed by $n$, where $L$ is fixed with respect to $n$ and each network $\ell=1,\dots,L$ has size $n_{\ell}$ satisfying $n_{\ell}/n\rightarrow c_{\ell}\in(0,\infty)$ as $n\rightarrow\infty$. Let $\hat{\mu}^{\ell}=n_{\ell}^{-1}\sum_{i=1}^{n_{\ell}}\psi_{i}(\mathcal{N}_{n_{\ell}})$, the network moment computed on network $\ell$. Assume there exists a universal population moment $\mu_{0}$ such that ${\bf E}[\hat{\mu}^{\ell}]=\mu_{0}+o(n^{-1/2})$ for all $\ell$. We seek to test the null $H_{0}\colon\mu_{0}=\mu$.
We consider the randomization test proposed by Canay et al. (2017). For $S_{n,\ell}=\sqrt{n_{\ell}}(\hat{\mu}^{\ell}-\mu)$ and $S_{n}=(S_{n,\ell})_{\ell=1}^{L}$, define the Wald statistic
$T(S_{n})=\left(\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}S_{n,\ell}^{\prime}\right)\left(\frac{1}{L}\sum_{\ell=1}^{L}S_{n,\ell}S_{n,\ell}^{\prime}\right)^{-1}\left(\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}S_{n,\ell}\right).$
We obtain critical values from the randomization distribution $\{T(\pi S_{n})\colon\pi\in\{-1,1\}^{L}\}$ where $\pi S_{n}=(\pi_{\ell}S_{n,\ell})_{\ell=1}^{L}$ for $\pi=(\pi_{\ell})_{\ell=1}^{L}$. Let $\alpha$ be the desired level of the test, $q=2^{L}(1-\alpha)$ rounded up to the nearest integer, and $c_{L,1-\alpha}$ be the $q$th largest value of $\{T(\pi S_{n})\colon\pi\in\{-1,1\}^{L}\}$. The test rejects if and only if
$T(S_{n})>c_{L,1-\alpha}.$
The test is asymptotically level $\alpha$ under the high-level condition that the limit distribution of the vector of network moments is asymptotically normal. We provide the first primitive sufficient conditions in the literature for strategic network formation. Intuitively, under $H_{0}$ and asymptotic normality, $\{S_{n,\ell}\}_{\ell=1}^{L}$ are independent draws from a mean-zero, approximately normal distribution. Hence, multiplying these draws by $\pm 1$ does not change the asymptotic distribution of $T(S_{n})$, which is the key justification for the validity of randomization tests.
5.2 Structural Inference
We revisit Example 6, which concerns inference on structural parameters using moment inequalities proposed by Sheng (2020). Recalling the setup there, let $\theta_{0}$ denote the true parameters of $V(\cdot)$. To test the hypothesis $H_{0}\colon\theta_{0}=\theta$, we test the moment inequality
$\mu_{0}\equiv{\bf E}[\psi_{1}(\mathcal{N}_{n})]\leq\bm{0}\quad\text{where}\quad\psi_{i_{1}}(\mathcal{N}_{n})=\sum_{i_{2}=1}^{n}\dots\sum_{i_{m}=1}^{n}(G_{i_{1},\ldots,i_{m}}-H_{i_{1},\ldots,i_{m}}(\theta_{0}))q_{i_{1},\ldots,i_{m}}(\bm{T}_{n}),$
When the data consists of a single large network, we can employ a test due to Leung (2022). Let $\mu_{k}$ be the $k$th component of $\mu_{0}$, $\psi_{ik}(\mathcal{N}_{n})$ the $k$th component of $\psi_{i}(\mathcal{N}_{n})$, and $T_{U,k}(\mu_{k};\pi)$ the U-type statistic defined in the previous subsection but computed with scalar data $\{\psi_{ik}(\mathcal{N}_{n})\}_{i=1}^{n}$. Also let $\hat{\mu}_{k}$ be the $k$th component of $\hat{\mu}$ and $\hat{\bm{V}}_{kk}$ the $k$th diagonal entry of $\hat{\bm{V}}$. Define the test statistic
$Q_{n}(\pi)=\max_{1\leq k\leq d_{\psi}}\big{\{}T_{U,k}(0;\pi)-\mu_{k}^{*}\bm{1}\{\hat{\mu}_{k}<0\}\big{\}}\quad\text{where}$
$\mu_{k}^{*}=\hat{\mu}_{k}\hat{\bm{V}}_{kk}^{-1}\frac{1}{\sqrt{d_{\psi}R_{n}}}\sum_{r=1}^{R_{n}}\left(\psi_{\pi_{r}(1),k}(\mathcal{N}_{n})+\psi_{\pi_{r}(2),k}(\mathcal{N}_{n})\right)-\sqrt{\frac{R_{n}}{d_{\psi}}}\hat{\bm{V}}_{kk}^{-1}\hat{\mu}_{k}^{2}.$
Let $\tilde{\pi}_{1},\dots,\tilde{\pi}_{L}$ be i.i.d. with the same distribution as $\pi$. Let $q=L(1-\alpha)$ rounded up to the nearest integer and $c_{L,1-\alpha}$ be $q$th largest value of $\{\max_{1\leq k\leq d_{\psi}}T_{U,k}(\hat{\mu}_{k};\tilde{\pi}_{\ell})\}_{\ell=1}^{L}$. The test rejects if and only if
$Q_{n}(\pi)>c_{L,1-\alpha}.$
Theorem 2 of Leung (2022) provides conditions under which the test is asymptotically level $\alpha$ under $H_{0}$. The main assumption that needs to be verified is $\sqrt{n}$-consistency of $\hat{\mu}$, which is a consequence of our CLT. Like the test in the previous subsection, this procedure is inefficient with a slower than $\sqrt{n}$-rate of convergence. To construct more powerful tests, we require either a consistent estimate of $\bm{\Sigma}_{n}$ or a valid resampling procedure, topics we leave to future research.
6 Simulation Study
We conduct a simulation study to assess the quality of the normal approximation and finite-sample performance of the inference procedures in § 5. We simulate data according to the model in Example 1 with $\lVert\cdot\rVert$ equal to the Euclidean norm, $\theta=(1,0.25,1)$, $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{U}([0,1]^{2})$, and $\zeta_{ij}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$. To satisfy Assumption 7, we set $\kappa=0.8\theta_{3}\pi^{-1}(q(\theta_{1}+\theta_{2})-q(\theta_{1}))^{-1}$ following the notation in Example 9, which implies $\lVert h^{*}\rVert_{\mathbf{m}}=0.8<1$. To satisfy Assumption 8, we select the equilibrium network using myopic best-response dynamics (Example 10) starting from the network $\bm{\Pi}$ defined in § 4.1. The model generates networks with fairly realistic properties; the largest component comprises about 75 percent of the network, while the average degree and clustering are respectively about 4.3 and 0.4.
We consider equality tests involving the following the network moment. Define the node statistic $\psi_{i}(\mathcal{N}_{n})=\psi_{i}^{c}-\psi_{i}^{f}$ where
$\psi_{i}^{c}=\frac{\sum_{j\neq k\neq i}A_{ij}A_{jk}A_{ik}}{\sum_{j\neq k\neq i}A_{ij}A_{ik}}\quad\text{and}\quad\psi_{i}^{f}=\frac{2}{n-1}\sum_{j\neq i}A_{ij}.$
Then $n^{-1}\sum_{i=1}^{n}\psi_{i}^{c}$ is the average clustering coefficient, a measure of triadic closure, while $n^{-1}\sum_{i=1}^{n}\psi_{i}^{f}$ is the link frequency. The motivation for the network moment $n^{-1}\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$ is as follows. Under the “null” Erdős-Rényi model, the moment is approximately zero. However, a well-known stylized fact is that real-world networks typically feature nontrivial clustering in that the statistic is far from zero (Barabási, 2015). While this fact is based only on the point estimate, the inference procedures in § 5 enable us to formulate this as a statistical test.
We simulate rejection rates for three different tests of hypotheses of the form
$H_{0}\colon{\bf E}[\psi_{1}]=c\mu\quad\text{against}\quad H_{1}\colon{\bf E}[\psi_{1}]\neq c\mu$
where $c\in(0,1]$ and $\mu$ is the true value of ${\bf E}[\psi_{1}]$, computed using 40k simulation draws. To simulate size, we set $c=1$, and to simulate power, we set $c\in\{0.8,0.9\}$. We consider the following tests at the 5-percent level. To assess the quality of the normal approximation, we use the “oracle” $t$-test with test statistic $(n^{-1}\sum_{i=1}^{n}\psi_{i}-\mu)/\sigma$, where $\sigma$ is the true standard error, computed by taking the standard deviation of the network moment across 40k simulation draws. We then consider the dependence-robust test (20) and randomization test from § 5.1. The randomization test will use data from either 5 or 8 independent networks of identical size, while the other tests only utilize data from one of the networks. We simulate size and power using 5k simulation draws.
Table 1: Simulation results
Size ($c=1$) Power ($c=0.9$) Power ($c=0.8$)
$n$ 250 500 1000 250 500 1000 250 500 1000
Oracle 0.050 0.054 0.048 0.337 0.627 0.915 0.856 0.995 1.000
DR 0.069 0.057 0.058 0.176 0.263 0.410 0.530 0.823 0.981
Rand (8) 0.047 0.045 0.044 0.947 0.999 1.000 1.000 1.000 1.000
Rand (5) 0.006 0.007 0.005 0.104 0.137 0.136 0.142 0.149 0.138
$\mu$ 0.393 0.409 0.416 0.393 0.409 0.416 0.393 0.409 0.416
$\sigma$ 0.026 0.018 0.013 0.026 0.018 0.013 0.026 0.018 0.013
  • 5000 simulations. DR corresponds to the dependence-robust test and Rand ($x$) to the randomization test using data from $x$ independent networks, all of size $n$. Oracle corresponds to the $t$-test using the true standard error $\sigma$.
Table 1 shows the results. From the first three columns, we see that the oracle $t$-test controls size well across all sample sizes, illustrating the quality of the normal approximation. The randomization test (the Rand rows in the table) controls size well across all $n$, while the dependence-robust test (the DR row in the table) exhibits some over-rejection, particularly in smaller samples. Section 6 of Leung (2022) provides additional simulation results for the dependence-robust test under a variety of data-generating processes, including a network formation model similar to the present design. His results show that the test has good properties for different choices of the tuning parameter $R_{n}$.
The last six columns of Table 1 compare power. The differences between the oracle and dependence-robust tests are due to the latter’s slower rate of convergence. A comparison of the power of the randomization and dependence-robust tests is not apples-to-apples because the former utilizes 5 to 8 times the sample size. However, we see that the randomization test with only 5 networks is extremely conservative and poor in terms of power, performing significantly worse than the dependence-robust test despite having 5 times the sample size. These results confirm the discussion in § 5.
7 Conclusion
This paper develops a large-sample theory for a model of network formation with strategic interactions and homophilous agents when the data consists of a small sample of large networks or possibly a single network. We prove a general CLT under a high-level weak dependence condition and provide a general methodology for its verification. We apply the methodology to obtain primitive conditions for a CLT for moments of pairwise stable networks.
There are several important directions for future work. Concentration inequalities for stabilization would be useful, for example, for deriving lower-level conditions for uniform convergence of nonparametric or high-dimensional estimators using network data. It is also of interest to develop efficient alternatives to the inference procedures in § 5 in the case where the data is a single network observation.
Appendix A Bounded Subnetwork Moments
[…]
Let $n]_{h}$ be the set of subsets of $\mathcal{N}_{n}$ of size $h$. The subnetwork count for $a_{h}$ is
$\sum_{H\in[n]_{h}}\bm{1}\{\bm{A}_{H}\cong a_{h}\}.$
This counts “unlabeled” subnetworks isomorphic to $a_{h}$. To rewrite it as a (scaled) network moment $\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$, we observe that this is proportional to the corresponding count of “labeled” subnetworks. Formally,
$h!\sum_{H\in[n]_{h}}\bm{1}\{\bm{A}_{H}\cong a_{h}\}=\sum_{i_{1}\in\mathcal{N}_{n}}\underbrace{\sum_{i_{2}\in\mathcal{N}_{n}}\dots\sum_{i_{h}\in\mathcal{N}_{n}}\bm{1}\{\bm{A}_{\{i_{1},\ldots,i_{h}\}}\cong a_{h}\}}_{\psi_{i_{1}}(\mathcal{N}_{n})}.$ (A.1)
The right-hand side counts labeled subnetworks isomorphic to $a_{h}$, which will be our object of analysis in what follows.
Equation (5.3) of Sheng (2020) uses subnetwork counts to define moment inequalities. She divides the counts by $\binom{n}{h}$ since she considers a setting with many small independent networks. With a single large sparse network, the correct scaling is instead to divide by $n$, resulting in the network moment $n^{-1}\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n})$ for $\psi_{i}(\mathcal{N}_{n})$ defined in (A.1).
Proposition A.1.
Under Assumption 1, Assumption 6(a) holds for node statistics $\psi_{i}(\mathcal{N}_{n})$ of the form given in (A.1).
[…]
$\psi_{1}(\mathcal{N}_{m+k})=\sum_{i_{1}\in\mathcal{N}_{m+k}}\dots\sum_{i_{h-1}\in\mathcal{N}_{m+k}}\bm{1}\{\bm{A}_{\{1,i_{1},\ldots,i_{h-1}\}}\cong a_{h}\}.$ (A.2)
Since $a_{h}$ is a connected network on $\{1,\dots,h\}$, any node in the network is at most path distance h1h-1 from node 1. Therefore, we can replace occurrences of $\mathcal{N}_{m+k}$ in (A.2) with $\mathcal{N}_{\bm{A}}(1,h-1)$. With this change, (A.2) is bounded above by $\lvert\mathcal{N}_{\bm{A}}(1,h-1)\rvert^{h-1}$.
Recall the definition of the network $\bm{M}$ from (SA.2.1). Since $_{ij}\leq M_{ij}, 𝑨\bm{A}$ is a subnetwork of $\bm{M}$, so
$\lvert\mathcal{N}_{\bm{A}}(1,h-1)\rvert^{h-1}\leq\lvert\mathcal{N}_{\bm{M}}(1,h-1)\rvert^{h-1}.$ (A.3)
By Lemma SA.2.2, for $m$ sufficiently large, $\lvert\mathcal{N}_{\bm{M}}(1,h-1)\rvert$ is stochastically dominated by the size of a branching process $\lvert\mathfrak{X}^{M}_{r_{n}}(X_{1},Z_{1};h-1)\rvert$. By Lemma SA.2.4, the distribution of $\lvert\mathfrak{X}^{M}_{r}(x,z;h-1)\rvert$ has exponential tails uniformly in $x,z,r$. It follows that $\lvert\mathcal{N}_{\bm{M}}(1,h-1)\rvert^{h-1}$ has uniformly bounded $p$ moments for any $p>2$.
Supplementary Appendix
This supplement is organized as follows. The first section (§ SA.1) concerns the proof of Theorem 1. In § SA.1.1, we prove an abstract CLT for the “Poissonized” model and “de-Poissonize” the result in § SA.1.2 to obtain a CLT for the original “binomial” model of interest (see § 3.2 for a sketch of the argument). Theorem 1 is a simple corollary of the latter result and is stated at the end of § SA.1.
The second section (§ SA.2) concerns the proof of Theorem 2. In § SA.2.1, we show that the sizes of strategic neighborhoods are stochastically dominated by those of certain branching processes. Then § SA.2.2 presents a series of lemmas culminating in a tail bound for the radius of stabilization. The proof of Theorem 2 is stated at the end of § SA.2 (see § 4.6 for a sketch).
Appendix SA.1 Theorem 1
SA.1.1 Poissonization
This subsection proves a CLT for the Poissonized model in which the set of nodes is $\mathcal{N}_{N_{n}}$ rather than $\mathcal{N}_{n}$, where $N_{n}$ is defined in (11). Throughout, we assume $d_{\psi}=1$. Define $\tilde{\sigma}_{n}^{2}=n^{-1}\text{Var}(\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}}))$.
Theorem SA.1.1.
Suppose $d_{\psi}=1$. Under Assumptions 5 and 6, $\sup_{n}\tilde{\sigma}_{n}^{2}<\infty$, and if additionally $\liminf_{n\rightarrow\infty}\tilde{\sigma}_{n}^{2}>0$, then
$\tilde{\sigma}_{n}^{-1}\frac{1}{\sqrt{n}}\left(\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})-{\bf E}\left[\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})\right]\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,1).$
The proof largely follows that of Theorem 2.1 of Penrose and Yukich (2008) (henceforth PY). We next present several auxiliary lemmas and state the proof at the end of this subsection.
Remark SA.1.1.
That PY’s proof essentially carries over to our setting is perhaps not obvious due to several differences between the setups. First, the definitions of exponential stabilization differ since, as discussed in § 3.3, we only require invariance of node statistics to the removal of nodes outside the radius of stabilization, whereas PY demand invariance to removal and addition. Second, PY’s model has no random-utility shocks $\zeta_{ij}$, but adding these to the model turns out to not affect the argument due to their high degree of independence. Third, PY define $\bm{T}_{N_{n}}$ as a marked Poisson point process, [^10] See e.g. Last and Penrose (2017) for the definition of a marked Poisson point process. which is equivalent to our representation because (a) our set of positions $\bm{X}_{N_{n}}=(X_{i})_{i=1}^{N_{n}}$ has the same distribution as $\mathcal{P}_{nf}$, the Poisson point process on $\mathbb{R}^{d}$ with intensity function $nf(\cdot)$ (Penrose, 2003, Proposition 1.5), and (b) the “marks” $Z_{i}$ associated with each $X_{i}$ are independent across nodes, each only potentially correlated with its own position $X_{i}$. Fourth, PY consider the simpler case where $X_{i}\perp\!\!\!\perp Z_{i}$ for each $i$ since this is sufficient for their applications. We allow for dependence, but this turns out not to affect the argument. Finally, PY consider moments of the form
$\frac{1}{n}\sum_{i=1}^{N_{n}}\xi\big{(}(X_{i},Z_{i}),r_{n}^{-1}\bm{T}_{N_{n}}\big{)}.$
This differs from $\psi_{n}(\cdot)$ because the network is not an argument (nor is $\bm{\zeta}_{N_{n}}$ for reasons previously discussed). However, since the network equals $\lambda_{N_{n}}(r_{n}^{-1}\bm{T}_{N_{n}},\bm{\zeta}_{N_{n}})$ by Assumption 3, we can define
$\xi(i,r_{n}^{-1}\bm{T}_{N_{n}},\bm{\zeta}_{N_{n}})\equiv\psi_{N_{n}}\big{(}i,r_{n}^{-1}\bm{T}_{N_{n}},\bm{\zeta}_{N_{n}},\lambda_{N_{n}}(r_{n}^{-1}\bm{T}_{N_{n}},\bm{\zeta}_{N_{n}})\big{)},$
in which case this coincides with PY’s setup except for the presence of $\bm{\zeta}_{N_{n}}$.
The proof technique is to approximate $\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}})$ by a sum $\sum_{i=1}^{n}W_{i}$ such that the dependence structure of the summands can be characterized in terms of a “dependency graph” $\bm{G}$ that links dependent observations.
Definition 2.
Let $\bm{G}$ be a network on $n$ nodes with self-links, meaning that $G_{ii}=1$ for all $i$. We say $\bm{G}$ is a dependency graph for data $\{W_{i}\}_{i=1}^{n}$ if for any $S_{1},S_{2}\subseteq\mathcal{N}_{n}$ such that $G_{ij}=0$ for all $i\in S_{1}$ and $j\in S_{2}$, we have $\{W_{i}\colon i\in S_{1}\}\perp\!\!\!\perp\{W_{j}\colon j\in S_{2}\}$.
Let $\lVert W\rVert_{p}$ be the $L_{p}$-norm of $W$ and $\Phi(\cdot)$ the standard normal CDF.
Lemma SA.1.1 (Dependency Graph CLT).
Let $q\in(2,3]$ and $W=\sum_{i=1}^{n}W_{i}$. Suppose $\bm{G}$ is a dependency graph for $\{W_{i}\}_{i=1}^{n}$, and let $\Gamma=\max_{i}\sum_{j\neq i}G_{ij}$. Further suppose ${\bf E}[W^{2}]=1$, ${\bf E}[W_{i}]=0$, and $\lVert W_{i}\rVert_{q}\leq\theta$ for all $i$ and some $\theta>0$. Then
$\sup_{t}\,\lvert{\bf P}(W\leq t)-\Phi(t)\rvert\leq 75\,\theta^{q}\,\Gamma^{5(q-1)}\,n.$
Proof. This is Theorem 2.7 of Chen and Shao (2004).
We next introduce some notation. Recall from Assumption 5 the definitions of $\tau_{p}(\cdot)$ and $\eta$. For $p$ in Assumption 6(a) and $d$ the dimension of $X_{1}$, let $q\in(2,3]$ such that $q<p$. Choose $\rho_{n}=(\alpha\log n)^{1/\eta}$ for some $\alpha$ sufficiently large that, for some $C>0$ and all $n\geq 1$,
$\rho_{n}^{d/p}(n\tau_{p}(\rho_{n}))^{(q-2)/(2q)}<Cn^{-4},\quad\tau_{p}(\rho_{n})<Cn^{-3},\quad\rho_{n}^{d}<Cn^{p/(p+2)}.$ (SA.1.1)
We provide an alternative labeling of nodes by position that is more convenient for the argument that follows. Recall from Remark SA.1.1 that $\bm{X}_{N_{n}}$ has the same distribution as the Poisson point process $\mathcal{P}_{nf}$. Let $s_{n}=r_{n}\rho_{n}$ and $V_{n}$ be the number of cubes of the form $Q=\prod_{i=1}^{d}[j_{i}s_{n},(j_{i}+1)s_{n})$ where $j_{i}\in\mathbb{Z}$ for all $i$, such that $Q$ has positive density under $f$. Label the cubes $Q_{1},\dots,Q_{V_{n}}$, and note that they have side lengths $s_{n}=o(1)$. Since $f$ has bounded support as assumed at the start of § 2, $V_{n}$ is finite, and
$V_{n}=O(n\rho_{n}^{-d}).$ (SA.1.2)
Fix any $1\leq i\leq V_{n}$. Note that $N_{i}\equiv\lvert\bm{X}_{N_{n}}\cap Q_{i}\rvert$ is distributed $\text{Poisson}(\nu_{i})$ for $\nu_{i}=n\int_{Q_{i}}f(x)\,\text{d}x$. Relabel the points of $\bm{X}_{N_{n}}\cap Q_{i}$ as $X_{i1},\dots,X_{iN_{i}}$, which, conditional on $N_{i}$, are i.i.d. draws from a density $f(\cdot)/\int_{Q_{i}}f(x)\,\text{d}x$. Under this relabeling,
[…]
In what follows, we will often use this double-indexed labeling for nodes, writing $X_{ij}$ for node positions and $\psi_{ij}(\mathcal{N}_{N_{n}})$ for node statistics. That is, suppose under this relabeling that node $k\in\{1,\ldots,N_{n}\}$ is relabeled $ij$. Then $\psi_{ij}(\mathcal{N}_{N_{n}})$ equals $\psi_{k}(\mathcal{N}_{N_{n}})$, $k$’s node statistic under the Poissonized model with $N_{n}$ nodes.
The next lemma is mostly a restatement of Lemma 4.3 of Penrose and Yukich (2008), translated to our notation. [^11] The translation works as follows: replace $\lambda$ (their notation) with $n$ (our notation), $\lambda^{-1/d}$ with $r_{n}$, $\kappa$ with $f$, $A_{\lambda}$ with $\text{supp}(f)$, $f$ with the function mapping any real number to 1, $\mathcal{P}_{\lambda}$ with $\bm{X}_{N_{n}}$, and $\xi_{\lambda}(X_{i,j},\mathcal{P}_{\lambda})$ with $\psi_{ij}(\mathcal{N}_{N_{n}})$. For any $B\subseteq\mathbb{R}^{d}$, define $\mathcal{N}_{N_{n}}^{\prime}(B)=\{i\in\mathcal{N}_{N_{n}}^{\prime}\colon X_{i}\in B\}$. Unlike the definition of $\mathcal{N}_{n}(B)$ prior to Definition 1, this involves a random number of nodes $N_{n}$ and unscaled node positions $X_{i}$.
Lemma SA.1.2.
Under Assumption 6(a), for $p$ given in the assumption and $1<q<p$, there exists $C>0$ such that $\lVert\sum_{j=1}^{\infty}\lvert\psi_{ij}(\mathcal{N}_{N_{n}}^{\prime}(S_{ni}))\rvert\bm{1}\{j\leq N_{i}\}\rVert_{q}\leq C\rho_{n}^{d(p+1)/p}$ for any $Q_{i}\subseteq S_{ni}\subseteq\mathbb{R}^{d}$, $1\leq i\leq V_{n}$, and $n\in\mathbb{N}$.
Since the lemma is a fairly straightforward consequence of Assumption 6(a) and properties of the Poisson distribution, we omit the proof. The main difference relative to PY is the presence of random-utility shocks $\bm{\zeta}_{N_{n}}$. The proof only involves quantities conditional on $N_{i}$ and integrating over $\bm{\zeta}_{N_{n}}$. Because the latter is independent of $N_{i}$ for all $i$, their presence does not materially affect PY’s original argument.
The last lemma is used to construct a dependency graph.
Lemma SA.1.3.
For any two disjoint sets $S_{1},S_{2}\subseteq\mathbb{R}^{d}$,
$\{(X_{i},X_{j},Z_{i},Z_{j},\zeta_{ij})\colon i,j\in\mathcal{N}_{N_{n}}(S_{1})\}\perp\!\!\!\perp\{(X_{i},X_{j},Z_{i},Z_{j},\zeta_{ij})\colon i,j\in\mathcal{N}_{N_{n}}(S_{2})\}.$ (SA.1.4)
Proof. This is a consequence of the spatial independence property of Poisson point processes (e.g. Last and Penrose, 2017, Theorem 5.2). Here we provide an elementary proof from first principles. Let $N(S)=\lvert\mathcal{N}_{N_{n}}(S)\rvert$, $T_{i}=(X_{i},Z_{i})$, $\bm{T}_{m}=(T_{i})_{i=1}^{m}$, $\bm{T}_{S}=(T_{i}\colon X_{i}\in S)$, $\bm{T}_{m,\dots,n}=(T_{i})_{i=m}^{n}$, and $\bm{X}_{m,\dots,n}=(X_{i})_{i=m}^{n}$. First consider the simple case where $S_{1},S_{2}$ partition $\text{supp}(f)$. Fix $m_{1},m_{2},m\in\mathbb{N}$ such that $m_{1}+m_{2}=m$. For any $\mathcal{U}_{m}\subseteq\mathcal{T}^{m}$ and $\mathcal{Z}_{m}\subseteq\text{supp}(\bm{\zeta}_{m})$,
${\bf P}(N(S_{1})$ $=m_{1},\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{S_{1}}\in\mathcal{Z}_{m_{1}},N(S_{2})=m_{2},\bm{T}_{S_{2}}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{S_{2}}\in\mathcal{Z}_{m_{2}})$
$\begin{aligned} ={\bf P}(N_{n}=m){\bf P}(N(S_{1})=m_{1},\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},&\bm{\zeta}_{S_{1}}\in\mathcal{Z}_{m_{1}},N(S_{2})=m_{2},\\ &\bm{T}_{S_{2}}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{S_{2}}\in\mathcal{Z}_{m_{2}}\mid N_{n}=m)\end{aligned}$
$\begin{aligned} =\frac{e^{-n}n^{m}}{m!}\frac{m!}{m_{1}!m_{2}!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\,&\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}},\bm{X}_{m_{1}+1,\dots,n}\in S_{2},\\ &\bm{T}_{m_{1}+1,\dots,n}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{m_{1}+1,\dots,n}\in\mathcal{Z}_{m_{2}})\end{aligned}$
$\begin{aligned} =\frac{e^{-n{\bf P}(X_{1}\in S_{1})}n^{m_{1}}}{m_{1}!}&{\bf P}(\bm{X}_{m_{1}}\in S_{1},\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}})\\ &\times\frac{e^{-n{\bf P}(X_{1}\in S_{2})}n^{m_{2}}}{m_{2}!}{\bf P}(\bm{X}_{m_{2}}\in S_{2},\bm{T}_{m_{2}}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{m_{2}}\in\mathcal{Z}_{m_{2}}).\end{aligned}$
The third equality is true because types are identically distributed, and $N(S_{1})=m_{1}$ and $N(S_{2})=m_{2}$ imply $N_{n}=m$ given that $S_{1},S_{2}$ partition the space. The last line is true because disjoint sets of nodes have independent types and random-utility shocks.
We perform similar calculations for the marginal distribution:
$\begin{aligned} {\bf P}(N(S_{1})=m_{1},\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{S_{1}}\in\,\,&\mathcal{Z}_{m_{1}})=\sum_{m\geq m_{1}}{\bf P}(N_{n}=m){\bf P}(N(S_{1})=m_{1},\\ &\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{S_{1}}\in\mathcal{Z}_{m_{1}},N(S_{1}^{c})=m-m_{1}\mid N_{n}=m)\end{aligned}$
$=\sum_{m\geq m_{1}}\frac{e^{-n}n^{m}}{m!}\frac{m!}{m_{1}!(m-m_{1})!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}},\bm{X}_{m_{1}+1,\dots,n}\in S_{1}^{c})$
$\begin{aligned} =\sum_{m\geq m_{1}}\frac{e^{-n{\bf P}(X_{1}\in S_{1})}n^{m_{1}}}{m_{1}!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\,\bm{T}_{m_{1}}\in\,\,&\mathcal{U}_{m_{1}},\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}})\\ &\times\frac{e^{-n{\bf P}(X_{1}\in S_{1}^{c})}n^{m-m_{1}}}{(m-m_{1})!}{\bf P}(\bm{X}_{m_{1}+1,\dots,n}\in S_{1}^{c})\end{aligned}$
$\begin{aligned} =\frac{e^{-n{\bf P}(X_{1}\in S_{1})}n^{m_{1}}}{m_{1}!}{\bf P}(\bm{X}_{m_{1}}\in S_{1},\bm{T}_{m_{1}}\in\mathcal{U}_{m_{1}},\,\,&\bm{\zeta}_{m_{1}}\in\mathcal{Z}_{m_{1}})\\ &\times\underbrace{\sum_{m\geq m_{1}}\frac{e^{-n{\bf P}(X_{1}\in S_{1}^{c})}(n{\bf P}(X_{1}\in S_{1}^{c}))^{m-m_{1}}}{(m-m_{1})!}}_{1}.\end{aligned}$
Therefore,
${\bf P}(N(S_{1})=m_{1},\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{S_{1}}\in\mathcal{Z}_{m_{1}},N(S_{2})=m_{2},\bm{T}_{S_{2}}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{S_{2}}\in\mathcal{Z}_{m_{2}})\\ ={\bf P}(N(S_{1})=m_{1},\bm{T}_{S_{1}}\in\mathcal{U}_{m_{1}},\bm{\zeta}_{S_{1}}\in\mathcal{Z}_{m_{1}}){\bf P}(N(S_{2})=m_{2},\bm{T}_{S_{2}}\in\mathcal{U}_{m_{2}},\bm{\zeta}_{S_{2}}\in\mathcal{Z}_{m_{2}}),$
which proves (SA.1.4) for the simple case where $S_{1},S_{2}$ partition $\text{supp}(f)$. For the general case in which $S_{1},S_{2}$ does not partition $\text{supp}(f)$, since the sets are disjoint, we can partition the support into two sets $S_{1}^{\prime},S_{2}^{\prime}$ such that $S_{1}\subseteq S_{1}^{\prime}$ and $S_{2}\subseteq S_{2}^{\prime}$ and apply the result for the simple case.
Proof of Theorem SA.1.1. We mostly follow PY’s proof of Theorem 2.1 with some minor modifications, reorganization, and elaboration. That $\sup_{n}\tilde{\sigma}_{n}^{2}<\infty$ is a consequence of Lemma SA.1.7 below. Abbreviate
$\Lambda_{n}=\sum_{i=1}^{N_{n}}\psi_{i}(\mathcal{N}_{N_{n}}).$
We prove the following stronger result: for $q\in(2,3]$ with $q<p$ and $\eta$ from Assumption 5, there exists $C>0$ such that for all $n$ sufficiently large,
$\sup_{t}\,\lvert{\bf P}\big{(}\text{Var}(\Lambda_{n})^{-1/2}(\Lambda_{n}-{\bf E}[\Lambda_{n}])\leq t\big{)}-\Phi(t)\rvert\leq C(\log n)^{dq/\eta}n\text{Var}(\Lambda_{n})^{-q/2}.$ (SA.1.5)
Since $\text{Var}(\Lambda_{n})^{-1}=O(n^{-1})$ by assumption, the right-hand side of (SA.1.5) tends to zero, which proves the theorem.
Under representation (SA.1.3), $\Lambda_{n}=\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}_{N_{n}})$. The key idea is to approximate this with a similar sum to which the dependency graph CLT (Lemma SA.1.1) more usefully applies. For any $1\leq i\leq V_{n}$, let $Q_{i}^{+}=\{x\in\mathbb{R}^{d}\colon\lVert x-y\rVert\leq r_{n}\rho_{n}\text{ for some }y\in Q_{i}\}$, the $r_{n}\rho_{n}$-neighborhood of $Q_{i}$. Recall the definition of $\mathcal{N}_{N_{n}}^{\prime}(S)$ prior to Lemma SA.1.2. Define
$\Lambda_{n}^{\prime}=\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}^{\prime}_{N_{n}}(Q_{i}^{+}))\quad\text{and}\quad\Psi_{i}=\text{Var}(\Lambda_{n}^{\prime})^{-1/2}\sum_{j=1}^{N_{i}}\psi_{ij}(\mathcal{N}^{\prime}_{N_{n}}(Q_{i}^{+})),$
so that $\text{Var}(\Lambda_{n}^{\prime})^{-1/2}\Lambda_{n}^{\prime}=\sum_{i=1}^{V_{n}}\Psi_{i}$.
Recall $\bm{R}_{i}(n,r_{n})$ from Definition 1. The relabeling introduced above indexes nodes by $ij$ for $1\leq i\leq V_{n}$ and $1\leq j\leq N_{i}$, so we instead write $\bm{R}_{ij}(n,r_{n})$. For $\rho_{n}$ defined prior to (SA.1.1), define the event $\mathcal{E}_{ij}=\{\bm{R}_{ij}(N_{n},r_{n})\leq\rho_{n}\}$.
[…]
By Lemma SA.1.2, Minkowski’s inequality, and (SA.1.2), there exists $C^{\prime}>0$ such that for all $n$,
$\max\big{\{}\lVert\Lambda_{n}\rVert_{q},\lVert\Lambda_{n}^{\prime}\rVert_{q}\big{\}}\leq CV_{n}\rho_{n}^{d(p+1)/p}\leq C^{\prime}n\rho_{n}^{d/p}.$ (SA.1.6)
By the Hölder and Minkowski inequalities,
$\lVert\Lambda_{n}-\Lambda_{n}^{\prime}\rVert_{2}\leq\lVert\Lambda_{n}-\Lambda_{n}^{\prime}\rVert_{q}{\bf P}(\mathcal{E}_{n}^{c})^{0.5-1/q}\leq(\lVert\Lambda_{n}\rVert_{q}+\lVert\Lambda_{n}^{\prime}\rVert_{q}){\bf P}(\mathcal{E}_{n}^{c})^{(q-2)/(2q)}.$ (SA.1.7)
By the law of iterated expectations,
${\bf P}\left(\mathcal{E}_{n}^{c}\right)\leq{\bf E}\left[\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}\bm{1}\{\mathcal{E}_{ij}^{c}\}\right]={\bf E}\left[\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}{\bf E}[\bm{1}\{\mathcal{E}_{ij}^{c}\}\mid\bm{X}_{N_{n}},N_{n}]\right]$
As in the setting of Penrose and Yukich (2008), this is an expectation of pointwise sums (i.e. over the points of $\bm{X}_{N_{n}}$) of functionals of a Poisson point process $\bm{X}_{N_{n}}$. We may then apply Theorem 1.6 of Penrose (2003) to obtain
${\bf P}\left(\mathcal{E}_{n}^{c}\right)\leq{\bf E}\left[\sum_{i=1}^{V_{n}}\sum_{j=1}^{N_{i}}{\bf E}[\bm{1}\{\mathcal{E}_{ij}^{c}\}\mid\bm{X}_{N_{n}},N_{n}]\right]\\ =n\int_{\mathbb{R}^{d}}{\bf P}\left(\bm{R}_{ij}(N_{n}+1,r_{n})>\rho_{n}\mid X_{ij}=x\right)f(x)\,\text{d}x\leq n\,\tau_{p}(\rho_{n})$ (SA.1.8)
for $n$ sufficiently large. [^12] The quantity $h(Y,\mathcal{P}_{\lambda})$ in Theorem 1.6 corresponds to ${\bf E}[\bm{1}\{\mathcal{E}_{ij}^{c}\}\mid\bm{X}_{N_{n}},N_{n}]$, which is a function of the Poisson process $\mathcal{P}_{\lambda}=\bm{X}_{N_{n}}$ and unit $ij$’s position $Y=X_{ij}$. By (SA.1.6), (SA.1.7), (SA.1.8), our choice of $\rho_{n}$ (SA.1.1), and Assumption 5, there exist $C,C^{\prime}>0$ such that for all $n$
$\lVert\Lambda_{n}-\Lambda_{n}^{\prime}\rVert_{2}\leq Cn\rho_{n}^{d/p}(n\tau_{p}(\rho_{n}))^{(q-2)/(2q)}\leq C^{\prime}n^{-3}.$ (SA.1.9)
Dependency Graph. We construct a dependency graph $\bm{G}$ for the summands $\{\Psi_{i}\}_{i=1}^{V_{n}}$ of $\Lambda_{n}^{\prime}$ by setting $G_{ij}=\bm{1}\{\inf\{\lVert x-y\rVert\colon x\in Q_{i},y\in Q_{j}\}\leq 2r_{n}\rho_{n}\}$. To see that this is a dependency graph, observe that the value of $\Psi_{i}$ is determined solely by the types of and random-utility shocks between nodes positioned in $Q_{i}^{+}$. Then whenever $\inf\{\lVert x-y\rVert\colon x\in Q_{i},y\in Q_{j}\}>2r_{n}\rho_{n}$, we have $\{\Psi_{k}\colon k\in Q_{i}\}\perp\!\!\!\perp\{\Psi_{\ell}\colon\ell\in Q_{j}\}$ by Lemma SA.1.3. Hence, $\bm{G}$ is a dependency graph.
Note that (SA.1.5) holds trivially when $\text{Var}(\Lambda_{n})<1$ for $n$ sufficiently large, so we henceforth consider the case where it is at least 1. We apply Lemma SA.1.1 to $\{W_{i}\}_{i=1}^{V_{n}}$ for $W_{i}=\Psi_{i}-{\bf E}[\Psi_{i}]$. To determine a good choice of $\theta$ for applying the lemma, notice by Lemma SA.1.2 that
[…]
We set $\theta$ to be the right-hand side. There are at most $5^{d}$ other cubes at most distance $2r_{n}\rho_{n}$ from cube $Q_{i}$, so $\Gamma\leq 5^{d}$ for $\Gamma=\max_{i}\sum_{j\neq i}G_{ij}$. By Lemma SA.1.1 and (SA.1.2), for $W=\sum_{i=1}^{V_{n}}(\Psi_{i}-{\bf E}[\Psi_{i}])$, there exists $C>0$ such that for all $n$
$\sup_{t}\,\lvert{\bf P}(W\leq t)-\Phi(t)\rvert\leq Cn\rho_{n}^{-d}\,\text{Var}(\Lambda_{n}^{\prime})^{-q/2}\rho_{n}^{dq(p+1)/p}$ (SA.1.10)
Right side of (SA.1.5). We next show that the right-hand side of (SA.1.10) is at most the right-hand side of (SA.1.5). First, we bound the discrepancy between $\text{Var}(\Lambda_{n}^{\prime})$ and $\text{Var}(\Lambda_{n})$. Trivially,
$\text{Var}(\Lambda_{n})=\text{Var}(\Lambda_{n}^{\prime})+\text{Var}(\Lambda_{n}-\Lambda_{n}^{\prime})+2\text{Cov}(\Lambda_{n}^{\prime},\Lambda_{n}-\Lambda_{n}^{\prime}).$ (SA.1.11)
We can bound the first term on the right-hand side as follows. By construction, $\Lambda_{n}^{\prime}$ is the sum of $V_{n}$ random variables that have second moments bounded by a constant times $\rho_{n}^{2d(p+1)/p}$ due to Lemma SA.1.2 for $q=2$. Also, as previously established, the covariance between any pair of these variables is zero when the corresponding nodes are respectively positioned in cubes $Q_{i}$ and $Q_{j}$ such that any element of $Q_{i}$ is more than $2r_{n}\rho_{n}$ away from any element of $Q_{j}$. Otherwise, the covariance is bounded by a constant times $\rho_{n}^{2d(p+1)/p}$ by Cauchy-Schwarz. Consequently, by (SA.1.2),
$\text{Var}(\Lambda_{n}^{\prime})=O(n\rho_{n}^{d(p+2)/p}).$ (SA.1.12)
Combining (SA.1.9), (SA.1.11), and (SA.1.12), by our choice of $\rho_{n}$ (SA.1.1) and Cauchy-Schwarz,
$\lvert\text{Var}(\Lambda_{n})-\text{Var}(\Lambda_{n}^{\prime})\rvert\leq Cn^{-2}.$ (SA.1.13)
Given $\text{Var}(\Lambda_{n})\geq 1$ from above, this implies that, for $n$ sufficiently large, $\text{Var}(\Lambda_{n}^{\prime})\geq\text{Var}(\Lambda_{n})/2$. Since $q<p$,
[…]
which delivers the bound on the right-hand side of (SA.1.5), given $\rho_{n}=(\alpha\log n)^{1/\eta}$.
Left side of (SA.1.5). However, (SA.1.10) and (SA.1.14) do not immediately deliver (SA.1.5) because they apply to $W=\text{Var}(\Lambda_{n}^{\prime})^{-1/2}(\Lambda_{n}^{\prime}-{\bf E}[\Lambda_{n}^{\prime}])$ rather than the target $\text{Var}(\Lambda_{n})^{-1/2}(\Lambda_{n}-{\bf E}[\Lambda_{n}])$. Let $\tilde{W}=\text{Var}(\Lambda_{n}^{\prime})^{-1/2}(\Lambda_{n}-{\bf E}[\Lambda_{n}])$. Using the Lipschitz property of $\Phi(\cdot)$ and (SA.1.14), elementary calculations yield
$\sup_{t}\,\lvert{\bf P}(\tilde{W}\leq t)-\Phi(t)\rvert\leq C\beta+Cn\text{Var}(\Lambda_{n})^{-q/2}\rho_{n}^{dq}+{\bf P}(\lvert\tilde{W}-W\rvert\geq\beta)$ (SA.1.15)
for any $\beta>0$ (Penrose and Yukich, 2008, eq. (4.19)). Also,
[…]
Since (SA.1.9) implies ${\bf E}[\lvert\Lambda_{n}-\Lambda_{n}^{\prime}\rvert]\leq Cn^{-3}$, we have (SA.1.16) $\leq Cn^{-3}$ under $\mathcal{E}_{n}$. On the other hand, ${\bf P}(\mathcal{E}_{n}^{c})\leq Cn^{-2}$ by (SA.1.8) and (SA.1.1). So choosing $\beta=Cn^{-3}$ in (SA.1.15), we obtain
$\sup_{t}\,\lvert{\bf P}(\tilde{W}\leq t)-\Phi(t)\rvert\leq Cn\text{Var}(\Lambda_{n})^{-q/2}\rho_{n}^{dq}+Cn^{-2}.$ (SA.1.17)
With these pieces in place, we may now prove (SA.1.5). We have
$\sup_{t}\,\lvert{\bf P}\big{(}\text{Var}(\Lambda_{n})^{-1/2}(\Lambda_{n}-{\bf E}[\Lambda_{n}]\big{)}\leq t)-\Phi(t)\rvert\\ \leq\sup_{t}\bigg{|}{\bf P}\left(\tilde{W}\leq t\left(\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}\right)^{1/2}\right)-\Phi\left(t\left(\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}\right)^{1/2}\right)\bigg{|}\\ +\sup_{t}\bigg{|}\Phi\left(t\left(\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}\right)^{1/2}\right)-\Phi(t)\bigg{|}.$ (SA.1.18)
For $n$ large, $\text{Var}(\Lambda_{n}^{\prime})\geq\text{Var}(\Lambda_{n})/2\geq 0.5$ since $\text{Var}(\Lambda_{n})\geq 1$. By (SA.1.13), there exists $C>0$ such that for all $n>0$,
$\bigg{|}t\left(\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}\right)^{1/2}-t\bigg{|}\leq\lvert t\rvert\bigg{|}\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}-1\bigg{|}\leq Cn^{-2}\lvert t\rvert$. (SA.1.19)
Letting $\phi(\cdot)$ denote the normal density, for all $s\leq t$, $\lvert\Phi(s)-\Phi(t)\rvert\leq(t-s)\max_{u\in[s,t]}\phi(u)$. Then using (SA.1.19),
$\sup_{t}\bigg{|}\Phi\left(t\left(\frac{\text{Var}(\Lambda_{n})}{\text{Var}(\Lambda_{n}^{\prime})}\right)^{1/2}\right)-\Phi(t)\bigg{|}\leq Cn^{-2}\sup_{t}\big{\{}\lvert t\rvert\max_{u\in[t-tCn^{-2},t+tCn^{-2}]}\phi(u)\big{\}}\leq Cn^{-2}.$
Combined with (SA.1.17) and (SA.1.18), we obtain
$\sup_{t}\,\lvert{\bf P}\big{(}\text{Var}(\Lambda_{n})^{-1/2}(\Lambda_{n}-{\bf E}[\Lambda_{n}])\leq t\big{)}-\Phi(t)\rvert\leq C(\log n)^{dq/\eta}n\text{Var}(\Lambda_{n})^{-q/2}+O(n^{-2}).$
By (SA.1.12) and (SA.1.13), $\text{Var}(\Lambda_{n})=O(n\rho_{n}^{d(p+2)/p})$, so by (SA.1.1) and the fact that $q\in(2,3]$, the first term on the right-hand side dominates the second term, and (SA.1.5) follows.
SA.1.2 de-Poissonization
As in the previous subsection, assume $d_{\psi}=1$. The next result shows that, given a CLT for the Poissonized model with node set $\mathcal{N}_{N_{n}}$, a CLT also holds for the original “binomial” model of interest with node set $\mathcal{N}_{n}$. Define $\sigma_{n}^{2}=n^{-1}\text{Var}(\sum_{i=1}^{n}\psi_{i}(\mathcal{N}_{n}))$.
Theorem SA.1.2.
Suppose $d_{\psi}=1$. Under Assumptions 5 and 6, $\sup_{n}\sigma_{n}^{2}<\infty$. Further suppose that $\liminf_{n\rightarrow\infty}\sigma_{n}^{2}>0$. Then $\liminf_{n\rightarrow\infty}\tilde{\sigma}_{n}^{2}>0$ and
$\sigma_{n}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\big{(}\psi_{i}(\mathcal{N}_{n})-{\bf E}[\psi_{i}(\mathcal{N}_{n})]\big{)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,1).$
The proof is stated at the end of this subsection and uses the next four lemmas. Lemma SA.1.4 establishes an asymptotic equivalence between $\bm{X}_{n}$ and $\bm{X}_{N_{n}}$ and is taken directly from Penrose and Yukich (2003). Lemma SA.1.5 proves that the node statistics of two nodes positioned a fixed distance apart are asymptotically uncorrelated. Lemma SA.1.6 proves that add-one costs of two nodes at a fixed distance apart are asymptotically uncorrelated. Finally, Lemma SA.1.7 shows that various moments are uniformly bounded.
Remark SA.1.2 (Related Literature).
These results are adaptations of lemmas in Penrose (2007), so we compare and contrast our setup with his. First, the differences discussed in Remark SA.1.1 persist here. Second, the proofs of Lemmas SA.1.5 and SA.1.6 need some modification because our setups differ, most notably the definition of stabilization. Third, to prove the lemmas, we define a coupling that follows the ideas in Penrose (2007) and Penrose and Yukich (2003), but the details are different because we need to accommodate random-utility shocks $\zeta_{ij}$, dependence between $X_{i}$ and $Z_{i}$, and differences in notation. Fourth, Penrose considers a generalization of $\psi_{n}(\cdot)$ that is a measure in that it takes as an argument a subset of $\mathbb{R}^{d}$. This generalization is unnecessary for our purposes.
The lemmas that follow utilize the following coupling whose essential idea follows Penrose and Yukich (2003), p. 298. As in § 2.2, define i.i.d. types $\{(X_{i},Z_{i})\}_{i\in\mathbb{N}}$ and i.i.d. random-utility shocks $\{\zeta_{ij}\colon i\neq j,\{i,j\}\subseteq\mathbb{N}\}$, with the two sets mutually independent. For any $n\in\mathbb{N}$, let $N_{n}^{\prime}$ be an independent copy of $N_{n}$ (defined in § 3.2), independent of all other primitives.
  • For any $m\in\mathbb{N}$, let each node $i\in\mathcal{N}_{m}$ be assigned type $T_{i}=(X_{i},Z_{i})$ and each node pair $i,j\in\mathcal{N}_{m}$ assigned random-utility shock $\zeta_{ij}$. Define
    $\bm{X}_{m}=(X_{5},X_{7},X_{9},\dots,X_{2m+3}).$ (SA.1.20)
    This will be the array of node positions, which notably excludes even-labeled nodes and nodes $1,\ldots,4$. The latter are reserved for particular use in the proof of Lemma SA.1.5 below, while the even-labeled nodes will be used in the next bullet point. Observe that (SA.1.20) has the same distribution as the more conventional-looking array $(X_{1},X_{2},\dots,X_{m})$.
  • Fix any $x,y\in\mathbb{R}^{d}$. Let $F_{x}$ be the set of points in $\mathbb{R}^{d}$ closer to $x$ than $y$ and $F_{y}$ the remaining set of points (break ties arbitrarily). Let $\mathcal{P}_{nf}^{*,x}$ be the restriction of $\tilde{\bm{X}}_{N_{n}}\equiv(X_{5},X_{7},X_{9},\dots,X_{2N_{n}+3})$ to $F_{x}$. Let $\mathcal{Q}_{nf}^{*,x}$ be the restriction of $\tilde{\bm{X}}_{N_{n}^{\prime}}\equiv(X_{6},X_{8},X_{10}\dots,X_{2N_{n}^{\prime}+4})$ to $F_{y}$. Construct $\mathcal{P}_{nf}^{x}$ by concatenating the two vectors $\mathcal{P}_{nf}^{*,x}$ and $\mathcal{Q}_{nf}^{*,x}$. As the notation suggests, for any $n$, the set consisting of the components of $\mathcal{P}_{nf}^{x}$ has the same distribution as an inhomogeneous Poisson point process on $\mathbb{R}^{d}$ with intensity function $nf(\cdot)$. This follows from Poisson superposition (Last and Penrose, 2017, Theorem 3.3) since $F_{x}$ and $F_{y}$ partition $\mathbb{R}^{d}$. Hence, $\mathcal{P}_{nf}^{x}$ has the same distribution as $\bm{X}_{N_{n}}$ used in § SA.1.1.
  • Similarly, let $\mathcal{P}_{nf}^{*,y}$ be the restriction of $\tilde{\bm{X}}_{N_{n}}$ to $F_{y}$ and $\mathcal{Q}_{nf}^{*,y}$ the restriction of $\tilde{\bm{X}}_{N_{n}^{\prime}}$ to $F_{x}$. Construct $\mathcal{P}_{nf}^{y}$ by concatenating the two vectors $\mathcal{P}_{nf}^{*,y}$ and $\mathcal{Q}_{nf}^{*,y}$. Then $\mathcal{P}_{nf}^{y}$ has the same distribution as $\bm{X}_{N_{n}}$.
The point of this construction is that, by switching the odds and evens on the two half-sets $F_{x}$ and $F_{y}$, we have $\mathcal{P}_{nf}^{x}\perp\!\!\!\perp\mathcal{P}_{nf}^{y}$. As a consequence, we obtain the following independence property. Let $r_{n}^{-1}\mathcal{P}_{nf}^{x}=\{r_{n}^{-1}X\colon X\in\mathcal{P}_{nf}^{x}\}$. For any $S_{1},S_{2}\subseteq\mathbb{R}^{d}$ with $S_{1}\cap S_{2}=\varnothing$, define $\mathcal{N}_{x}(S_{1})=\{k\in\mathbb{N}\colon r_{n}^{-1}X_{k}\in r_{n}^{-1}\mathcal{P}_{nf}^{x}\cap S_{1}\}$ and $\mathcal{N}_{y}(S_{2})$ similarly. Then
$\{(X_{i},X_{j},Z_{i},Z_{j},\zeta_{ij})\colon i,j\in\mathcal{N}_{x}(S_{1})\}\perp\!\!\!\perp\{(X_{i},X_{j},Z_{i},Z_{j},\zeta_{ij})\colon i,j\in\mathcal{N}_{y}(S_{2})\}.$ (SA.1.21)
This is similar to (SA.1.4) but using two “interleaved” Poisson processes $\mathcal{P}_{nf}^{x}$ and $\mathcal{P}_{nf}^{y}$. The interleaving idea of taking two different “halves” on $F_{x}$ and $F_{y}$ of two independent Poisson processes (in our case, the odds and the evens) is taken from Penrose and Yukich (2003). Our construction only differs by building the two processes from the odd and even elements of the same countable set of random vectors.
For any $x,y\in\mathbb{R}^{d}$, $r>0$, and $\bm{X}\subseteq\mathbb{R}^{d}$, let $\tau_{x,r}y=x+r^{-1}(y-x)$ and $\tau_{x,r}\bm{X}=\{x+r^{-1}(y-x)\colon y\in\bm{X}\}$. This operator dilates the point set $\bm{X}$ relative to point $x$, whereas in the main paper, we use $r^{-1}\bm{X}=\tau_{\bm{0},r}\bm{X}$, which specifically dilates $\bm{X}$ relative to the origin. For technical reasons, it will be convenient to consider dilations relative to other locations. The next coupling lemma shows that the “binomial process” $\tau_{x,r_{n}}\bm{X}_{n}$ and Poisson process $\tau_{x,r_{n}}\mathcal{P}_{nf}^{x}$ coincide on neighborhoods of $x$.
Lemma SA.1.4.
For any $R>0$, $x,y\in\mathbb{R}^{d}$, and sequence $\{\ell_{n}\}_{n\in\mathbb{N}}$ with $\ell_{n}/n\rightarrow 1$,
${\bf P}\big{(}\tau_{x,r_{n}}\bm{X}_{\ell_{n}}\cap B(x,R)=\tau_{x,r_{n}}\mathcal{P}_{nf}^{x}\cap B(x,R)\big{)}\rightarrow 1.$
Proof. This is Lemma 3.1 of Penrose and Yukich (2003).
Remark SA.1.3.
The coupling will be used to show that the node statistic of a node positioned at $x$ is asymptotically uncorrelated with that of a node positioned at $y$, for any fixed $x,y\in\mathbb{R}^{d}$ (Lemma SA.1.5). The idea is that by stabilization, $x$’s statistic is primarily determined by nodes near $x$ and likewise for $y$, and by Lemma SA.1.4 and stabilization, nodes near $x$ will be elements of $\mathcal{P}_{nf}^{x}$ and nodes near $y$ of $\mathcal{P}_{nf}^{y}$, which are independent. Note this does not mean node statistics are essentially i.i.d. because in the continuum limit, there are many nodes near $x$ and many near $y$.
Recall that $\Phi_{z}(\cdot\mid x)$ is the conditional distribution of $Z_{1}$ given $X_{1}=x$. Fix any $w,x,y,z\in\text{supp}(f)$, $z_{1}\in\text{supp}(\Phi_{z}(\cdot\mid x))$, $z_{2}\in\text{supp}(\Phi_{z}(\cdot\mid y))$, $z_{3}\in\text{supp}(\Phi_{z}(\cdot\mid x+r_{n}z))$, and $z_{4}\in\text{supp}(\Phi_{z}(\cdot\mid y+r_{n}w))$. Consider a modification of the coupling construction in which we redefine
$X_{1}\equiv x,X_{2}\equiv y,X_{3}\equiv x+r_{n}z,X_{4}\equiv y+r_{n}w,\quad\text{and}$
$Z_{k}\equiv z_{k}\quad\text{for all }k=1,\ldots,4.$
That is, we replace the types associated with the first four positions in the coupling construction with non-random values. We refer to this as “DGP $(\star)$” in what follows. Note that the DGP implicitly depends on $n$ and the eight non-random vectors defined above. It will be used to construct objects analogous to quantities defined prior to Assumptions 5 and 6(a) involving a $\star$ superscript.
For any finite $J\subseteq\mathbb{N}$, $i\in J$, and $v\in\mathbb{R}^{d}$, define
$\psi_{i,v}(J)=\psi_{\lvert J\rvert}\big{(}i,\tau_{v,r_{n}}\bm{T}_{J},\bm{\zeta}_{J},\lambda_{\lvert J\rvert}(\tau_{v,r_{n}}\bm{T}_{J},\bm{\zeta}_{J})\big{)}$
where $\tau_{v,r_{n}}\bm{T}_{J}=((\tau_{v,r_{n}}X_{i},Z_{i}))_{i\in J}$ and $\bm{\zeta}_{J}=(\zeta_{ij})_{i\neq j;i,j\in J}$. In other words, $\psi_{i,v}(J)$ is $i$’s node statistic under a model in which the node set is $J$ and positions are dilated relative to $v$. For $v=\bm{0}$, $\psi_{i,v}(J)$ is a representation of $\psi_{i}(J)$ defined in (9). Recall from (SA.1.21) the definition of $\mathcal{N}_{x}(S_{1})$, and abbreviate $\mathcal{N}_{x}\equiv\mathcal{N}_{x}(\mathbb{R}^{d})$. Then for $u\in\{x,y\}\subseteq\mathbb{R}^{d}$, $\psi_{i,x}(\mathcal{N}_{u})$ corresponds to $i$’s statistic under the Poissonized model where the number of nodes $N_{n}$ is random. When $u=m\in\mathbb{N}$, it corresponds to $i$’s statistic under the usual “binomial” model with non-random node set $\mathcal{N}_{m}$. Finally, for any $a,b\in\mathbb{N}$ with $a\leq b$, define $\mathcal{N}_{a;b}=\{a,\ldots,b+a-1\}$.
Lemma SA.1.5.
Let $\{\ell_{n}\}_{n\in\mathbb{N}}$ and $\{m_{n}\}_{n\in\mathbb{N}}$ be sequences of natural numbers such that $\ell_{n}/n$ and $m_{n}/n$ tend to one. Define
$\begin{aligned} V_{1}(x,y)=\big{(}&\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1\}),\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,2\}),\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,3\}),\\ &\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,2,3\}),\psi_{1,x}(\mathcal{N}_{5;m_{n}}\cup\{1\}),\psi_{1,x}(\mathcal{N}_{5;m_{n}}\cup\{1,2\}),\\ &\psi_{1,x}(\mathcal{N}_{5;m_{n}}\cup\{1,2,3\}),\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{2\}),\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,2\}),\\ &\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,2,3\}),\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,\dots,4\})\big{)}\quad\text{and}\end{aligned}$
$\begin{aligned} V_{2}(x,y)=\big{(}&\psi_{1,x}(\mathcal{N}_{x}\cup\{1\}),\psi_{1,x}(\mathcal{N}_{x}\cup\{1\}),\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\}),\\ &\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\}),\psi_{1,x}(\mathcal{N}_{x}\cup\{1\}),\psi_{1,x}(\mathcal{N}_{x}\cup\{1\}),\\ &\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\}),\psi_{2,y}(\mathcal{N}_{y}\cup\{2\}),\psi_{2,y}(\mathcal{N}_{y}\cup\{2\}),\\ &\psi_{2,y}(\mathcal{N}_{y}\cup\{2\}),\psi_{2,y}(\mathcal{N}_{y}\cup\{2,4\})\big{)}.\end{aligned}$
Under DGP $(\star)$ and Assumption 5,
$\lVert V_{1}(x,y)-V_{2}(x,y)\rVert_{\infty}\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
where $\lVert\cdot\rVert_{\infty}$ denotes the component-wise maximum. Furthermore, the terms in $V_{2}(x,y)$ with $x$ subscripts are independent of the terms with $y$ subscripts by (SA.1.21).
Remark SA.1.4.
The vector $V_{1}(x,y)$ only involves binomial models, whereas $V_{2}(x,y)$ only involves Poissonized models. By the coupling construction, $\mathcal{N}_{x}$ and $\mathcal{N}_{y}$ only involve nodes with labels 5 or greater, and by definition, $\mathcal{N}_{5;\ell_{n}}$ only involves nodes with labels 5 through $\ell_{n}+5$. Then $\psi_{1,x}(\mathcal{N}_{5;\ell_{n}}\cup\{1,2,3\})$, for example, involves nodes 1, 2, 3, and 5 through $\ell_{n}+5$, so it depends on $(X_{1},Z_{1})\equiv(x,z)$ and $(X_{2},Z_{2})\equiv(y,z_{2})$, and $(X_{3},Z_{3})\equiv(x+r_{n}z,z_{3})$, as well as the types of nodes 5 through $\ell_{n}+5$, which are random.
Remark SA.1.5.
In this lemma, we have four nodes of interest: two “focal” nodes $1,2$ and two “ancillary” nodes $3,4$. Since $r_{n}\rightarrow 0$, node $3$ is local to $1$ and $4$ is local to $2$ in terms of positions, while $1$ and $2$ are a fixed distance $\lVert x-y\rVert$ apart and in this sense non-local. The lemma shows that, in various models with different node sets, node $1$’s statistic is asymptotically uncorrelated with node $2$’s statistic since all terms in $V_{2}(x,y)$ involving node 1 are independent of those involving node 2 by the coupling construction. Node $1$’s statistic asymptotically only depends on nodes $1$ and $3$ and not on nodes $2$ and $4$, and vice versa for node $2$.
Proof of Lemma SA.1.5. This lemma is analogous to Lemma 3.7 of Penrose (2007), but the arguments need modification for reasons discussed in Remark SA.1.2. The key step is establishing an analog of his (3.4), which is accomplished in (SA.1.26) below.
We will only prove convergence of arguably the most complicated of the eleven components in $V_{1}(x,y)$ and $V_{2}(x,y)$, namely that under DGP $(\star)$, for any $\epsilon>0$,
${\bf P}\big{(}\lVert\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,\dots,4\})-\psi_{2,y}(\mathcal{N}_{y}\cup\{2,4\})\rVert_{\infty}>\epsilon\big{)}\rightarrow 0.$ (SA.1.22)
Convergence of the other components follows from the same arguments. We will make use of the following fact, which is a consequence of Assumption 6(b):
$\psi_{m}(i,\tau_{z,r_{n}}\bm{T}_{m},\bm{\zeta}_{m},\lambda_{m})$ $\equiv\psi_{m}\big{(}i,\tau_{z,r_{n}}\bm{T}_{m},\bm{\zeta}_{m},\lambda_{m}(\tau_{z,r_{n}}\bm{T}_{m},\bm{\zeta}_{m})\big{)}$
$=\psi_{m}(i,\tau_{z^{\prime},r_{n}}\bm{T}_{m},\bm{\zeta}_{m},\lambda_{m})\quad\forall z,z^{\prime}\in\mathbb{R}^{d},i\in\mathcal{N}_{m},m\in\mathbb{N}.$ (SA.1.23)
Maintain DGP $(\star)$ in what follows. Let $\bm{X}_{5;m_{n}}\equiv\{X_{i}\}_{i=5}^{m_{n}+4}$, and define for any $R>0$
$\mathcal{E}_{n}(R)=\big{\{}\tau_{y,r_{n}}\bm{X}_{5;m_{n}}\cap B(y,R)=\tau_{y,r_{n}}\mathcal{P}_{nf}^{y}\cap B(y,R)\big{\}}$
and
$\bm{R}^{*}_{2}=\max\big{\{}\bm{R}(2,\tau_{y,r_{n}}\bm{T}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\bm{\zeta}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\lambda_{m_{n}+4}),\\ \bm{R}(2,\tau_{y,r_{n}}\bm{T}_{\mathcal{N}_{y}\cup\{2,4\}},\bm{\zeta}_{\mathcal{N}_{y}\cup\{2,4\}},\lambda_{\lvert\mathcal{N}_{y}\rvert+2})\big{\}}.$
The latter is the larger of the radii of stabilization (Definition 1) corresponding to the node statistics $\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,\dots,4\})$ and $\psi_{2,y}(\mathcal{N}_{y}\cup\{2,4\})$, respectively.
For any $R>0$, define $\mathcal{N}(y,R)=\{j\in\mathbb{N}\colon\tau_{y,r_{n}}X_{j}\in B(y,R)\}$, the set of nodes whose dilated positions lie within the $R$-ball of $y$. This is analogous to the definition of $\mathcal{N}_{n}(\mathcal{B}_{i})$ in Definition 1. By Definition 1, if $y=\bm{0}$, then under the event $\{R>\bm{R}^{*}_{2}\}$,
$\psi_{2,y}\big{(}\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}\big{)}=\psi_{2,y}\big{(}(\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\})\cap\mathcal{N}(y,R)\big{)}.$ (SA.1.24)
By (SA.1.23), this also holds for $y\neq\bm{0}$.
Recall that under DGP $(\star)$, the positions of nodes $1$, 33, and $4$ are $x$, $x+r_{n}z$, and $y+r_{n}w$, respectively. Since $\tau_{y,r_{n}}(\{x,x+r_{n}z,y+r_{n}w\})=\{y+r_{n}^{-1}(x-y),y+r_{n}^{-1}(x+r_{n}z-y),w+y\}$, the intersection of this set and $B(y,R)$ is $\{w+y\}$ for $n,R$ sufficiently large. Then
[…]
for such $n,R$ under $\{R>\bm{R}^{*}_{2}\}$. Furthermore, under the event $\mathcal{E}_{n}(R)\cap\{R>\bm{R}^{*}_{2}\}$,
[…]
where the first equality uses the event $\mathcal{E}_{n}(R)$, which allows us to replace the “binomial” model with the “Poissonized” model, and the second equality follows from the argument for (SA.1.24). Therefore, by the law of total probability,
${\bf P}\big{(}\lVert\psi_{2,y}(\mathcal{N}_{5;m_{n}}\cup\{1,\dots,4\})-\psi_{2,y}(\mathcal{N}_{y}\cup\{2,4\})\rVert_{\infty}>\epsilon\big{)}\\ \leq{\bf P}(\mathcal{E}_{n}(R)^{c})+{\bf P}(\bm{R}^{*}_{2}>R),$ (SA.1.26)
where $\mathcal{E}_{n}(R)^{c}$ is the complement of $\mathcal{E}_{n}(R)$. It remains to show that the right-hand side can be made arbitrarily small by choosing $n,R$ large enough.
We first claim that by Assumption 5, $\bm{R}^{*}_{2}=O_{p}(1)$. This is not quite immediate from the assumption because, for instance, $\bm{R}^{*}_{2}$ is defined in terms of the point sets $\tau_{y,r_{n}}\bm{X}_{5;m_{n}}$ and $\tau_{y,r_{n}}\mathcal{P}_{nf}^{y}$ rather than $r_{n}^{-1}\bm{X}_{5;m_{n}}=\tau_{\bm{0},r_{n}}\bm{X}_{5;m_{n}}$ and $r_{n}^{-1}\mathcal{P}_{nf}^{y}=\tau_{\bm{0},r_{n}}\mathcal{P}_{nf}^{y}$ as in the assumption. Invoking (SA.1.23) with $z=y$ and $z^{\prime}=\bm{0}$, we have
$\psi_{m_{n}+4}(2,\tau_{y,r_{n}}\bm{T}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\bm{\zeta}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\lambda_{m_{n}+4})\\ =\psi_{m_{n}+4}(2,\tau_{\bm{0},r_{n}}\bm{T}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\bm{\zeta}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\lambda_{m_{n}+4}).$
Consequently,
$\bm{R}(2,\tau_{y,r_{n}}\bm{T}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\bm{\zeta}_{\mathcal{N}_{m_{n}+4}\cup\{1,\ldots,4\}},\lambda_{m_{n}+4})\\ =\bm{R}(2,\tau_{\bm{0},r_{n}}\bm{T}_{\mathcal{N}_{5;m_{n}}\cup\{1,\ldots,4\}},\bm{\zeta}_{\mathcal{N}_{m_{n}+4}\cup\{1,\ldots,4\}},\lambda_{m_{n}+4}).$
Since $r_{n}^{-1}z=\tau_{\bm{0},r_{n}}z$, the right-hand side of the previous display is $O_{p}(1)$ by Assumption 5. By the same argument,
$\bm{R}(2,\tau_{y,r_{n}}\bm{T}_{\mathcal{N}_{y}\cup\{2,4\}},\bm{\zeta}_{\mathcal{N}_{y}\cup\{2,4\}},\lambda_{\lvert\mathcal{N}_{y}\rvert+2})=\bm{R}(2,\tau_{\bm{0},r_{n}}\bm{T}_{\mathcal{N}_{y}\cup\{2,4\}},\bm{\zeta}_{\mathcal{N}_{y}\cup\{2,4\}},\lambda_{\lvert\mathcal{N}_{y}\rvert+2})=O_{p}(1).$
Hence, $\bm{R}^{*}_{2}=O_{p}(1)$, as desired.
It follows that, for any $\epsilon>0$, we can choose $R$ large enough such that for all $n$ sufficiently large, ${\bf P}(\bm{R}^{*}_{2}>R)<\epsilon/2$. Furthermore, by Lemma SA.1.4, for any such $R$, we can choose $n$ large enough such that ${\bf P}(\mathcal{E}_{n}(R)^{c})<\epsilon/2$. Combining these facts with (SA.1.26), we obtain (SA.1.22).
The next lemma shows that the add-one costs $\Xi_{\ell}$ and $\Xi_{m}$ defined in (14) are asymptotically uncorrelated for $\ell,m$ large.
Lemma SA.1.6.
Let $\{\nu_{n}\}_{n\in\mathbb{N}}\subseteq\mathbb{N}$ be a diverging sequence such that $\nu_{n}/n\rightarrow 0$. Under Assumptions 5 and 6,
$\sup_{n-\nu_{n}\leq\ell<m\leq n+\nu_{n}}\big{|}\,{\bf E}[\Xi_{\ell}\Xi_{m}]-{\bf E}[\Xi_{N_{n}}]^{2}\big{|}\rightarrow 0.$
Proof. We follow the proof of Lemma 5.1 of Penrose (2007). The main difference is that we use our Lemma SA.1.5 in place of their Lemma 3.7. For $i\leq m$, let
$\psi_{m}=\psi_{m}(m,r_{n}^{-1}\bm{T}_{m},\bm{\zeta}_{m},\lambda_{m})\\ \text{and}\quad\Delta_{i,m}=\psi_{m+1}(i,r_{n}^{-1}\bm{T}_{m+1},\bm{\zeta}_{m+1},\lambda_{m+1})-\psi_{m}(i,r_{n}^{-1}\bm{T}_{m},\bm{\zeta}_{m},\lambda_{m}).$
Then
${\bf E}[\Xi_{\ell}$ $\Xi_{m}]={\bf E}\left[\left(\psi_{\ell+1}+\sum_{i=1}^{\ell}\Delta_{i,\ell}\right)\left(\psi_{m+1}+\sum_{j=1}^{\ell}\Delta_{j,m}+\Delta_{\ell+1,m}+\sum_{k=\ell+2}^{m}\Delta_{k,m}\right)\right]$
$\begin{aligned} ={\bf E}[\psi_{\ell+1}\psi_{m+1}]&+\ell\,{\bf E}[\Delta_{1,\ell}\psi_{m+1}]+\ell\,{\bf E}[\psi_{\ell+1}\Delta_{1,m}]+\ell(\ell-1){\bf E}[\Delta_{1,\ell}\Delta_{2,m}]\\ &+\ell\,{\bf E}[\Delta_{1,\ell}\Delta_{1,m}]+{\bf E}[\psi_{\ell+1}\Delta_{\ell+1,m}]+\ell\,{\bf E}[\Delta_{1,\ell}\Delta_{\ell+1,m}]\\ &+(m-\ell-1){\bf E}[\psi_{\ell+1}\Delta_{\ell+2,m}]+\ell(m-\ell-1){\bf E}[\Delta_{1,\ell}\Delta_{\ell+2,m}].\end{aligned}$ (SA.1.27)
By the law of iterated expectations,
${\bf E}[\Xi_{N_{n}}]={\bf E}\left[\psi_{N_{n}+1}+\sum_{i=1}^{N_{n}}{\bf E}[\Delta_{i,N_{n}}\mid\bm{X}_{N_{n}},N_{n}]\right].$
This is the expectation of a pointwise sum (i.e. over the points of $\bm{X}_{N_{n}}$) of functionals of a Poisson point process $\bm{X}_{N_{n}}$. We may then apply Corollary 3.2.3 of Schneider and Weil (2008) and use a change of variables to obtain
${\bf E}[\Xi_{N_{n}}]=\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]\,\text{d}\Phi_{z}(z_{1}\mid x)f(x)\,\text{d}x\\ +\kappa\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]\\ \times\,\text{d}\Phi_{z}(z_{3}\mid x+r_{n}z)\,\text{d}\Phi_{z}(z_{1}\mid x)f(x+r_{n}z)f(x)\,\text{d}z\,\text{d}x,$ (SA.1.28)
where $\bf E$$^{\star}$$\cdot$] denotes an expectation taken under DGP $(\star)$ defined prior to Lemma SA.1.5. Recall that under this DGP, $\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})$ is a function of $(x,z_{1})$ and $\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})$ is a function of $(x,z_{1})$ and $(x+r_{n}z,z_{3})$.
We need to derive the “limits” of each of the terms in (SA.1.27) and show they match with the corresponding term of ${\bf E}[\Xi_{N_{n}}]^{2}$. We only demonstrate for the first two terms as the argument for the others is similar.
Term 1. This term provides the simplest illustration of the logic. Using (SA.1.23) and $m>\ell\geq 1$, the first term ${\bf E}[\psi_{\ell+1}\psi_{m+1}]$ equals
$\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\big{[}\psi_{1,x}(\mathcal{N}_{5;\ell}\cup\{1\})\psi_{2,y}(\mathcal{N}_{5;m-1}\cup\{1,2\})\big{]}\,\text{d}\Phi_{z}(z_{1}\mid x)\,\text{d}\Phi_{z}(z_{2}\mid y)f(x)f(y)\,\text{d}x\,\text{d}y,$
where $\bf E$$^{\star}$$\cdot$] denotes an expectation taken under DGP $(\star)$ defined prior to Lemma SA.1.5. Recall that under this DGP, $\psi_{1,x}(\mathcal{N}_{5;\ell}\cup\{1\})$ is a function of $(x,z_{1})$ and $\psi_{2,y}(\mathcal{N}_{5;m-1}\cup\{1,2\})$ a function of $(x,z_{1})$ and $(y,z_{2})$. Since $\ell,m$ are order $n$, by Lemma SA.1.5,
$\lvert\psi_{1,x}(\mathcal{N}_{5;\ell}\cup\{1\})\psi_{2,y}(\mathcal{N}_{5;m-1}\cup\{1,2\})-\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\psi_{2,y}(\mathcal{N}_{y}\cup\{2\})\rvert\stackrel{{\scriptstyle p}}{{\longrightarrow}}0$
and $\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\perp\!\!\!\perp\psi_{2,y}(\mathcal{N}_{y}\cup\{2\})$. By the Vitali convergence theorem,
$\lvert{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{5;\ell}\cup\{1\})\psi_{2,y}(\mathcal{N}_{5;m-1}\cup\{1,2\})\right]\\ -{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]{\bf E}^{\star}\left[\psi_{2,y}(\mathcal{N}_{y}\cup\{2\})\right]\rvert\rightarrow 0.$
Then using Assumption 6 and the dominated convergence theorem,
$\bigg{|}{\bf E}[\psi_{\ell+1}\psi_{m+1}]-\left(\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]\,\text{d}\Phi_{z}(z_{1}\mid x)f(x)\,\text{d}x\right)^{2}\bigg{|}\rightarrow 0.$
Term 2. Consider the second term on the right-hand side of (SA.1.27). For $m>\ell\geq 1$,
$\ell\,{\bf E}[\Delta_{1,\ell}\psi_{m+1}]=\ell\,{\bf E}\big{[}\left(\psi_{\ell+1}(1,r_{n}^{-1}\bm{T}_{\ell+1},\bm{\zeta}_{\ell+1},\lambda_{\ell+1})-\psi_{\ell}(1,r_{n}^{-1}\bm{T}_{\ell},\bm{\zeta}_{\ell},\lambda_{\ell})\right)\\ \cdot\psi_{m+1}(m+1,r_{n}^{-1}\bm{T}_{m+1},\bm{\zeta}_{m+1},\lambda_{m+1})\big{]}\\ =\ell\,{\bf E}\left[\left(\psi_{1,\bm{0}}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})-\psi_{1,\bm{0}}(\mathcal{N}_{5;\ell-1}\cup\{1\})\right)\psi_{2,\bm{0}}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right].$
By (SA.1.23) and a change of variables, this equals
$\ell\,r_{n}^{d}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\big{(}\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})\right.\\ \left.-\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\big{)}\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right]\\ \times\,\text{d}\Phi_{z}(z_{1}\mid x)\,\text{d}\Phi_{z}(z_{2}\mid y)\,\text{d}\Phi_{z}(z_{3}\mid x+r_{n}z)f(x)f(y)f(x+r_{n}z)\,\text{d}z\,\text{d}x\,\text{d}y,$ (SA.1.29)
where $\bf E$$^{\star}$$\cdot$] denotes an expectation taken under DGP $(\star)$ defined prior to Lemma SA.1.5. Recall that under this DGP, $\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})$, $\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})$, and $\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})$ are functions of $(x,z_{1})$, $(x+r_{n}z,z_{3})$, and/or $(y,z_{2})$. Following the argument for Term 1, by Lemma SA.1.5, Assumption 6, and the Vitali convergence theorem,
${\bf E}^{\star}\left[\big{(}\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\big{)}\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right]\\ -{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]{\bf E}^{\star}\left[\psi_{2,y}(\mathcal{N}_{y}\cup\{2\})\right]\rightarrow 0.$ (SA.1.30)
Next, to apply the dominated convergence theorem, we show that the integrand of (SA.1.29) is uniformly bounded by an integrable function. Under DGP $(\star)$ and the event
$\lVert z\rVert>\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\lambda_{\ell+1}),$
the distance between the scaled positions of nodes 1 and 3 (respectively $r_{n}^{-1}x$ and $r_{n}^{-1}(x+r_{n}z)$) exceeds the right-hand side since $r_{n}^{-1}\lVert x-(x+r_{n}z)\rVert=\lVert z\rVert$. By Definition 1, this implies $\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})=\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})$, in which case the integrand of (SA.1.29) is zero. By Assumption 6 and the Hölder and Minkowski inequalities, there exist $C,\epsilon>0$ such that for $n$ large,
$\lvert{\bf E}^{\star}\left[\big{(}\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\big{)}\psi_{2,y}(\mathcal{N}_{5;m-2}\cup\{1,2,3\})\right]\rvert\\ \leq\big{(}\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1,3\})\rVert_{p}^{\star}+\lVert\psi_{1,x}(\mathcal{N}_{5;\ell-1}\cup\{1\})\rVert_{p}^{\star}\big{)}\cdot\lVert\psi_{2,y}(\mathcal{N}_{5;m-2},\cup\{1,2,3\})\rVert_{p}^{\star}\\ \times{\bf P}^{\star}\big{(}\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{5;\ell-1}\cup\{1,3\}},\lambda_{\ell+1})>\lVert z\rVert\big{)}^{1-2/p}\leq C\tau_{b,\epsilon}(\lVert z\rVert)^{1-2/p},$
where $p$ is defined in Assumption 6 and, for any random variable $W$, $\lVert W\rVert_{p}^{\star}\equiv\sup_{x,z_{1},y,z_{2},z,z_{3}}{\bf E}^{\star}[\lvert W\rvert^{p}]^{1/p}$, and $\bf P$$^{\star}$ denotes the distribution under DGP $(\star)$. By Assumption 5,
$C\ell r_{n}^{d}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\tau_{b,\epsilon}(\lVert z\rVert)^{1-2/p}f(x)f(y)f(x+r_{n}z)\,\text{d}x\,\text{d}y\,\text{d}z\\ \leq C\kappa\sup_{w}f(w)\int_{\mathbb{R}^{d}}\tau_{b,\epsilon}(\lVert z\rVert)^{1-2/p}\,\text{d}z<\infty,$
which verifies the dominance condition. Therefore, by (SA.1.29), (SA.1.30) and dominated convergence,
$\bigg{|}\ell\,{\bf E}[\Delta_{1,\ell}\psi_{m+1}]-\kappa\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{2,y}(\mathcal{N}_{y}\cup\{2\})\right]\,\text{d}\Phi_{z}(z_{2}\mid y)f(y)\,\text{d}y\\ \times\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})-\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]\\ \times\,\text{d}\Phi_{z}(z_{1}\mid x)\,\text{d}\Phi_{z}(z_{3}\mid x+r_{n}z)f(x)f(x+r_{n}z)\,\text{d}z\,\text{d}x\bigg{|}\rightarrow 0.$
The last lemma shows that add-one costs have bounded moments.
Lemma SA.1.7.
Let $\{\nu_{n}\}_{n\in\mathbb{N}}\subseteq\mathbb{N}$ be a diverging sequence such that $\nu_{n}/n\rightarrow 0$. Under Assumptions 5 and 6, for $\tilde{\sigma}_{n}^{2}$ defined in § SA.1.1,
$\limsup_{n\rightarrow\infty}\sup_{n-\nu_{n}\leq m\leq n+\nu_{n}}{\bf E}[\Xi_{m}^{2}]<\infty\quad\text{and}\quad\sup_{n}\max\big{\{}{\bf E}[\Xi_{N_{n}}],\tilde{\sigma}_{n}^{2}\big{\}}<\infty.$
Proof. Recall the characterization of ${\bf E}[\Xi_{N_{n}}]$ in (SA.1.28). The first term on the right-hand side of this equation is uniformly bounded in $n$ by Assumption 6. For the second term, consider the quantity inside $\bf E$$^{\star}$$\cdot$]. Under DGP $(\star)$ and the event $\lVert z\rVert>\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{x}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{x}\cup\{1,3\}},\lambda_{\lvert\mathcal{N}_{x}\rvert+2})$, since node 1 has position $x$ and node 3 has position $x+r_{n}z$, their scaled distance satisfies $r_{n}^{-1}\lVert x-(x+r_{n}z)\rVert>\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{x}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{x}\cup\{1,3\}},\lambda_{\lvert N_{x}\rvert+2})$. Hence, by definition of the radius of stabilization, $\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})=\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})$. Therefore, by Assumption 6 and the Hölder and Minkowski inequalities, there exist $C,\epsilon>0$ such that for $n$ sufficiently large, the integrand of the second term on the right-hand side is bounded above in absolute value by
$\big{(}\lVert\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})\rVert_{p}^{\star}+\lVert\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\rVert_{p}^{\star}\big{)}f(x)f(x+r_{n}z)\\ \times{\bf P}^{\star}\big{(}\bm{R}(1,\tau_{x,r_{n}}\bm{T}_{\mathcal{N}_{x}\cup\{1,3\}},\bm{\zeta}_{\mathcal{N}_{x}\cup\{1,3\}},\lambda_{\lvert N_{x}\rvert+2})>\lVert z\rVert\big{)}^{1-1/p}\\ \leq Cf(x)\sup_{w}f(w)\tau_{p}(\lVert z\rVert)^{1-1/p},$ (SA.1.31)
where for any random variable $W$, $\lVert W\rVert_{p}^{\star}\equiv\sup_{x,z_{1},z,z_{3}}{\bf E}^{\star}[\lvert W\rvert^{p}]^{1/p}$, and $\bf P$$^{\star}$ and $\bf E$$^{\star}$$\cdot$] are respectively the probability and expectation taken under DGP $(\star)$. The right-hand side of (SA.1.31) does not depend on $n$ and has a finite integral over $x,z$ by Assumption 5. Therefore $\sup_{n}{\bf E}[\Xi_{N_{n}}]<\infty$.
The result for ${\bf E}[\Xi_{m}^{2}]$ can be established using similar arguments; see Lemma 5.2 of Penrose (2007). Since the effort of translating the proof to our notation is the same as what was done for the previous two lemmas, we omit the argument.
Finally, similar to (SA.1.28),
$\tilde{\sigma}_{n}^{2}=\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})^{2}\right]\,\text{d}\Phi_{z}(z_{1}\mid x)f(x)\,\text{d}x\\ +\kappa\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d_{z}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\bigg{(}{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1,3\})\psi_{3,x}(\mathcal{N}_{x}\cup\{3,1\})\right]\\ -{\bf E}^{\star}\left[\psi_{1,x}(\mathcal{N}_{x}\cup\{1\})\right]{\bf E}^{\star}\left[\psi_{3,x}(\mathcal{N}_{x+r_{n}z}\cup\{3\})\right]\bigg{)}\\ \times\,\text{d}\Phi_{z}(z_{1}\mid x)\,\text{d}\Phi_{z}(z_{3}\mid x+r_{n}z)f(x)f(x+r_{n}z)\,\text{d}z\,\text{d}x.$
The first term is uniformly bounded by Assumption 6. The second term is also uniformly bounded by an argument similar to (SA.1.31).
Proof of Theorem SA.1.2. We largely follow the proof of Theorem 2.12 of Penrose (2003) and Theorem 2.3 in Penrose (2007), although the characteristic function argument at the end is different since we avoid deriving the limiting variance. Abbreviate $\alpha_{n}={\bf E}[\Xi_{N_{n}}]$ and $\Lambda_{m}=\sum_{i=1}^{m}\psi_{i}(\mathcal{N}_{m})$. We first prove that
${\bf E}\left[n^{-1}\left(\Lambda_{N_{n}}-\Lambda_{n}-(N_{n}-n)\alpha_{n}\right)^{2}\right]\rightarrow 0.$ (SA.1.32)
The left-hand side equals
$\sum_{m\colon|m-n|\leq n^{3/4}}{\bf E}\left[n^{-1}\left(\Lambda_{m}-\Lambda_{n}-(m-n)\alpha_{n}\right)^{2}\right]{\bf P}(N_{n}=m)\\ +{\bf E}\left[n^{-1}\left(\Lambda_{N_{n}}-\Lambda_{n}-(N_{n}-n)\alpha_{n}\right)^{2}\mathbf{1}\{\lvert N_{n}-n\rvert>n^{3/4}\}\right].$ (SA.1.33)
Let $\varepsilon>0$. For $n$ sufficiently large and $m\in[n,n+n^{3/4}]$,
${\bf E}\left[\left(\Lambda_{m}-\Lambda_{n}-(m-n)\alpha_{n}\right)^{2}\right]={\bf E}\left[\left(\sum_{\ell=n}^{m-1}(\Xi_{\ell}-\alpha_{n})\right)^{2}\right]\\ =\sum_{\ell=n}^{m-1}{\bf E}\left[(\Xi_{\ell}-\alpha_{n})^{2}\right]+\sum_{\ell=n}^{m-1}\sum_{\begin{subarray}{c}k=n\\ k\neq\ell\end{subarray}}^{m-1}{\bf E}\left[(\Xi_{\ell}-\alpha_{n})(\Xi_{k}-\alpha_{n})\right]\\ \leq\varepsilon\sqrt{n}(m-n)+\varepsilon(m-n)^{2},$
where the inequality follows from Lemmas SA.1.6 and SA.1.7. A similar argument also applies to $m\in[n-n^{3/4},n]$. Then the first term of (SA.1.33) is bounded by
$n^{-1}{\bf E}\left[\varepsilon\,n^{1/2}\lvert N_{n}-n\rvert+\varepsilon(N_{n}-n)^{2}\right]\leq n^{-1}\varepsilon\left(n^{1/2}\left({\bf E}\lvert N_{n}-n\rvert^{2}\right)^{1/2}+n\right)=2\varepsilon.$
Following the argument in the proof of Theorem 2.3 in Penrose (2007), the second term of (SA.1.33) can be bounded by a constant times
$n^{-1}n^{(2p+1)/p}{\bf P}\left(\lvert N_{n}-n\rvert>n^{3/4}\right)^{1-2/p^{\prime}}$
for some $p^{\prime}\in(2,p)$ and $p$ defined in Assumption 6. This argument uses Assumption 6 and Lemma SA.1.2. The previous display tends to zero since the probability decays exponentially with $n$ by Lemma 1.4 of Penrose (2003). This establishes (SA.1.32).
Next consider the identity
$n^{-1/2}\Lambda_{N_{n}}=n^{-1/2}\Lambda_{n}+n^{-1/2}(N_{n}-n)\alpha_{n}+n^{-1/2}\left(\Lambda_{N_{n}}-\Lambda_{n}-(N_{n}-n)\alpha_{n}\right).$ (SA.1.34)
The variance of the last term tends to zero by (SA.1.32). The second term on the right-hand side has variance $\alpha_{n}^{2}$ and is independent of the first term by the coupling construction. Therefore,
$\frac{1}{n}\text{Var}\left(\Lambda_{N_{n}}\right)=\frac{1}{n}\text{Var}\left(\Lambda_{n}\right)+\alpha_{n}^{2}+o(1)\quad\Rightarrow\quad\sigma_{n}^{2}=\tilde{\sigma}_{n}^{2}-\alpha_{n}^{2}+o(1),$ (SA.1.35)
and $\sup_{n}\sigma_{n}^{2}<\infty$ since $\sup_{n}\max\{\alpha_{n},\tilde{\sigma}_{n}^{2}\}<\infty$ by Lemma SA.1.7.
Since $\sigma_{n}^{2}$ has strictly positive limit infimum by assumption, so does $\tilde{\sigma}_{n}^{2}$. We can then rewrite (SA.1.34) as
$n^{-1/2}\tilde{\sigma}_{n}^{-1}(\Lambda_{N_{n}}-{\bf E}[\Lambda_{N_{n}}])-o_{p}(1)=\frac{\sigma_{n}}{\tilde{\sigma}_{n}}n^{-1/2}\sigma_{n}^{-1}(\Lambda_{n}-{\bf E}[\Lambda_{N_{n}}])+\frac{\alpha_{n}}{\tilde{\sigma}_{n}}n^{-1/2}(N_{n}-n).$ (SA.1.36)
Let $\tilde{\varphi}_{n}(\cdot)$ be the characteristic function of the left-hand side, $\varphi_{n}(\cdot)$ that of $n^{-1/2}\sigma_{n}^{-1}(\Lambda_{n}-{\bf E}[\Lambda_{N_{n}}])$, and $\varphi_{p}(\cdot)$ that of $n^{-1/2}(N_{n}-n)$. By independence of $N_{n}$, due to the coupling construction, the characteristic function of the right-hand side of (SA.1.36) is
$\varphi_{n}\left(\frac{\sigma_{n}}{\tilde{\sigma}_{n}}x\right)\varphi_{p}\left(\frac{\alpha_{n}}{\tilde{\sigma}_{n}}x\right),\quad\text{implying}\quad\varphi_{n}\left(y\right)=\tilde{\varphi}_{n}\left(\frac{\tilde{\sigma}_{n}}{\sigma_{n}}y\right)\varphi_{p}\left(\frac{\alpha_{n}}{\sigma_{n}}y\right)^{-1}$
for $y=\sigma_{n}\tilde{\sigma}_{n}^{-1}x$. By Theorem SA.1.1 and the Poisson CLT,
$\tilde{\varphi}_{n}(x)\rightarrow e^{-x^{2}/2}\quad\text{and}\quad\varphi_{p}(x)\rightarrow e^{-x^{2}/2}.$
Furthermore, this convergence is uniform over $x$ in a compact set (Klenke, 2013, Theorem 15.23). As previously discussed, $\sup_{n}\max\{\alpha_{n},\tilde{\sigma}_{n}\}<\infty$ and $\liminf_{n\rightarrow\infty}\min\{\sigma_{n},\tilde{\sigma}_{n}\}>0$. Thus, for any $y\in\mathbb{R}^{d}$, there exists a compact set $K$ such that for $n$ sufficiently large, $y\tilde{\sigma}_{n}/\sigma_{n}$ and $y\alpha_{n}/\sigma_{n}$ lie in $K$. Combining these facts,
$\varphi_{n}(y)=\frac{\text{exp}\left\{-\frac{\tilde{\sigma}_{n}^{2}}{\sigma_{n}^{2}}\frac{y^{2}}{2}\right\}+o(1)}{\text{exp}\left\{-\frac{\alpha_{n}^{2}}{\sigma_{n}^{2}}\frac{y^{2}}{2}\right\}+o(1)}=\frac{\text{exp}\left\{-\frac{\tilde{\sigma}_{n}^{2}-\alpha_{n}^{2}}{\sigma_{n}^{2}}\frac{y^{2}}{2}\right\}+o(1)}{1+o(1)}\rightarrow e^{-y^{2}/2}$
for any $y\in\mathbb{R}$, where convergence follows from (SA.1.35). Therefore,
$n^{-1/2}\sigma_{n}^{-1}\left(\Lambda_{n}-{\bf E}\left[\Lambda_{N_{n}}\right]\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,1).$
Finally, in the previous expression, we can replace ${\bf E}[\Lambda_{N_{n}}]$ with ${\bf E}[\Lambda_{n}]$ since (SA.1.32) implies that ${\bf E}[n^{-1/2}(\Lambda_{N_{n}}-\Lambda_{n}-(N_{n}-n)\alpha_{n})]\rightarrow 0$.
SA.1.3 Proof of Theorem 1
[…]
Appendix SA.2 Theorem 2
Because we are verifying Assumption 5, we consider a model with nodes $\mathcal{N}_{m}$ and sparsity parameter $r_{n}$ with $m,n\in\mathbb{N}$. Let $\bm{A}=\lambda(r_{n}^{-1}\bm{T}_{m},\bm{\zeta}_{m})$, per Assumption 3. From the same primitives $(r_{n}^{-1}\bm{T}_{m},\bm{\zeta}_{m})$, construct $\bm{D}\equiv\bm{D}(m)$, the network on $\mathcal{N}_{m}$ with no self links and $ij$th entry (15). Let $C_{i}\equiv C_{i}(m)$ be $i$’s component in $\bm{D}(m)$, and define $C_{i}^{+}\equiv C_{i}^{+}(m)$ as in (16). Let $\bm{M}\equiv\bm{M}(m)$ be the network on $\mathcal{N}_{m}$ with no self links and $ij$th entry
$M_{ij}=\bm{1}\big{\{}\sup_{s}V(r_{n}^{-1}\lVert X_{i}-X_{j}\rVert,s,Z_{i},Z_{j},\zeta_{ij})>0\big{\}}.$ (SA.2.1)
This is the network of potential links that are not robustly absent. Finally, for any $J\subseteq\mathcal{N}_{m}$ and $i\in J$, abbreviate
$\psi_{i}(J)\equiv\psi_{\lvert J\rvert}(i,r_{n}^{-1}\bm{T}_{J},\bm{\zeta}_{J},\lambda_{\lvert J\rvert})\equiv\psi_{\lvert J\rvert}\big{(}i,r_{n}^{-1}\bm{T}_{J},\bm{\zeta}_{J},\lambda_{\lvert J\rvert}(r_{n}^{-1}\bm{T}_{J},\bm{\zeta}_{J})\big{)}.$
The following lemma constructs for each node $i$ a set $J_{i}$ satisfying (18).
Lemma SA.2.1 (Leung (2019b), Lemma 1).
Under Assumptions 2, 4, and 8, for any $m,n\in\mathbb{N}$ and $i\in\mathcal{N}_{m}$, $\psi_{i}(\mathcal{N}_{m})=\psi_{i}(J_{i})$ for
$J_{i}\equiv J_{i}(m)=\bigcup\left\{C_{j}^{+}(m)\colon j\in\mathcal{N}_{\bm{M}(m)}(i,K)\right\}.$
[…]
$\tilde{J}_{i}\equiv\tilde{J}_{i}(m)=\bigcup\left\{\tilde{C}_{j}(m)\colon j\in\mathcal{N}_{\bm{M}(m)}(i,K)\right\}.$ (SA.2.2)
SA.2.1 Branching Processes
We next define branching processes used to stochastically bound $\lvert\tilde{C}_{i}\rvert$ and $\lvert\mathcal{N}_{\bm{M}}(i,K)\rvert$. These are used to construct a process whose size stochastically dominates $\lvert\tilde{J}_{i}\rvert$ from (SA.2.2). A branching process is a discrete-time stochastic process in which each time period is associated with a set of elements that are the “offspring” of elements of the previous time period. The process starts at time zero with only one element.
The use of branching processes to bound component and neighborhood sizes is a well-known technique in random graph theory (e.g. Bollobás and Riordan, 2012). The basic idea is to explore the component of a node $i$ using a breadth-first search where at each step we uncover the neighbors of nodes uncovered in the previous step. If the network formation model is dyadic (as is the case for $\bm{D}$ and $\bm{M}$), then the set of nodes uncovered at a given step can be stochastically bounded by the number of offspring produced at a given time period of an appropriate branching process. If the branching process is subcritical in that number of offspring declines over time, then the size of $i$’s component will be stochastically bounded.
Let $x\in\mathbb{R}^{d}$, $z\in\mathbb{R}^{d_{z}}$, $r>0$, and $\bar{f}=\sup_{x\in\mathbb{R}^{d}}f(x)$. Recall the definition of $\Phi^{*}$ from Assumption 7 and $p_{1}(x,z;x^{\prime},z^{\prime})$ from (17). Denote by $\mathfrak{X}^{D}_{r}(x,z)$ the multi-type Galton-Walton branching process (see e.g. Bollobás et al., 2007, §2.1) with type space $\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}$ starting at a particle of type $(x,z)$ with the following two properties. First, each particle of type $(x^{\prime},z^{\prime})$ is replaced in the next generation by a set of particles (its “offspring”) distributed as a Poisson point process on $\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}$ with intensity
$\text{d}\varphi_{r}(x^{\prime},z^{\prime};x^{\prime\prime},z^{\prime\prime})=\kappa\bar{f}(1+r)p_{1}(x^{\prime},z^{\prime};x^{\prime\prime},z^{\prime\prime})\,\text{d}\Phi^{*}(z^{\prime\prime})\,\text{d}x^{\prime\prime}.$ (SA.2.3)
Second, conditional on particle types at generation t1t-1, the Poisson point processes that create the particles of generation $t$ are independent across generation t1t-1 particles.
We can interpret $\mathfrak{X}^{D}_{r}(x,z)$ as a tree network. The root of the tree, corresponding to the first generation of the process, consists of a single node (particle) of type $(x,z)$. Its neighbors are its offspring, which constitute the second generation and have types distributed according to the Poisson point process described above. Conditional on the types of the second generation, each node in the second generation of type $(x^{\prime},z^{\prime})$ independently generates neighbors with types distributed as the Poisson point process above. These constitute the third generation. The process is repeated indefinitely. Let $\lvert\mathfrak{X}^{D}_{r}(x,z)\rvert$ denote the number of particles ultimately generated, which may be infinite. We write $(x^{\prime},z^{\prime})\in\mathfrak{X}^{D}_{r}(x,z)$ to mean that $(x^{\prime},z^{\prime})$ is a type of a particle generated at some point in the process, viewing $\mathfrak{X}^{D}_{r}(x,z)$ as the set of all such particles.
We next define a “fixed-depth” branching process, which is required to terminate at a predetermined generation. Recalling the definitions in Assumption 1, let
$\bar{p}_{r}(x;x^{\prime})={\bf P}\big{(}\bar{V}(r^{-1}\lVert x-x^{\prime}\rVert,\zeta_{ij})>0\}\big{)}.$ (SA.2.4)
Let $\mathfrak{X}^{M}_{r}(x,z;K)$ denote the branching process on $\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}$ starting at a particle of type $(x,z)$ that terminates at generation $K+2$, with the following two properties. First, any particle of type $(x^{\prime},z^{\prime})$ is replaced in the next generation by offspring with types distributed as a Poisson point process on $\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}$ with intensity
$\text{d}\bar{\varphi}_{r}(x^{\prime},z^{\prime};x^{\prime\prime},z^{\prime\prime})=\kappa\bar{f}(1+r)\bar{p}_{1}(x^{\prime};x^{\prime\prime})\,\text{d}\Phi^{*}(z^{\prime\prime})\,\text{d}x^{\prime\prime}.$ (SA.2.5)
Second, conditional on the types of generation t1t-1, the point processes that constitute generation $t$ are independent. Thus, the process is generated the same way as $\mathfrak{X}^{D}_{r}(x,z)$, except the intensity measure is different, and once the $(K+1)$-th generation is born, no further offspring are generated. Define $\lvert\mathfrak{X}^{M}_{r}(x,z)\rvert$ and $(x^{\prime},z^{\prime})\in\mathfrak{X}^{M}_{r}(x,z)$ as we did for $\mathfrak{X}^{D}_{r}(x,z)$.
We use $\lvert\mathfrak{X}^{D}_{r_{n}}(X_{i},Z_{i})\rvert$ and $\lvert\mathfrak{X}^{M}_{r_{n}}(X_{i},Z_{i};K)\rvert$ to stochastically bound $\lvert C_{i}\rvert$ and $\lvert\mathcal{N}_{\bm{M}}(i,K)\rvert$, respectively. The main idea is that the number of offspring of a particle stochastically dominates the degree of a node with the same type.
Lemma SA.2.2.
Let $(x,z)\in\mathcal{T}\equiv\text{supp}(X_{1},Z_{1})$.
  1. (a)
    Under Assumption 1, for any $m,n\in\mathbb{N}$ sufficiently large and $i\in\mathcal{N}_{m}$, $\lvert\mathcal{N}_{\bm{M}(m)}(i,K)\rvert$ is stochastically dominated by $\lvert\mathfrak{X}^{M}_{r_{n}}(x,z;K)\rvert$. The same result holds if we replace $m$ with $N_{n}+1$.
  2. (b)
    Under Assumptions 1, 7, and 9, part (a) holds if we replace $\lvert\mathcal{N}_{\bm{M}(m)}(i,K)\rvert$ with $\lvert C_{i}(m)\rvert$ and $\lvert\mathfrak{X}^{M}_{r_{n}}(x,z;K)\rvert$ with $\lvert\mathfrak{X}^{D}_{r_{n}}(x,z)\rvert$.
Proof. Part (b) follows from Lemma SA.3.1 of Leung (2020). The key argument in the lemma is that we can stochastically bound the degree distribution of any node $j$ in $\bm{D}(m)$, which is conditionally binomial, with the Poisson point process with intensity (SA.2.3). The case $m=N_{n}+1$ is simpler since the degree distribution in this case is exactly Poisson, so it follows from essentially the same arguments. The corresponding results for part (a) are simpler still because we terminate the branching process after $K+1$ generations rather than branching indefinitely. Hence, they also follow from similar arguments. For part (a), the analogous proof of Lemma SA.3.1 of Leung (2020) only needs his Assumption SA.3.2(a), which is implied by our Assumption 1. Our Assumptions 7 or 9 are unnecessary for (a) since these are used to control branching processes with an infinite number of periods.
We lastly construct a “hybrid” branching process
$\mathfrak{X}^{H}_{r_{n}}(X_{i},Z_{i};K)$ (SA.2.6)
whose size stochastically dominates $\lvert\tilde{J}_{i}\rvert$ for sufficiently large $m,n$. First, for any $(x,z)\in\mathcal{T}$, define the smaller process
$\mathfrak{X}_{r_{n}}^{+}(x,z),$
constructed by first generating $\mathfrak{X}^{D}_{r_{n}}(x,z)$ and then for each particle (in any generation) of type $(x^{\prime},z^{\prime})\in\mathfrak{X}^{D}_{r_{n}}(x,z)$ generating a branching process $\mathfrak{X}^{M}_{r_{n}}(x^{\prime},z^{\prime};1)$ independently across particles conditional on $\mathfrak{X}^{D}_{r_{n}}(x,z)$. Using Lemma SA.2.2, $\lvert\mathfrak{X}_{r_{n}}^{+}(X_{i},Z_{i})\rvert$ stochastically dominates $\lvert\tilde{C}_{i}\rvert$.
Next construct (SA.2.6) by generating the fixed-depth process $\mathfrak{X}^{M}_{r_{n}}(X_{i},Z_{i};K)$ and then for each particle (in any generation) of type $(x^{\prime},z^{\prime})$ in $\mathfrak{X}^{M}_{r_{n}}(X_{i},Z_{i};K)$ generating the smaller process $\mathfrak{X}_{r_{n}}^{+}(x^{\prime},z^{\prime})$ independently across particles conditional on $\mathfrak{X}^{M}_{r_{n}}(X_{i},Z_{i};K)$. Using Lemma SA.2.2, $\lvert\mathfrak{X}^{H}_{r_{n}}(X_{i},Z_{i};K)\rvert$ stochastically dominates $\lvert\tilde{J}_{i}(m)\rvert$, and therefore $\lvert J_{i}(m)\rvert$, as desired. The previous claims are also true if we replace $m$ with $N_{n}+1$.
SA.2.2 Tail Bounds
The first two results respectively show that $\lvert\mathfrak{X}^{D}_{r_{n}}(x,z)\rvert$ and $\lvert\mathfrak{X}^{M}_{r_{n}}(x,z;K)\rvert$, defined in § SA.2.1, have distributions with exponential tails. Let $\bar{f}=\sup_{x}f(x)$, and for $\alpha,r>0$, define $g^{\alpha}_{r}(x,z)={\bf E}[\alpha^{\lvert\mathfrak{X}^{D}_{r}(x,z)\rvert}$] and
$\psi_{r}(x,z)=\kappa\bar{f}(1+r)\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d_{z}}}p_{1}(x,z;x^{\prime},z^{\prime})^{2}\,\text{d}\Phi^{*}(z^{\prime})\right)^{1/2}\,\text{d}x^{\prime}.$
Lemma SA.2.3.
Under Assumptions 1, 7, and 9, there exists $\alpha>1$ such that, for $r>0$ sufficiently small, $\sup_{(x,z)\in\mathcal{T}}g^{\alpha}_{r}(x,z)<\infty$.
Proof. We apply Lemma SA.3.3 of Leung (2020), which directly delivers the conclusion. His Assumption SA.3.1 corresponds to our Assumption 7. His Assumption SA.3.2 requires $\inf_{(x,z)\in\mathcal{T}}\psi_{0}(x,z)>0$ and $\sup_{(x,z)\in\mathcal{T}}\psi_{0}(x,z)<\infty$. Assumption 1 implies the second inequality, and Assumption 9 implies the first because
$\lim_{n\rightarrow\infty}\psi_{r_{n}}(x,z)\geq\lim_{n\rightarrow\infty}n\,{\bf E}[p_{r_{n}}(x,z;X_{2},Z_{2})].$
Lemma SA.2.4.
Under Assumption 1, $\sup_{r\leq\kappa}\sup_{(x,z)\in\mathcal{T}}{\bf E}[\alpha^{\lvert\mathfrak{X}^{M}_{r}(x,z;K)\rvert}$$\infty$ for any $\alpha>1$ and $K\in\mathbb{N}$.
Proof. Let $\bar{g}^{\alpha}_{r}(K)$ be the function that maps $(x,z)\in\mathcal{T}$ to ${\bf E}[\alpha^{\lvert\mathfrak{X}^{M}_{r}(x,z;K)\rvert}$] and $\bar{T}_{r}$ the functional that maps $h\colon\mathbb{R}^{d}\times\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}$ to
$(\bar{T}_{r}h)(x,z)=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}h(x^{\prime},z^{\prime})\text{d}\bar{\varphi}_{r}(x,z;x^{\prime},z^{\prime}),$
where $\bar{\varphi}_{r}(\cdot)$ is defined in (SA.2.5). Define $\Gamma\colon h\mapsto\alpha\,\text{exp}\{\bar{T}_{r}(h-1)\}$. A standard branching process argument yields the recursion
$\bar{g}^{\alpha}_{r}(K)(x,z)\\ =\alpha\,\text{exp}\left\{\kappa\bar{f}(1+r)\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}(\bar{g}^{\alpha}_{r}(K-1)(x^{\prime},z^{\prime})-1)\bar{p}_{r}(x;x^{\prime})\,\text{d}\Phi^{*}(z^{\prime})\,\text{d}x^{\prime}\right\}\\ =\alpha\,\text{exp}\{(\bar{T}_{r}(\bar{g}^{\alpha}_{r}(K-1)-1))(x,z)\}=\alpha(\Gamma\bar{g}^{\alpha}_{r}(K-1))(x,z).$ (SA.2.7)
See for example Chapter 6 of Mode (1971). Then for any $(x,z)\in\mathcal{T}$,
$\bar{g}^{\alpha}_{r}(K)(x,z)=\Gamma^{K}\alpha,$ (SA.2.8)
by (SA.2.7) and the fact that $\lvert\mathfrak{X}^{M}_{r}(x,z;0)\rvert=1$.
By (SA.2.7) and (SA.2.8),
$\Gamma\alpha=\alpha e^{\alpha-1}\text{exp}\left\{\kappa(1+r)\bar{f}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d_{z}}}\bar{p}_{r}(x;x^{\prime})\,\text{d}\Phi^{*}(z^{\prime})\,\text{d}x^{\prime}\right\}.$ (SA.2.9)
By Assumption 1, $\sup_{x}\int_{\mathbb{R}^{d}}\bar{p}_{1}(x;x^{\prime})\,\text{d}x^{\prime}<\infty$ since $\bar{p}_{1}(x;x^{\prime})\leq\bar{\Phi}_{\zeta}(\bar{V}^{-1}(\lVert x-x^{\prime}\rVert,0))$ decays exponentially in $\lVert x-x^{\prime}\rVert$. Consequently, there exists $C<\infty$ such that, for any $r\leq\kappa$, (SA.2.9) is bounded by $\alpha e^{\alpha-1}C$. Likewise, $\Gamma^{2}\alpha<\alpha e^{\alpha e^{\alpha-1}C-1}C<\infty$, and repeating this argument, we obtain $\sup_{(x,z)\in\mathcal{T}}\Gamma^{K}\alpha<\infty$. Combined with (SA.2.8), this proves the claim.
Using the previous two lemmas, we establish an exponential tail bound for the distribution of $\lvert\mathfrak{X}^{H}_{r_{n}}(x,z;K)\rvert$ defined in (SA.2.6).
Lemma SA.2.5.
Under Assumptions 1, 7, and 9, there exists $n_{0}\in\mathbb{N}$ such that
$\limsup_{w\rightarrow\infty}w^{-1}\log\left(\sup_{n>n_{0}}\sup_{(x,z)\in\mathcal{T}}{\bf P}\big{(}\lvert\mathfrak{X}^{H}_{r_{n}}(x,z;K)\rvert>w\big{)}\right)<0.$
[…]
By the law of total probability, for any $w^{\prime}>0$,
[…]
For any $\alpha>1$, the second term on the right-hand side equals
${\bf E}\bigg{[}{\bf P}\bigg{(}\sum_{(x^{\prime},z^{\prime})\in\Gamma_{n}(x,z)}\lvert\mathfrak{X}^{M}_{r_{n}}(x^{\prime},z^{\prime};1)\rvert>w\,\bigg{|}\,\Gamma_{n}(x,z)\bigg{)}\mathbf{1}\big{\{}\lvert\Gamma_{n}(x,z)\rvert\leq w^{\prime}\big{\}}\bigg{]}\\ \leq\alpha^{-w}\,{\bf E}\bigg{[}\prod_{(x^{\prime},z^{\prime})\in\Gamma_{n}(x,z)}{\bf E}\left[\alpha^{\lvert\mathfrak{X}^{M}_{r_{n}}(x^{\prime},z^{\prime};1)\rvert}\,\big{|}\,\Gamma_{n}(x,z)\right]\mathbf{1}\big{\{}\lvert\Gamma_{n}(x,z)\rvert\leq w^{\prime}\big{\}}\bigg{]}\\ \leq\alpha^{-w}\,\left(\sup_{n>n_{0}}\sup_{(x^{\prime},z^{\prime})\in\mathcal{T}}{\bf E}\left[\alpha^{\lvert\mathfrak{X}^{M}_{r_{n}}(x^{\prime},z^{\prime};1)\rvert}\right]\right)^{w^{\prime}},$ (SA.2.11)
where the second line uses Markov’s inequality and conditional independence of the branching processes, and the third line uses independence to drop the conditioning.
By Lemma SA.2.4, we can choose $\alpha$ and $n_{0}$ such that the supremum term in (SA.2.11) is bounded above by some $C>1$. Then setting $w^{\prime}=c\,w$ for any $c>0$,
[…]
Choosing $c$ such that $\log\alpha>c\log C$, the second element on the right-hand side of (SA.2.10) is bounded above by $e^{-\beta w}$ for some $\beta>0$. By a similar argument,
${\bf P}\left(\lvert\Gamma_{n}(x,z)\rvert>w^{\prime}\right)\leq\alpha^{-w^{\prime}}\left(\sup_{n>n_{0}}\sup_{(x^{\prime},z^{\prime})\in\mathcal{T}}{\bf E}\left[\alpha^{\lvert\mathfrak{X}^{D}_{r_{n}}(x^{\prime},z^{\prime})\rvert}\right]\right)^{w^{\prime\prime}}\\ +{\bf P}\left(\mathfrak{X}^{M}_{r_{n}}(x,z;K)>w^{\prime\prime}\right).$ (SA.2.12)
By Lemma SA.2.3, we can choose $\alpha,n_{0}$ such that the supremum term is bounded by a finite constant. Setting $w^{\prime\prime}=c^{\prime}w^{\prime}$ for some small enough $c^{\prime}>0$, the first term on the right-hand side can be bounded above by $e^{-\beta^{\prime}w^{\prime}}=e^{-\beta^{\prime}cw}$ for some $\beta^{\prime}>0$, recalling $w^{\prime}=c\,w$. By Lemma SA.2.4, the second term on the right-hand side of (SA.2.12) is $O(e^{-\beta^{\prime\prime}w^{\prime\prime}})$ for some $\beta^{\prime\prime}>0$, uniformly over all $n$ sufficiently large. We have therefore established that (SA.2.10) is $O(e^{-\gamma w})$ for some $\gamma>0$.
Recall the definition of $\lvert\tilde{J}_{i}(m)\rvert$ from (SA.2.2). We next obtain an exponential tail bound for its distribution using the previous lemmas. Let $m,k\in\mathbb{N}$ and $\bm{t}_{k}\in\mathcal{T}^{k}$. Recall that $\tilde{J}_{i}(m+k)$ is a function of node types $r_{n}^{-1}\bm{T}_{m+k}$. Construct $\tilde{J}_{i}^{\star}(m+k;\bm{t}_{k})$ from $\tilde{J}_{i}(m+k)$ by replacing the node types with $r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k})$ defined prior to Assumption 5, so that ${\bf P}(\lvert\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})\rvert>w)$ is a version of the conditional probability ${\bf P}(\lvert\tilde{J}_{1}(m+k)\rvert>w\mid\bm{T}_{k}=\bm{t}_{k})$.
Lemma SA.2.6.
Under Assumptions 1, 7, and 9, for any $k\in\{1,\dots,4\}$ and $k^{\prime}\in\{1,2\}$, there exist $n_{0},\epsilon>0$ such that $\limsup_{w\rightarrow\infty}w^{-1}\max\{\log\tilde{\tau}_{b,\epsilon}(w),\log\tilde{\tau}_{p}(w)\}<0$, where
$\tilde{\tau}_{b,\epsilon}(w)=\sup_{n>n_{0}}\sup_{m\in((1-\epsilon)n,(1+\epsilon)n)}\sup_{\bm{t}_{k}\in\mathcal{T}^{k}}{\bf P}\big{(}\lvert\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})\rvert>w\big{)},$
$\tilde{\tau}_{p}(w)=\sup_{n>n_{0}}\sup_{\bm{t}_{k^{\prime}}\in\mathcal{T}^{k^{\prime}}}{\bf P}\big{(}\lvert\tilde{J}_{1}^{\star}(N_{n}+k^{\prime};\bm{t}_{k^{\prime}})\rvert>w\big{)}.$
[…]
$\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})\subseteq Q_{1}(\{2\})\cup Q_{2}(\{1\}).$ (SA.2.13)
Let $\bm{M}^{-}$ and $\bm{D}^{-}$ be the networks obtained from $\bm{M}$ and $\bm{D}$ respectively by deleting node 1 and its links. Fix $i\in\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})$. If $i=2$, then clearly $i\in Q_{2}(\{1\})$, so suppose $i\neq 2$. Then either (a) $i\in\mathcal{N}_{\bm{M}}(1,K)$ or (b) $i\in\tilde{C}_{j}$ for some $j\in\mathcal{N}_{\bm{M}}(1,K)$. For case (a), if $i\not\in Q_{1}(\{2\})$, then node 2 must lie on a path of length at most $K$ in $\bm{M}$ connecting $i$ and node 1. Therefore $i\in\mathcal{N}_{\bm{M}^{-}}(2,K)$, so $i\in Q_{2}(\{1\})$. For case (b), if $i\not\in Q_{1}(\{2\})$, then there are two possibilities. The first is that nodes 2 and $i$ are linked in $\bm{M}$, in which case $i\in\mathcal{N}_{\bm{M}^{-}}(2,K)$. The second is that node 2 lies on some path in $\bm{D}$ connecting $i$ and some node $j$, in which case nodes 2 and $i$ are connected through a path in $\bm{D}^{-}$. In both cases, $i\in Q_{2}(\{1\})$.
Applying the logic of (SA.2.13) several times yields
$\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})$ $\subseteq Q_{1}(\{3\})\cup Q_{3}(\{1\})$
$\subseteq\big{(}Q_{1}(\{2,3\})\cup Q_{2}(\{1,3\})\big{)}\cup\big{(}Q_{3}(\{1,2\})\cup Q_{2}(\{1,3\})\big{)}.$ (SA.2.14)
As shown in § SA.2.1, for sufficiently large $m,n$, the size of $Q_{1}(\{2,3\})$ (say) is stochastically dominated by $\lvert\mathfrak{X}^{H}_{r_{n}}(x,z;K)\rvert$ defined in (SA.2.6) where $(x,z)$ is node 1’s type under $\bm{t}_{k}$. By Lemma SA.2.5, the distribution of $\lvert\mathfrak{X}^{H}_{r_{n}}(x,z;K)\rvert$ satisfies the desired exponential tail bound.
The next lemma provides a tail bound for the distribution of the maximal distance between the (scaled) position of a node $i\in\mathcal{N}_{m}$ and the positions of nodes in
$\mathcal{N}_{\bm{J}(m)}(i,K,K^{\prime})=\bigcup_{j\in\mathcal{N}_{\bm{M}(m)}(i,K)}\bigcup_{k\in\mathcal{N}_{\bm{D}(m)}(j,K^{\prime})}\mathcal{N}_{\bm{M}(m)}(k,1).$ (SA.2.15)
The relevance of the latter object is its relation to $\tilde{J}_{i}(m)$ defined in (SA.2.2). If $K^{\prime}$ equals the diameter of $\bm{D}(m)$, which is the maximum path distance between any node pair, then $\mathcal{N}_{\bm{J}(m)}(i,K,K^{\prime})=\tilde{J}_{i}(m)$.
Let $m,k\in\mathbb{N}$ and $\bm{t}_{k}\in\mathcal{T}^{k}$. Let $\bm{D}(m+k)$, $\bm{M}(m+k)$, (SA.2.15), and $\tilde{C}_{i}(m+k)$ defined prior to (SA.2.2) be defined for nodes with types $r_{n}^{-1}\bm{T}_{m+k}$. Let $\bm{D}^{\star}(m+k;\bm{t}_{k})$, $\bm{M}^{\star}(m+k;\bm{t}_{k})$, $\mathcal{N}_{\bm{J}^{\star}(m+k;\bm{t}_{k})}(i,K,K^{\prime})$, and $\tilde{C}_{i}^{\star}(m+k;\bm{t}_{k})$ be the respective analogs of these objects obtained by replacing the node types with $r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k})$ defined prior to Assumption 5. Then, for example, ${\bf P}(\lvert\mathcal{N}_{\bm{J}^{\star}(m+k;\bm{t}_{k})}(i,K,K^{\prime})\rvert>w)$ is a version of the conditional probability ${\bf P}(\lvert\mathcal{N}_{\bm{J}(m+k)}(i,K,K^{\prime})\rvert>w\mid\bm{T}_{k}=\bm{t}_{k})$. The lemma below uses this notation for the case $k=1$ with $\bm{t}_{1}=(x,z)\in\mathcal{T}$.
Lemma SA.2.7.
Under Assumptions 1, 7, and 9, for any $\epsilon>0$ and $K\in\mathbb{N}$, there exist constants $n_{0},c>0$ such that for $w,w^{\prime}\rightarrow\infty$ with $w/w^{\prime}\rightarrow\alpha\in(0,\infty)$,
$\max\left\{\sup_{n>n_{0}}\sup_{m\in((1-\epsilon)n,(1+\epsilon)n)}\sup_{(x,z)\in\mathcal{T}}{\bf P}\big{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(m+1;(x,z))}(1,K,w^{\prime})}r_{n}^{-1}\lVert x-X_{i}\rVert>w\big{)},\right.\\ \left.\sup_{n>n_{0}}\sup_{(x,z)\in\mathcal{T}}{\bf P}\big{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}(1,K,w^{\prime})}r_{n}^{-1}\lVert x-X_{i}\rVert>w\big{)}\right\}\\ =O(w^{d+2}\text{exp}\{-c\,w^{1/2}\}).$
Proof. We prove the bound for the term involving $\bm{J}^{\star}(N_{n}+1;(x,z))$. The argument for the other term is almost identical except it does not require use of Corollary 3.2.3 to simplify the expectation in (SA.2.18) below.
Abbreviate $\bm{M}=\bm{M}(N_{n}+1)$ with $ij$th entry $M_{ij}$ and $\bm{D}=\bm{D}(N_{n}+1)$ with $ij$th entry $D_{ij}$. Let $\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,\ell,\ell^{\prime},\ell^{\prime\prime})$ be the set of nodes $k$ for which there exist nodes $i,j$ such that the path distance between 1 and $i$ in $\bm{M}$ is $\ell$, the path distance between $i$ and $j$ in $\bm{D}$ is $\ell^{\prime}$, and the path distance between $j$ and $k$ in $\bm{M}$ is $\ell^{\prime\prime}$. Then
${\bf P}\bigg{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}(1,K,w^{\prime})}$ $r_{n}^{-1}\lVert x-X_{i}\rVert>w\bigg{)}$
$\leq\sum_{\ell=1}^{K}\sum_{\ell^{\prime\prime}=0}^{1}{\bf P}\bigg{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,\ell,0,\ell^{\prime\prime})}r_{n}^{-1}\lVert x-X_{i}\rVert>w\bigg{)}$
$+\sum_{\ell^{\prime}=0}^{w^{\prime}}\sum_{\ell^{\prime\prime}=0}^{1}{\bf P}\bigg{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,0,\ell^{\prime},\ell^{\prime\prime})}r_{n}^{-1}\lVert x-X_{i}\rVert>w\bigg{)}$
$+\sum_{\ell=1}^{K}\sum_{\ell^{\prime}=1}^{w^{\prime}}{\bf P}\bigg{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,\ell,\ell^{\prime},0)}r_{n}^{-1}\lVert x-X_{i}\rVert>w\bigg{)}$
$+\sum_{\ell=1}^{K}\sum_{\ell^{\prime}=1}^{w^{\prime}}{\bf P}\bigg{(}\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,\ell,\ell^{\prime},1)}r_{n}^{-1}\lVert x-X_{i}\rVert>w\bigg{)}.$ (SA.2.16)
We begin by bounding the fourth term on the right-hand side. Under the event that $\max_{i\in\mathcal{N}_{\bm{J}^{\star}(N_{n}+1;(x,z))}^{\partial}(1,\ell,\ell^{\prime},1)}r_{n}^{-1}\lVert x-X_{i}\rVert>w$, there must exist distinct nodes $j_{0}\neq j_{1}\neq\dots\neq j_{\ell+\ell^{\prime}+1}$ with $j_{0}=1$ such that (a) for some $m=1,\dots,\ell+\ell^{\prime}+1$, we have $r_{n}^{-1}\lVert X_{j_{m-1}}-X_{j_{m}}\rVert>w/(\ell+\ell^{\prime})$ where $X_{1}\equiv x$; (b) for all $k=1,\dots,\ell$, we have $M_{j_{k-1}j_{k}}=1$; (c) for all $k^{\prime}=\ell+1,\dots,\ell+\ell^{\prime}$, we have $D_{j_{k^{\prime}-1}j_{k^{\prime}}}=1$; and (d) $M_{j_{\ell+\ell^{\prime}}j_{\ell+\ell^{\prime}+1}}=1$.
By Assumption 1, these facts imply $\zeta_{j_{m-1}j_{m}}>\bar{V}^{-1}(w/(\ell+\ell^{\prime}),0)$ for some $m=1,\dots,\ell+\ell^{\prime}+1$, so
[…]
Let $X_{1}\equiv x$, $\bar{p}_{r_{n}}(\cdot)$ be defined as in (SA.2.4), $p_{r_{n}}(\cdot)$ be defined as in (17), and for $x,x^{\prime}\in\mathbb{R}^{d}$,
$\rho_{n}(x,x^{\prime},w/(\ell+\ell^{\prime}))\equiv{\bf P}\big{(}\zeta_{j_{m-1}j_{m}}>\max\{\bar{V}^{-1}(w/(\ell+\ell^{\prime}),0),\bar{V}^{-1}(r_{n}^{-1}\lVert x-x^{\prime}\rVert,0)\}\big{)}.$
Using the law of iterated expectations conditioning on all node types,
[…]
$\times\prod_{\begin{subarray}{c}k=1\\ k\neq m\end{subarray}}^{\ell}\bar{p}_{r_{n}}(X_{j_{k-1}},X_{j_{k}})\cdot\prod_{\begin{subarray}{c}k^{\prime}=\ell+1\\ k^{\prime}\neq m\end{subarray}}^{\ell+\ell^{\prime}}p_{r_{n}}(X_{j_{k^{\prime}-1}},Z_{j_{k^{\prime}-1}},X_{j_{k^{\prime}}},Z_{j_{k^{\prime}}})$
$\times\big{(}\bm{1}\{m=\ell+\ell^{\prime}+1\}+\bm{1}\{m\neq\ell+\ell^{\prime}+1\}\bar{p}_{r_{n}}(X_{j_{\ell+\ell^{\prime}}},X_{\ell+\ell^{\prime}+1})\big{)}\bigg{]}.$ (SA.2.18)
By Corollary 3.2.3 of Schneider and Weil (2008),
[…]
$\times\int_{\mathbb{R}^{d}}\rho_{n}\big{(}x_{j_{\ell}},x_{j_{\ell}+1},w/(\ell+\ell^{\prime})\big{)}\,nf(x_{\ell+1})\,\text{d}x_{\ell+1}$
$\times\underbrace{\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}}_{\ell^{\prime}-2\text{ times}}\tilde{\bf E}\bigg{[}\prod_{k^{\prime}=\ell+2}^{\ell+\ell^{\prime}}p_{r_{n}}(x_{j_{k^{\prime}-1}},Z_{j_{k^{\prime}-1}},x_{j_{k^{\prime}}},Z_{j_{k^{\prime}}})\bigg{]}\prod_{k^{\prime}=\ell+2}^{\ell+\ell^{\prime}}nf(x_{j_{k^{\prime}}})\,\text{d}x_{j_{k^{\prime}}}$
$\times\int_{\mathbb{R}^{d}}\bar{p}_{r_{n}}(x_{j_{\ell+\ell^{\prime}}},x_{j_{\ell+\ell^{\prime}+1}})\,nf(x_{\ell+\ell^{\prime}+1})\,\text{d}x_{\ell+\ell^{\prime}+1},$
where we abbreviate $\tilde{\bf E}[\cdot]={\bf E}[\cdot\mid X_{j_{\ell}+2}=x_{j_{\ell}+2},\dots,X_{j_{\ell+\ell}}=x_{j_{\ell+\ell^{\prime}}}]$. By repeatedly applying the Cauchy-Schwarz inequality and Assumption 7(a), we can bound this by
$(\ell+\ell^{\prime}+1)\left(\sup_{x}n\int_{\mathbb{R}^{d}}\bar{p}_{r_{n}}(x,y)f(y)\,\text{d}y\right)^{\ell+1}\lVert h^{*}\rVert_{\mathbf{m}}^{\ell^{\prime}-2}\sup_{x}n\,{\bf E}\big{[}\rho_{n}\big{(}x,X_{2},w/(\ell+\ell^{\prime})\big{)}\big{]}.$
Note that $\lVert h^{*}\rVert_{\mathbf{m}}<1$ by the assumption. Using Assumption 1, for sufficiently large $w$ there exist $c_{1}>0$ and $c_{2}>1$ such that this is bounded above by
$(\ell+\ell^{\prime}+1)\cdot c_{1}^{\ell+1}\cdot c_{2}^{-\ell^{\prime}+2}\cdot\kappa\bar{f}\bigg{(}\int_{w/(\ell+\ell^{\prime})}^{\infty}u^{d-1}\bar{\Phi}_{\zeta}(\bar{V}^{-1}(u,0))\text{d}u\\ +\frac{(w/(\ell+\ell^{\prime}))^{d}}{d}\bar{\Phi}_{\zeta}(\bar{V}^{-1}(w/(\ell+\ell^{\prime}),0))\bigg{)}$ (SA.2.19)
by a change of variables to polar coordinates. Since $\bar{\Phi}_{\zeta}(\bar{V}^{-1}(u,0))\leq c_{3}e^{-c_{4}u}$ for some $c_{3},c_{4}>0$ by Assumption 1, the second term involving $\bar{\Phi}_{\zeta}$ is at least as large as the first term for sufficiently $w$ large, as can be seen from integration by parts. Then
[…]
for $w$ sufficiently large and some $C,c>0$ that do not depend on $w,\ell,\ell^{\prime}$.
Therefore, for $w$ large enough, the fourth term on the right-hand side of (SA.2.16) is bounded above by
$C\sum_{\ell=1}^{K}\sum_{\ell^{\prime}=1}^{w^{\prime}}(\ell+\ell^{\prime})(w/(\ell+\ell^{\prime}))^{d}\,\text{exp}\{-c(w/(\ell+\ell^{\prime})-\ell+\ell^{\prime})\}\\ <CK(w^{\prime}+K)w^{d}\sum_{\ell^{\prime}=1}^{w^{\prime}}\text{exp}\{-c(w/(K+\ell^{\prime})+\ell^{\prime}-K)\}\\ \leq C^{\prime}w^{\prime}w^{d}\sum_{\ell^{\prime}=1}^{w^{\prime}}\text{exp}\{-c(w/(K+\ell^{\prime})+\ell^{\prime})\}.$
for some $C^{\prime}>0$ that does not depend on $w,w^{\prime}$. Since $w/(K+\ell^{\prime})+\ell^{\prime}$ is optimized at $\ell^{\prime}=\sqrt{w}-K$, for $w$ large, the right-hand side is bounded above by
$C^{\prime}w^{\prime}w^{d}\cdot w^{\prime}\text{exp}\left\{-c\left(2\sqrt{w}-K\right)\right\}=O(w^{d+2}\text{exp}\{-c\,w^{1/2}\}).$
By a similar argument, the other terms of (SA.2.16) are also of this order.
The main result of this subsection is the following tail bound for a quantity that upper bounds the radius of stabilization. Recall the definitions of $r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k})$ prior to Assumption 5, $\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})$ prior to Lemma SA.2.6, and the objects prior to Lemma SA.2.7. For each $i\in\mathcal{N}_{m+k}$, let $X_{i}(\bm{t}_{k})$ denote node $i$’s position according to the type array $r_{n}^{-1}\bm{T}_{m+k}^{\star}(\bm{t}_{k})$. Then ${\bf P}(\max_{i\in\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w)$ is a version of the conditional probability ${\bf P}(\max_{i\in\tilde{J}_{1}(m+k)}r_{n}^{-1}\lVert X_{1}-X_{i}\rVert>w\mid\bm{T}_{k}=\bm{t}_{k})$ for $\tilde{J}_{i}$ defined in (SA.2.2).
Lemma SA.2.8.
Under Assumptions 1, 7, and 9, for $n,m\in\mathbb{N}$ sufficiently large, $k\in\{1,\dots,4\}$, and $k^{\prime}\in\{1,2\}$, there exist $n_{0},\epsilon>0$ such that
$\limsup_{w\rightarrow\infty}w^{-1/2}\max\{\log\tau_{b,\epsilon}^{*}(w),\log\tau_{p}^{*}(w)\}<0,\quad\text{where}$
$\tau_{b,\epsilon}^{*}(w)$ $=\sup_{n>n_{0}}\sup_{m\in((1-\epsilon)n,(1+\epsilon)n)}\sup_{\bm{t}_{k}\in\mathcal{T}^{k}}{\bf P}\big{(}\max_{i\in\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{)},$
$\tau_{p}^{*}(w)$ $=\sup_{n>n_{0}}\sup_{\bm{t}_{k^{\prime}}\in\mathcal{T}^{k^{\prime}}}{\bf P}\big{(}\max_{i\in\tilde{J}_{1}^{\star}(N_{n}+k^{\prime};\bm{t}_{k^{\prime}})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k^{\prime}})-X_{i}(\bm{t}_{k^{\prime}})\rVert>w\big{)}.$
Proof. We prove the result for $\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})$ with $k=3$. The argument is very similar for the other cases. Let $w^{\prime}=\alpha w$ for some $\alpha>0$. By the law of total probability,
[…]
The first term on the right-hand side decays exponentially in $w$ uniformly over $n,m$ by Lemma SA.2.6.
For the second term, by (SA.2.14),
$\big{\{}\max_{i\in\tilde{J}_{1}^{\star}(m+k;\bm{t}_{k})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$ $\subseteq\big{\{}\max_{i\in Q_{1}(\{2,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$
$\cup\big{\{}\max_{i\in Q_{2}(\{1,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$ $\cup\big{\{}\max_{i\in Q_{3}(\{1,2\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}$
$\cup\big{\{}\max_{i\in Q_{2}(\{1,3\})}r_{n}^{-1}\lVert X_{1}(\bm{t}_{k})-X_{i}(\bm{t}_{k})\rVert>w\big{\}}.$
We proceed to bound
[…]
noting that the argument is identical for analogous probabilities involving the other $Q$ sets. Let $\bm{t}_{1}=(x,z)$. Given $k=3$, the previous expression equals
[…]
By Lemma SA.2.7 and the choice of $w^{\prime}$, $\delta(w)=O(w^{d+2}\text{exp}\{-c\,w^{1/2}\})$ for some $c>0$, so $\limsup_{w\rightarrow\infty}w^{-1/2}\log\delta(w)<0$.
SA.2.3 Proof of Theorem 2
Lemma SA.2.1 constructs for each node $i$, a set $J_{i}$ satisfying (18), which is contained in $\tilde{J}_{i}$ defined in (SA.2.2). Using the branching process results in § SA.2.1 and § SA.2.2, Lemma SA.2.8 derives a tail bound on the distribution of $\max_{i\in\tilde{J}_{1}}r_{n}^{-1}\lVert X_{1}-X_{i}\rVert$. This plus one is an upper bound on the radius of stabilization by (19), which proves the result.